### 06 Mar 2026 v1

# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

Abstract
We prove that embedded infinite plane triangulations in ergodic scale-free environments are close to their circle packing and Riemann uniformization embedding on a large scale, as long as suitable moment and connectivity conditions are satisfied. Ergodic scale-free environments were earlier considered by Gwynne, Miller and Sheffield (2018) in the context of the invariance principles for random walk, and they arise naturally in the study of random planar maps and Liouville quantum gravity.
1 Introduction
A planar map is, roughly speaking, a planar graph where, for each vertex, the cyclic ordering of its edges in a planar embedding is given. Planar maps are studied in many different branches of both mathematics and physics, including combinatorics, probability, geometry, string theory, and conformal field theory.
Given a planar map $\mathcal{M}$, how is it natural to draw it in the plane? In this paper we consider two well-studied ways to embed $\mathcal{M}$ which give a discrete conformal structure to $\mathcal{M}$: One is the Koebe-Andreev-Thurston circle packings, where the vertices of $\mathcal{M}$ are represented by a collection of closed disks with disjoint interior, such that two disks are tangent if and only if their corresponding vertices on $\mathcal{M}$ are adjacent. The second way, which is called the Riemann uniformization embedding, is to view each face of $\mathcal{M}$ as a regular polygon and glue them isometrically along their edges. This defines a Riemann surface $M$, and one uses the Riemann uniformization theorem to draw the vertices and edges of $\mathcal{M}$ via a conformal map $\varphi:M\to\mathbb{C}$.
In this paper, we work with random infinite triangulations under the two aforementioned conformal embeddings. Circle packing for random triangulations have been studied in e.g. [BS01, GN13, ABG+16, GJN19, IM23, BG24], and the Riemann uniformization embeddings for random planar maps have been considered in e.g. [GR13, CUR15]. Our main result is, for a family of infinite plane triangulations in ergodic scale-free environments, they are close to their circle packing and Riemann uniformization embedding on a large scale. Our result holds under a moment condition on the degrees of the triangulation and the regularity of the environment, along with a condition guaranteeing that the embedded triangulation is sufficiently well-connected.
Ergodic scale-free environments were originally considered by Gwynne, Miller and Sheffield [GMS22] where they proved the convergence of random walk to planar Brownian motion. In subsequent work [GMS21] they used this to prove that a particular family of random planar maps known as mated-CRT maps converge under the Tutte embedding to continuum random surfaces arising in the study of Liouville quantum gravity. Liouville quantum gravity (LQG) was introduced by Polyakov [POL81] and is conjectured and proven to describe the scaling limit of several models of random planar maps; see e.g. the book [BP24] for an introduction to the mathematical theory of LQG. Other recent works which establish convergence results of random planar maps to LQG under conformal embedding include [HS23, GMS20, BGS25].
The rest of this section is organized as follows. In Section 1.1 we introduce some backgrounds on circle packings and the Riemann uniformization embedding. In Section 1.2 we introduce ergodic scale-free cell configurations. In Section 1.3 we state our main results, Theorems 1.7 and 1.8. In Section 1.4 we discuss some applications and future directions related to this work. Finally in Section 1.5, we give a detailed outline of our proof of Theorems 1.7 and 1.8 in the rest of the paper.
1.1 Discrete conformal embeddings
Let $\mathcal{G}$ be a locally finite planar (multi)graph with vertex set $\mathcal{V}\mathcal{G}$ and edge set $\mathcal{E}\mathcal{G}$. Multiple edges connecting a pair of vertices are allowed, while loops (i.e., edges starting and ending at the same vertex) are not allowed. We say that a disjoint collection of points $\{z_{v}:v\in\mathcal{V}\mathcal{G}\}\subseteq\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ and a collection of simple (undirected) curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{G}\}$ on $\widehat{\mathbb{C}}$ form a planar embedding of $\mathcal{G}$, if the endpoints of curves $\gamma_{e}$ for $e$ connecting vertices $v,v^{\prime}$ are at $z_{v}$ and $z_{v^{\prime}}$, and different curves $\gamma_{e}$ and $\gamma_{e^{\prime}}$ may only intersect each other at starting and ending points. A planar map $\mathcal{M}$ is an embedding of a planar graph $\mathcal{G}$ modulo orientation-preserving homeomorphisms of $\widehat{\mathbb{C}}$. A face is a connected component of the complement of the planar map. A triangular face is a face with three edges on its boundary. A (planar) triangulation is the union of the vertices and edges of a planar map $\mathcal{M}$, as well as (a subset of) the triangular faces of $\mathcal{M}$. We will work with infinite plane triangulations, where the underlying planar map is one-ended and every face has degree three and belongs to the triangulation, and disk triangulations, where the triangulation is finite and topologically homeomorphic to the unit disk. See Section 2.1 for more details.
A planar map $\mathcal{M}$ can be embedded in $\widehat{\mathbb{C}}$ (as described in the previous paragraph) in a number of different ways. In this paper, we focus on two natural embeddings: the circle packing and the Riemann uniformization embedding. These two embeddings are canonical examples of discrete conformal embeddings. This means in particular that embeddings of the same planar map in different domains induce (or is believed to induce) a map that approximates the conformal map between the two domains. For circle packing, this was conjectured by Thurston and solved in [RS87] in the uniformly bounded degree case [^1] The paper only considers the case of the hexagonal lattice, but it is commented in [HS96] that the proof in [RS87] also works for uniformly bounded degrees. and by [HS96] in full generality. For Riemann uniformization embedding, this follows from the definition.
Let $\mathcal{M}$ be a simple planar map with vertex set $\mathcal{V}\mathcal{M}$. We say that a collection $\mathcal{P}=(D_{v}:v\in\mathcal{V}\mathcal{M})$ of closed disks with disjoint interiors is a circle packing for $\mathcal{M}$, if two disks $D_{v}$ and $D_{v^{\prime}}$ are tangent if and only if the vertices $v$ and $v^{\prime}$ are connected by an edge. This gives an embedding of the planar map, with the vertices embedded at the centers of the disks and the edges are straight line segments connecting the centers of tangent disks. By the celebrated Koebe-Andreev-Thurston circle packing theorem, any finite simple planar map has a circle packing, and for disks triangulations, there exists a unique circle packing in the unit disk $\mathbb{D}$ up to Möbius transforms. For infinite plane triangulations, the existence of a circle packing is a straightforward consequence of the tightness from the Ring Lemma [RS87] (Lemma 3.1), and the uniqueness up to Möbius transforms was proved by Schramm [SCH91]. In this case, there is a dichotomy: an infinite plane triangulation $\mathcal{T}$ is circle packing parabolic (resp. circle packing hyperbolic), if there exists a circle packing $\mathcal{P}$ for $\mathcal{T}$, such that the union of all the faces and edges (a.k.a. the carrier of $\mathcal{P}$) is equal to the whole plane (resp. unit disk). In [HS95], it was further shown that, if the simple random walk on $\mathcal{T}$ is recurrent, then $\mathcal{T}$ is circle packing parabolic, while if $\mathcal{T}$ has uniformly bounded degree and the simple random walk on $\mathcal{T}$ is transient, then $\mathcal{T}$ is circle packing hyperbolic. There are numerous applications of circle packings. For instance, circle packing is closely related to He and Schramm’s work on the Koebe conjecture [HS93], and has played an important role on the proof of recurrence of random walks on random graphs [BS01, GN13]. See also [STE05, NAC20] for a comprehensive introduction on circle packing and related topics.
A second natural way to embed random planar maps is through the Riemann uniformization theorem. Roughly speaking, for a triangulation $\mathcal{T}$, one associate each face with an equilateral triangle and glue them isometrically along their edges to get a Riemann surface $M(\mathcal{T})$; see Section 2.1 for details. If $\mathcal{T}$ is an infinite plane triangulation, there are again two possible cases, where the Riemann surface $M(\mathcal{T})$ is parabolic (resp. hyperbolic) if it is conformally equivalent to the whole plane (resp. the unit disk). As commented in [GR13], similar to the circle packing case, this is equivalent to the recurrence/transience of the Brownian motion on the surface $M(\mathcal{T})$. See e.g. [GR13, CUR15] for some works in this direction.
There exist a number of discrete conformal embeddings beyond the two discussed above, including the Tutte embedding, the Smith embedding [BSS+40], the Cardy-Smirnov embedding [HS23], and the t-embedding/Coulomb gauge/s-embedding [CHE18, CLR23, KLR+22].
1.2 Ergodic scale-free cell configurations
In [GMS22], Gwynne, Miller and Sheffield introduced a certain type of ergodic scale-free cell configurations, and proved the convergence of random walks on such cell configurations to Brownian motion in quenched sense, which is an important step towards proving the convergence of random planar maps to LQG under Tutte embedding. In this subsection, we introduce cell configurations [GMS22], as well as a number of conditions we need in our main results.
Definition $1.$
A cell configuration on $\mathbb{C}$ consists of the following objects.
  1. 1.
    A locally finite collection $\mathcal{H}$ of compact connected subsets of $\mathbb{C}$ (“cells”) with non-empty interiors whose union is all of $\mathbb{C}$ and such that the intersection of any two elements of $\mathcal{H}$ has zero Lebesgue measure.
  2. 2.
    A planar map $\mathcal{M}$, which we call the associated map of $\mathcal{H}$, with vertices $\{v_{H}:H\in\mathcal{H}\}$. We require that if $v_{H}$ and $v_{H^{\prime}}$ are connected by an edge of $\mathcal{M}$ then $H\neq H^{\prime}$ and $H\cap H^{\prime}\neq\emptyset$.
  3. 3.
    A function $\mathfrak{c}=\mathfrak{c}_{\mathcal{H}}$ (“conductance”) on the edges $\mathcal{M}$ with values in $(0,\infty)$.
We say $H\sim H^{\prime}$ if $H$ and $H^{\prime}$ are connected by at least one edge in $\mathcal{M}$. Note that we use a different definition of a cell configuration than in [GMS22, Definition 1.1], where a cell configuration is defined to be a tuple $(\mathcal{H},\sim,\mathfrak{c})$. We need to also include $\mathcal{M}$ in the definition since our goal is to prove closeness between $\mathcal{H}$ and conformal embeddings of the map $\mathcal{M}$ and, furthermore, will work with settings where $\mathcal{M}$ has multiple edges. We will typically slightly abuse notation by making the associated map $\mathcal{M}$ and the function $\mathfrak{c}$ implicit, so we write $\mathcal{H}$ instead of $(\mathcal{H},\mathcal{M},\mathfrak{c})$.
For $A\subseteq\mathbb{C}$, let $\mathcal{H}(A)=\{H\in\mathcal{H}:H\cap A\neq\emptyset\}$, and we view $\mathcal{H}(A)$ as an edge-weighted planar map with edge set $\mathcal{E}\mathcal{H}(A)$ consisting of all of the edges in $\mathcal{E}\mathcal{H}$ joining elements in $\mathcal{H}(A)$ and conductances given by the restriction of $\mathfrak{c}$. We also define a metric on the space of cell configurations by
$\begin{split}\mathbb{d}^{\rm CC}(\mathcal{H},\mathcal{H}^{\prime}):=\int_{0}^{\infty}e^{-r}\wedge\inf\Big\{\sup_{z\in\mathbb{C}}|z-f_{r}(z)|+\sup_{e\in\mathcal{E}\mathcal{H}(B(0;r))}\big|\mathfrak{c}(e)-\mathfrak{c}^{\prime}(f_{r}(e))\big|\Big\}\,dr,\end{split}$ (1.1)
where the infimum is over all homeomorphisms $f_{r}:\mathbb{C}\to\mathbb{C}$ such that $f_{r}$ takes each cell in $\mathcal{H}(B(0;r))$ to $\mathcal{H}^{\prime}(B(0;r))$ and preserves the associated map when restricted to cells intersecting $B(0;r)$, and $f_{r}^{-1}$ does the same with $\mathcal{H}$ and $\mathcal{H}^{\prime}$ reversed. Here $f_{r}(\mathfrak{e})$ is understood as the image of the edge $e$ under map isomorphisms. We will typically be working with a random cell configuration, i.e., a random variable taking values in the space of cell configurations equipped with the Borel $\sigma$-algebra generated by the above metric.
Definition 1.2.
A random cell configuration $\mathcal{H}$ is translation invariant modulo scaling if it satisfies the following condition. There is a (possibly random and $\mathcal{H}$-dependent) increasing sequence of open sets $U_{j}\subseteq\mathbb{C}$, each of which is either a square or a disk, whose union is all of $\mathbb{C}$, such that the following is true. Conditional on $\mathcal{H}$ and $U_{j}$, let $z_{j}$ for $j\in\mathbb{N}$ be sampled uniformly from Lebesgue measure on $U_{j}$. Then there exist random numbers $C_{j}>0$ (possibly depending on $\mathcal{H}$ and $z_{j}$) such that $C_{j}(\mathcal{H}-z_{j})\to\mathcal{H}$ in law with respect to the metric (1.1).
In [GMS22], the authors introduced four more conditions that are equivalent to the translation invariant invariant modulo scaling condition in Definition 1.2; see [GMS22, Definition 1.2 and Lemma 2.3] along with Lemma 2.3 below.
Definition 1.3.
Let $\mathcal{H}$ be a random cell configuration that is translation invariant modulo scaling in the sense of Definition 1.2. We say $\mathcal{H}$ is ergodic modulo scaling, if for every real-valued measurable functions $F=F(\mathcal{H})$ which is invariant under translation and scaling, i.e., $F(C(\mathcal{H}-z))=F(\mathcal{H})$ for each $z\in\mathbb{C}$ and $C>0$, is a.s. equal to a deterministic constant.
We emphasize that, in the definition of cell configurations, it is allowed that two cells $H$ and $H^{\prime}$ have nonempty intersections while $H$ and $H^{\prime}$ are not connected by any edge in $\mathcal{E}\mathcal{H}$. See Figure 1 for an illustration. On the other hand, we will need the following property on the connectedness of $\mathcal{H}$.
Definition 1.4.
A random cell configuration $\mathcal{H}$ is connected along macroscopic lines, if almost surely, for each given $\varepsilon>0$, there exists $R>0$ such that for $r>R$ and each horizontal or vertical line segment $\ell\subseteq B(0;r)$ with length at least $\varepsilon r$, the subgraph of $\mathcal{H}$ induced by the set of cells which intersect $\ell$ is connected.
In [GMS22], it is shown that for cell configurations satisfying the constraints in Definition 1.3 and Definition 1.4, along with a moment condition, random walks on $\mathcal{H}$ converges to the Brownian motion (see Theorem 3.5). This will be applied to the proof of the closeness of $\mathcal{H}$ and its circle packing, where in Section 3.2 we establish a variant of this result for a certain conductance from circle packing.
1.3 Main results
In this subsection, we state our main results on the closeness between cell configurations and their discrete conformal embeddings. We work on the case where the associated map $\mathcal{M}$ of $\mathcal{H}$ is an infinite plane triangulation. We emphasize that the cells in $\mathcal{H}$ do not need to be simply connected, and different cells may cross each other; see Figure 1. However, we do need the following almost planarity assumption, which says that $\mathcal{H}$ is close to some planar embedding of $\mathcal{M}$. In the rest of the paper, without specific mentioning, we will always assume that all the conductances are equal to 1, i.e., $\mathfrak{c}(e)=1$ for every edge $e\in\mathcal{E}\mathcal{H}$.
Figure 1: An illustration of five cells in a cell configuration $\mathcal{H}$. The cells in $\mathcal{H}$ may not be simply connected and could cross from each other. We allow that the green/blue and red/purple cells to touch each other while not being connected by an edge in $\mathcal{E}\mathcal{H}$.
Definition 1.5.
Let $\mathcal{H}$ be a cell configuration as in Definition $1.$ We say $\mathcal{H}$ is almost planar, if there exists a collection of points $\{z_{H}:H\in\mathcal{H}\}$ and simple curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in $\mathbb{C}$ forming a planar embedding of the associated map of $\mathcal{H}$, and almost surely as $r\to\infty$,
$\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}\max_{e}\,\operatorname{dist}(\gamma_{e},H)\to 0,$ (1.2)
where the second maximum is over edges with $H$ being an endpoint and the distance is the Hausdorff distance.
Definition 1.6.
[…]
We write $H_{0}$ for the (a.s. unique) cell in $\mathcal{H}$ that contains 0, and $\deg(H_{0})$ for the degree of $H_{0}$ in the associated map.
Theorem 1.7.
Let $\mathcal{H}$ be a random cell configuration such that its associated map is a simple infinite plane triangulation. Suppose $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound
$\mathbb{E}\Big[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\deg(H_{0})^{4}\Big]<\infty.$ (1.3)
Then $\mathcal{H}$ is a.s. circle packing parabolic, and there exists a circle packing $\mathcal{P}$ for $\mathcal{H}$ and a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ such that the following is true. For $H\in\mathcal{H}$, let $c(H)=\frac{1}{\mathrm{area}(H)}\int_{H}z\,dz$ be the Euclidean center of $H$, and $o(H)$ be the center of the disk for $H$ in $\mathcal{P}$. Then almost surely,
$\lim_{r\to\infty}\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}|\mathbf{A}c(H)-o(H)|=0.$ (1.4)
Furthermore, if $\mathcal{H}$ satisfies the rotation invariance in Definition 1.6, then $\mathbf{A}$ can be taken to be the identity matrix.
Theorem 1.8.
Let $\mathcal{H}$ be a random cell configuration such that its associated map is an infinite plane triangulation. Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound (1.3). Then the Riemann surface $M(\mathcal{H})$ formed by the associated map of $\mathcal{H}$ is a.s. parabolic, and there exists a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and a conformal map $\varphi_{0}:M(\mathcal{H})\to\mathbb{C}$ for $\mathcal{H}$ such that the following is true. For $H\in\mathcal{H}$, let $c(H)=\int_{H}z\,dz$ be the Euclidean center of $H$, and $v_{H}$ be the vertex for $H$ in $M(\mathcal{H})$. Then almost surely,
$\lim_{r\to\infty}\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}|\mathbf{A}c(H)-\varphi_{0}(v_{H})|=0.$ (1.5)
Furthermore, for each $e\in\mathcal{E}\mathcal{H}$, we write $\mathfrak{e}_{e}$ for the edge for $e$ on $M(\mathcal{H})$,
$\lim_{r\to\infty}\frac{1}{r}\max_{e\in\mathcal{E}\mathcal{H}(B(0;r))}\mathrm{diam}(\varphi_{0}(\mathfrak{e}_{e}))=0.$ (1.6)
Finally, if $\mathcal{H}$ further satisfies the rotation invariance in Definition 1.6, then $\mathbf{A}$ can be taken to be the identity matrix.
In Section 5.2, we will prove that Theorem 1.7 also holds for general planar maps (non-triangulations) for a particular choice of the circle-packing and under a moment condition, where we recall that circle packings are typically not unique for general planar maps. In Section 5.3, we will show that Theorem 1.8 also holds if $\mathcal{M}$ is a $p$-angulation for any $p\geq 3$, where we recall that a $p$-angulation is a planar map where all faces have degree pp.
We comment that the convergence results of random planar maps under discrete conformal embeddings in [GMS20, GMS21, HS23, BGS25] are all stated for finite planar maps, whereas Theorems 1.7 and 1.8 are stated for infinite plane triangulations. In our subsequent work, we will apply Theorems 1.7 and 1.8 to work on the convergence of random planar maps under circle packing and Riemann uniformization embedding in both finite volume and infinite volume setting. More precisely, we will use a version of Theorems 1.7 and 1.8 (Theorem 5.1) with a weaker version of the line connectivity in Definition 1.4, which will be proved in Section 5.1.
1.4 Applications and future directions
In this section, we discuss some applications and future directions for this work. We start with the following two simple examples, which provide applications of Theorems 1.7 and 1.8. We believe the results in these examples are new, except that a finite-volume version of Example 1.9 for circle packings was proven in [IM23].
Example 1.9 (Voronoi tessellation of a homogeneous Poisson point process).
Let $\lambda>0$, and consider a Poisson point process $\Lambda$ on $\mathbb{C}$ with intensity measure equal to $\lambda$ times the Lebesgue measure on $\mathbb{C}$. Let $\mathcal{H}$ be the cell configuration induced by the Voronoi tessellation associated with $\Lambda$, where the cells are the Voronoi cells, and two cells are connected by an edge in $\mathcal{E}\mathcal{H}$ if and only if they have nonempty intersection. Almost surely, there does not exist four points in $\Lambda$ that are jointly on a circle, so the associated map of $\mathcal{H}$ is a.s. a simple infinite plane triangulation. Since $\Lambda$ is translation invariant in law, one can verify that the conditions in Theorem 1.7 and Theorem 1.8 holds. To see that the moment bound (1.3) holds, one can first prove that $\mathbb{P}[\mathrm{area}(H_{0})\leq s]\leq cs^{4/3}$ for some constant $c>0$. On the event
$E:=\{\Lambda\cap B(M^{1/4}(m+ni);2M^{1/4})\neq\emptyset,\text{ for all }m,n\in[-10,10]\cap\mathbb{Z}\}\cap\{\#\Lambda\cap B(0;100M^{1/4})\leq M^{3/4}\},$
we have $\mathrm{diam}(H_{0})\leq M$ and $\deg(H_{0})\leq M$. On the other hand, one can check that for some constants $c_{1},c_{2},c_{3}>0$, we have $\mathbb{P}[E]\geq 1-c_{1}e^{-c_{2}M^{c_{3}}}$, and thus
$\mathbb{E}\big[\frac{1}{\mathrm{area}(H_{0})^{p}}\big]<\infty,\ \ \mathbb{E}\big[\mathrm{diam}(H_{0})^{q}\big]<\infty,\ \ \mathbb{E}\big[\deg(H_{0})^{q}\big]<\infty,\ \ \text{ for $1\leq p<4/3$ and every $q>0$},$
which further implies (1.3). Therefore by Theorem 1.7 and Theorem 1.8, the circle packing of $\mathcal{H}$ and the Riemann uniformization embedding of $\mathcal{H}$ must be close to $\mathcal{H}$ on a large scale.
Example 1.10 (Percolation clusters).
Consider a hexagon lattice. Uniformly pick some point $w$ in some hexagon and shift the lattice by $w$. Consider an independent Bernoulli percolation with probability $p$ on the edges of the dual lattice. Let $\mathcal{H}_{0}$ be the cell configuration where the cells are given by the unions of the (closed) hexagons in each cluster. Assume $p$ is subcritical such that there is a.s. no infinite cluster. For each cell $H$ in $\mathcal{H}_{0}$, we join $H$ with the hexagons that are separated from infinity by $H$ and let $\mathcal{H}$ be the cell configuration we obtain. Here for each connected component of $H_{1}\cap H_{2}$ with $H_{1}\neq H_{2}$ and $H_{1}\cap H_{2}\neq\emptyset$, we connect $H_{1}$ and $H_{2}$ by an edge crossing this component of $H_{1}\cap H_{2}$, defining an associated planar map with possibly multiple edges. Then the associated map of $\mathcal{H}$ is an infinite plane triangulation. Using the exponential decay over $n$ for the probability that 0 is connected to hexagons intersecting $\partial B(0;n)$, translation invariance, and a union bound, one can check that the moment bound (1.3) holds, and $\mathcal{H}$ satisfies all the conditions in Theorem 1.8, as $\mathcal{H}$ is invariant in law under a rotation of $2\pi/3$. Therefore by Theorem 1.8, the Riemann uniformization embedding of $\mathcal{H}$ is close to $\mathcal{H}$ on a large scale. If we further collapse the double edges, in the sense that for every cell $H\in\mathcal{H}$ that is separated from infinity by different cells $H_{1}$ and $H_{2}$, we independently join $H$ into $H_{1}$ or $H_{2}$ with equal probability, then we get a cell configuration $\widehat{\mathcal{H}}$ whose associated map is a simple infinite plane triangulation. In this case, by Theorem 1.7 and Theorem 1.8, $\widehat{\mathcal{H}}$ is close to its circle packing and Riemann uniformization embedding on a large scale.
An important motivation this work is that we, in future work, will apply the results of this paper to planar maps arising in the study of Liouville quantum gravity. Another future direction is to generalize our results to planar maps which are not triangulations. Let us discuss the latter direction in more detail.
For infinite plane triangulations, the circle packing is unique. For general planar maps, the circle packing is not necessarily unique; however, one can get uniqueness if one imposes further conditions; see e.g. [BS93, SCH92, THU02]. There are several different kinds of conditions that one can impose. One of the simplest ways to define a unique circle packing for general simple planar maps is as follows. Let $\mathcal{M}$ be a one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$. For each face $f\in\mathcal{M}$, we add a vertex $v_{f}$ and add edges to connect $v_{f}$ to the vertices incident to $f$. This turns $\mathcal{M}$ into a simple infinite plane triangulation $\mathcal{T}_{\mathcal{M}}$. We can consider the circle packing for $\mathcal{T}_{\mathcal{M}}$ and remove the disks for all the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. This defines a circle packing for $\mathcal{M}$ and it is the unique, up to Möbius transform, circle packing constructed in this way. In Section 5.2, we will prove a variant of Theorem 1.7 for the version of the circle packing defined here.
On the other hand, for the Riemann uniformization embedding one has existence and uniqueness of the embedding for any planar map. In Section 5.3, we will show that Theorem 1.8 still holds if we assume our planar map is a $p$-angulation for $p=4,5,6,\dots$.
Problem 1.11.
Prove Theorem 1.7 for other variants of circle packings for non-triangulations, and prove Theorem 1.8 for non-$p$-angulations.
Regarding the Riemann uniformization embedding, we expect that general planar maps without loops can be treated by slightly modifying the argument of the current paper, possibly under a suitable probabilistic bound on the face degrees. The key step would be to prove that Lemma 4.2 still holds in this context. Furthermore, we believe it is also possible to treat maps with loops; in this case, it might additionally be necessary to give a different definition of $\phi_{0}$ in Section 4.2.
1.5 Outline
In this section we provide a moderately detailed overview of the content of the rest of the paper, as well as the outline of the proof of Theorems 1.7 and 1.8.
In Section 2, we present various preliminaries. In Section 2.1, we review various properties of circle packings and the Riemann surface associated with triangulations. In Section 2.2, gather some results on the ergodic averages of the ergodic modulo scaling cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2]. We will repeatedly use the following two facts: the maximum diameter of the cells which intersect a large origin-containing square is typically of smaller order than its side length (Lemma 2.6), and the sum of $\mathrm{diam}(H)^{2}\deg(H)^{4}$ over all the cells $H$ intersecting a large square $S$ which is not too far from origin is bounded from above by a deterministic constant times the area of $S$ (Lemma 2.7).
In Section 3, we prove Theorem 1.7. In Section 3.4 we prove that if the random walk on two different embeddings of the same infinite planar map both converge to planar Brownian motion, then the two embeddings must be close. In order to see this, we consider the wrapping around events for the random walk, and use this to prove that under appropriate scaling, the natural homeomorphisms induced by the embeddings are tight on compact sets. Then under the image of the limiting homeomorphism, the trace of a 2D Brownian motion still has the law of the trace of a 2D Brownian motion. In Proposition 3.14 we prove that any limiting homeomorphism must be a linear transform. By the almost planarity assumption it follows that in order to conclude the proof, it is sufficient to prove that the random walk with Dubejko weights on the cell configuration $\mathcal{H}$ and on the circle packing $\mathcal{P}$ of $\mathcal{H}$ both converge to planar Brownian motion.
To prove that the random walk on $\mathcal{H}$ with Dubejko weights converges to Brownian motion, which is handled in Section 3.2, a natural first attempt is to try to directly use the convergence of random walks on $\mathcal{H}$ with generic conductance from [GMS22] (Theorem 3.5). To use this result, one needs to check that the conductance satisfies the translation invariance modulo scaling property as in Definition 1.2, and the corresponding moment bounds (3.8). For the first condition, we use Rodin and Sullivan’s Ring Lemma [RS87] along with the uniqueness of circle packing for triangulations to prove that if triangulations $\mathcal{T}_{n}$ converges to an infinite plane triangulation $\mathcal{T}$ in Benjamini-Schramm topology, then the radii of the disks in the circle packing of $\mathcal{T}_{n}$ also converges to the radii of the disks in the circle packing of $\mathcal{T}$ (Proposition 3.4). For the moment bound condition, the first half of (3.8) is trivial since the Dubejko conductance is always bounded from above by 1. For the second half of (3.8) involving the inverse conductance, if we directly use the Ring Lemma, which bounds the ratio of the radii of tangent disks in a circle packing in terms of degree, then the inverse conductance is bounded from above by $\exp(C\deg(H_{0}))$ for some $C>0$. It is not possible to obtain the second half of (3.8) using this bound; in order to do this we would have needed a stronger version of (1.3) which would have been hard to verify in the relevant applications.
To circumvent the aforementioned issue, in Section 3.1 we establish a variant of the Ring Lemma (Lemma 3.3), where for three mutually tangent disks in a circle packing, the ratio of between the radii of the second smallest and the smallest disks is bounded above by 100 times the square of the degree. The proof of Lemma 3.3 is based on a geometric argument and Descartes’ theorem (Lemma 3.2). With Lemma 3.3, we modify the proof of Theorem 3.5 from  [GMS22], where every edge $\{H,H^{\prime}\}$ with large inverse Dubejko conductance can be replaced by a path with total inverse conductance controlled by $\deg(H)^{2}$. This is carried out in Proposition 3.8, where we are able to reduce the exponential degree bound from the Ring Lemma to a polynomial bound which allows us to apply (1.3). Proposition 3.8 is a variant of Theorem 3.5 from [GMS22] for the case where the conductances are given by the Dubejko weights and where the moment bound (3.8) has been replaced by (1.3).
It remains to show that the random walk with Dubejko weights on the circle packing $\mathcal{P}$ converges to Brownian motion, which is carried out in Section $3.$ Following work by Bou-Rabee and Gwynne [BG24] it is sufficient to check that up to scaling, there are no macroscopic disks in $\mathcal{P}$. In [GJN19], Gurel-Gurevich, Jerison and Nachmias proved that the maximal diameter of the disks in a circle packing can be controlled by the inverse of Cannon’s vertex extremal length [CAN94]. In Proposition 3.13, using this and the averaging over $\mathrm{diam}(H)^{2}$ from Lemma 2.7, we prove that for any number $q_{0}$, there are at most finitely many cells with associated vertex extremal length at most $q_{0}$, which eventually leads to the convergence of the random walk on $\mathcal{P}$.
In Section 4 we prove Theorem 1.8. In [GMS22], Theorem 3.5 is proven by studying discrete harmonic functions on the graph defined by $\mathcal{H}$. Our proof is partially following their strategy, with the difference that we will instead be working with (continuum) harmonic functions defined intrinsically on the manifold $M(\mathcal{H})$.
In Section 4.1 we establish a number of regularity estimates on the Riemann surface $M(\mathcal{H})$ under a conformal map $\varphi$. In particular, in Lemma 4.3 and Lemma 4.4, we prove exact estimates on the continuity of the mapping between $\mathcal{H}$ and $\varphi(M(\mathcal{H}))$. In order to prove this, we assign each vertex $H$ a semi-flower $P_{H}\subseteq M(\mathcal{H})$ (see Figure 6) where the union of the semi-flowers is equal to the whole surface. In Lemma 4.2, we use the Koebe’s distortion theorem to prove that, the ratio between the outradius and inradius of the conformal image $\varphi(P_{H})$ is bounded by a constant times $\deg(H)^{2}$. Another ingredient for the proof of Lemma 4.3 and Lemma 4.4 is a length-area argument, which was introduced by He and Schramm [HS96] in the context of circle packings and originated from the Length-Area Lemma in [RS87]. The terms in our length-area argument will be controlled by the average of $\mathrm{diam}(H)^{2}\deg(H)^{4}$ as in Lemma 2.7. Given Lemma 4.3 and Lemma 4.4, in Lemma 4.9 we prove that the surface $M(\mathcal{H})$ is a.s. parabolic, and in Lemma 4.10 we prove that the mapping between $\mathcal{H}$ and $\varphi(M(\mathcal{H}))$ has polynomial growth.
In Section 4.2 we construct a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$. Similarly as in  [GMS22, Section 2.3], we start from a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ which serves as the (non-injective) a priori embedding of $M(\mathcal{H})$. Much care needs to be taken with the exact definition of $\phi_{0}$ and we choose to define it by mapping each vertex of $M(\mathcal{H})$ to a uniform point in the corresponding cell of $\mathcal{H}$, and then extend to every equilateral triangle in $M(\mathcal{H})$ via linear interpolation. For each $m\in\mathbb{N}$ we let $\phi_{m}$ be equal to $\phi_{0}$ on the boundary of the $\phi_{0}^{-1}$ image of certain large squares, and harmonic elsewhere. The Dirichlet energy of the piecewise linear function $\phi_{0}$ can be controlled via the average in Lemma 2.7, and the Dirichlet energy of $\phi_{m}-\phi_{m^{\prime}}$ is small when $m,m^{\prime}$ is large (Lemma 4.19). By standard properties of harmonic functions, this implies that the $\phi_{m}$’s are Cauchy in probability in the local uniform topology, and the limit $\phi_{\infty}$ of $\phi_{m}$ exists.
In Section 4.3, we prove that the harmonic function $\phi_{\infty}$ is close to $\phi_{0}$ on a large scale (Proposition 4.20). We first follow the same outline as [GMS22, Section 2.4] to prove that $\phi_{\infty}$ is close to $\phi_{0}$ on a set of line segments, where in our setting the line segments are contained in the image of a suitable subset of $M(\mathcal{H})$ under a conformal map. To prove the closeness of $\phi_{\infty}$ to $\phi_{0}$ on the complementary set, the oscillation of $\phi_{\infty}$ in each connected component of the complement is equal to the oscillation of $\phi_{\infty}$ on the boundary of that component thanks to the maximum principle for harmonic functions, whereas the oscillation of $\phi_{0}$ in each connected component can be bounded by the continuity from Lemma $4.$ This shows that $|\phi_{0}-\phi_{\infty}|$ is also small in every connected component of the complementary set, which concludes the proof of Proposition 4.20.
In Section 4.4, we conclude our proof of Theorem 1.8 by showing that $\phi_{\infty}$ has to be the composition of a deterministic linear transform, a random rotation and a scaling. Using the polynomial growth from Lemma 4.10, the harmonic function $\phi_{\infty}\circ\varphi^{-1}$ on $\mathbb{C}$ has polynomial growth and must be a polynomial (Lemma 4.24). Furthermore, in Lemma 4.27, we prove that this polynomial must have degree 1 because if this is not the case, then one can find two cells that are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.3 and Lemma 4.6. In Lemmas 4.29,  4.30 and  4.32, we further use the ergodic theorem for the cell configuration (Lemma 2.4) to show that, this linear map $\phi_{\infty}\circ\varphi^{-1}$ can be written as a composition of a deterministic linear transform $\mathbf{A}$, a random rotation and a scaling, and in Lemma 4.33, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.6, then $\mathbf{A}$ must be a rotation. This finishes the proof of Theorem 1.8. Our proof in Section 4.4 works under quite mild assumptions and gives a general method for proving that a random harmonic function is of a particular form (namely, the composition of a deterministic linear transformation and a random rotation and scaling) as long as it has polynomial growth, is almost injective, and the averages of the Dirichlet energy over large regions are approximately a deterministic constant.
In Section 5.1, we prove Theorem 5.1, a version of Theorem 1.7 and Theorem 1.8 under a weaker version of the line connectivity property. In Section 5.2, we prove a version of Theorem 1.7 for general planar maps (non-triangulations); more specifically, we consider the variant of circle packings mentioned in Section 1.4. Finally, in Section 5.3 we prove the extension of Theorem 1.8 to $p$-angulations.
Acknowledgements. We thank Ewain Gwynne for mentioning to us at an earlier stage of the project that Proposition 3.11 hold under the assumptions of [GMS22], which weakened our initial assumptions. N.H. was supported by grant DMS-2246820 of the National Science Foundation and the Simons Collaboration Grant Probabilistic Paths to Quantum Field Theory.
2 Preliminaries
2.1 Triangulations, circle packings and Riemann surfaces
We follow the papers [BS01, AS03, GR13] for the definition on random plane triangulations. A planar map $\mathcal{M}$ is a proper (without edge-crossings) embedding of a connected planar (multi)graph $\mathcal{G}$ in the two-dimensional sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, where we identify two planar maps if one can be obtained from the other via an orientation-preserving homeomorphism of the sphere. We will always assume that $\mathcal{M}$ is locally finite, i.e., each vertex and each face has finite degree. For a planar map $\mathcal{M}$ on the sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, we call a connected component of $\widehat{\mathbb{C}}\backslash\mathcal{M}$ a triangular face if its boundary is consisted of precisely three edges of $\mathcal{M}$. An embedded (planar) triangulation $\mathcal{T}$ consists of a finite or infinite connected planar map $\mathcal{M}$ in the sphere, together with a subset of the triangular faces, and its support $S(\mathcal{T})\subseteq\widehat{\mathbb{C}}$ is defined to be the union of $\mathcal{M}$ and all the triangular faces in $\mathcal{M}$. Two embedded triangulations $T$ and $T^{\prime}$ are considered equivalent if there is an orientation preserving homeomorphism between their supports which maps vertices (resp., edges, faces) to vertices (resp., edges faces). For convenience we abbreviate “equivalence class of embedded triangulations” to “triangulation”. Following the terminology in [AS03, Section 1.2], we call $\mathcal{T}$ a type II triangulation if $\mathcal{T}$ has no loops but possibly multiple edges, and call $\mathcal{T}$ a type III triangulation if $\mathcal{T}$ has no loops and no multiple edges. We call a vertex $v\in\mathcal{T}$ an interior (resp. boundary) vertex of $\mathcal{T}$ if $v$ is in the interior (resp. on the boundary) of the support $S(\mathcal{T})$. We call a type II (resp. type III) triangulation $\mathcal{T}$ an infinite plane triangulation (resp. simple infinite plane triangulation), if the underlying graph is one-ended, and there are no boundary vertices. We call a type II (resp. type III) triangulation $\mathcal{T}$ a (finite) disk triangulation (resp. simple disk triangulation), if the underlying planar map $\mathcal{M}$ is finite and the support $S(\mathcal{T})$ is homeomorphic to a disk. A rooted triangulation $(\mathcal{T},e)$ is a triangulation $\mathcal{T}$ together with an oriented edge $e=(x,y)$ of $\mathcal{T}$, called the root edge. The vertex $x$ is the root vertex.
[…]
Theorem 2.1 ([HS95]).
Let $\mathcal{T}$ be a simple infinite plane triangulation. If the simple random walk on $\mathcal{T}$ is recurrent, then $\mathcal{T}$ is circle packing parabolic. Moreover, if $\mathcal{T}$ has bounded degree and the simple random walk on $\mathcal{T}$ is transient, then $\mathcal{T}$ is circle packing hyperbolic.
Figure 2: Two adjacent vertices $u,v$. The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\big(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\big)$.
One can also consider random walks on circle packings. Consider the circle packing $\mathcal{P}$ of some simple infinite triangulation $\mathcal{T}$. Let $u,v$ be adjacent vertices in $\mathcal{T}$, and $w_{1},w_{2}$ be the vertices on the two faces of $\mathcal{T}$ containing the edge uvuv other than $u$ and $v$. Let $o_{u},o_{v},o_{w_{1}},o_{w_{2}}$ be the centers of the circles for $u,v,w_{1},w_{2}$, and $o_{a},o_{b}$ be the centers of circles which are inscribed to the triangles $o_{u}o_{v}o_{w_{1}}$ and $o_{u}o_{v}o_{w_{2}}$, respectively. The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}$; see Figure 2 for an illustration. If we write $r_{u},r_{v},r_{w_{1}},r_{w_{2}}$ for the radii of the disks $D_{u},D_{v},D_{w_{1}},D_{w_{2}}$, then
$\mathfrak{c}(u,v)=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\bigg(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\bigg).$ (2.1)
One can infer from (2.1) that $\mathfrak{c}(u,v)<\frac{1}{2}$ for any edge uvuv using $2\sqrt{r_{u}r_{v}}\leq r_{u}+r_{v}$. Furthermore, by Dubejko’s theorem (see e.g. [STE05, Theorem 18.3]), for a circle packing $\mathcal{P}$ of a simple infinite triangulation, if we consider the random walk $(Z_{n})_{n\geq 0}$ moving between the centers of the circles in $\mathcal{P}$, then $(Z_{n})_{n\geq 0}$ is a martingale.
Next we follow the description of [GR13, Section 2.2] to define the Riemann surfaces associated with triangulations. Let $\mathcal{T}$ be a disk triangulation or an infinite plane triangulation. We associate $\mathcal{T}$ a metric space $(M(\mathcal{T}),d_{M})$, by regarding each triangle of $\mathcal{T}$ as an equilateral triangle of side length 1 and defining the distance between two points as the length of the shortest path joining them. Equivalently, $M(\mathcal{T})$ is obtained by gluing equilateral triangles according to their adjacency pattern in $\mathcal{T}$. The metric space $M(\mathcal{T})$ can be equipped with a compatible Riemann surface structure by defining coordinate charts as follows. If $x$ is an interior point of a triangle or interior to an edge on the boundary of the support, then we use this triangle, viewed as a subset of the plane, as the chart, and use the identity map as the projection. Similarly, if $x$ is interior to an edge and not on the boundary of the support, then we place two adjacent equilateral triangles in the plane and use the identity as the projection. Now let $x$ be an interior vertex of degree $n$, and let $\Delta_{1},...,\Delta_{n}$ be the equilaterals which meet at $x$ in cyclic order. We place them in the plane such that for j=1,,nj=1,...,n, $\Delta_{j}$ is the triangle with vertices $0,e^{\pi(j-1)i/3},e^{\pi ji/3}$, and define the coordinate map for $z=re^{i\theta}\in\Delta_{j}$ by $\phi(z)=r^{6/n}e^{6i\theta/n}$, using $\pi(j-1)/3<\theta<\pi j/3$ on $\Delta_{j}$. By further setting $\phi(0)=0$, this defines a homeomorphism between a neighborhood of $x$ in $M(\mathcal{T})$ and a neighborhood of 0 in $\mathbb{C}$. Moreover, for the atlas chosen as above, the transition map between two overlapping charts are analytic. The case where $x$ be an boundary vertex of degree $n$ can be treated similarly using the map $z\mapsto z^{3/(n-1)}$. This defines the Riemann surface structure for $M(\mathcal{T})$. With slight abuse of language, for a vertex $v\in\mathcal{V}\mathcal{T}$, we will also view $v$ as a point on the Riemann surface $M(\mathcal{T})$.
Following the uniformization theorem, for an infinite plane triangulation $\mathcal{T}$, the surface $M(\mathcal{T})$ is either conformally equivalent to the complex plane and thus parabolic, or conformally equivalent to the unit disk and thus hyperbolic. As pointed out by [GR13], an equivalent characterization is that the Brownian motion on the surface $M(\mathcal{T})$ is either recurrent (in the parabolic case) or transient (in the hyperbolic case). A function $f:M(\mathcal{T})\to\mathbb{C}$ is harmonic if and only if $f\circ\varphi^{-1}$ is a harmonic function in usual sense, where $\varphi:M(\mathcal{T})\to\mathbb{C}$ is some conformal map.
2.2 Ergodic averages of cell configurations using dyadic systems
In this subsection we gather some results on the ergodic averages of the cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2]. We start with the definition of dyadic systems and uniform dyadic systems as in [GMS22]. For a square $S$, we write $|S|$ for the side length of $S$. A dyadic child of $S$ is one of the 4 squares with half side length of $S$ whose corners include one corner of $S$ and the center of $S$. A dyadic parent of $S$ is one of the four squares with $S$ as a dyadic child. Then each square has 4 dyadic children and 4 dyadic parents. A dyadic descendant (resp. dyadic ancestor) of $S$ is a square which can be obtained from $S$ by iteratively choosing dyadic children (resp. parents) finitely many times. A dyadic system is a collection of closed squares with their sides parallel to the $x$-axis or $y$-axis with the following properties:
  1. 1.
    If $S\in\mathcal{D}$, then each of the four dyadic children of $S$ is in $\mathcal{D}$.
  2. 2.
    If $S\in\mathcal{D}$, then exactly one of the dyadic parents of $S$ is in $\mathcal{D}$.
  3. 3.
    Any two squares in $\mathcal{D}$ have a common dyadic ancestor.
As explained in [GMS22, Section 2.1], one can define a topology on the space of dyadic systems by looking at the local Hausdorff distance on the union of the origin-containing squares.
Definition $2.$
Let $s$ be uniformly sampled from $[0,1]$, and conditioned on $s$, we uniformly randomly sample $w$ in $0,2^{s}]\times[0,2^{s}]$. Let $S_{0}=[0,2^{s}]\times[0,2^{s}]-w$. For $k\in\mathbb{N}$, inductively let $S_{k}$ be sampled uniformly from the four dyadic parents of $S_{k-1}$. The uniform dyadic system $\mathcal{D}$ is defined by the set of all dyadic descendants of $S_{k}$ for each $k\in\mathbb{N}$. For $z\in\mathbb{C}$ and $m>0$, we define
$\widehat{S}_{m}^{z}:=\{\text{largest square }S\in\mathcal{D}:z\in S\text{ and }\sum_{H\in\mathcal{H}(S)}\frac{\mathrm{area}(H\cap S)}{\mathrm{area}(H)}\leq m\}$ (2.2)
which is well defined on the probability one event that $z$ is not on the boundary of any square in $\mathcal{D}$, and we abbreviate $\widehat{S}_{m}:=\widehat{S}_{m}^{0}$. Finally, for $C>0$ and $z\in\mathbb{C}$, we write $C(\mathcal{D}-z)$ for the collection of the images the squares in $\mathcal{D}$ under the map $w\mapsto C(w-z)$.
In other words, $\widehat{S}_{m}^{z}$ is the largest square containing $z$ which contains at most $m$ cells, with fractional parts of cells which intersect the boundary of the square also counted.
The following is the dyadic system resampling property for random cell configurations.
Lemma 2.3 (Lemma 2.3 of [GMS22]).
The condition in Definition 1.2 on random cell configurations $\mathcal{H}$ is equivalent to the following. Let $\mathcal{D}$ be a uniform dyadic system independent from $\mathcal{H}$ and define $\widehat{S}_{m}=\widehat{S}_{m}^{0}$ for $m>0$ as in (2.2). For each $m>0$, conditional on $\mathcal{H}$ and $\mathcal{D}$, let $z$ be sampled uniformly from the Lebesgue measure on $\widehat{S}_{m}$. Then there exists a random $C>0$ possibly depending on $z$, such that
$(C(\mathcal{H}-z),C(\mathcal{D}-z))\overset{d}{=}(\mathcal{H},\mathcal{D}).$
In the rest of this section, we will restrict to the ergodic modulo scaling cell configurations as in Definition 1.3. The following ergodicity statement will be a key tool in what follows.
Lemma 2.4 (Lemma 2.7 of [GMS22]).
Let $F=F(\mathcal{H},\mathcal{D})$ be a measurable function on the space of cell configuration/dyadic system pairs which is scale invariant (i.e., $F(C\mathcal{H},C\mathcal{D})=F(\mathcal{H},\mathcal{D})$ for every $C>0$). If either $\mathbb{E}[|F|]<\infty$ or $F\geq 0$ a.s., then a.s.,
$\lim_{k\to\infty}\frac{1}{|S_{k}|^{2}}\int_{S_{k}}F(\mathcal{H}-z,\mathcal{D}-z)\,dz=\mathbb{E}[F].$ (2.3)
We comment that our version of Lemma 2.3 and Lemma 2.4 are slightly different from the ones in [GMS22] since there they only have the relation $\sim$ instead of our associated map $\mathcal{M}$ in the definition of cell configurations, but the same proof carries to our setting.
For $k>\ell\geq 1$, we write $\mathcal{S}_{k,\ell}$ for the collection of squares in $\mathcal{D}$ consisting of the dyadic ancestors $S$ of $S_{k}$ of side length at most $2^{\ell}|S_{k}|$, as well as each of dyadic descendant of such $S$ of side length at least $2^{-\ell}|S_{k}|$. In particular, the 8 squares in $\mathcal{D}$ with length $|S_{k}|$ and non-empty intersection with $S_{k}$ are in $\mathcal{S}_{k,\ell}$. The results below are stated in terms of $S_{k}$, but they are true for $\widehat{S}_{k}$ as well since each $\widehat{S}_{k}$ is equal to $S_{m(k)}$ for some $m(k)$ and $m(k)\to\infty$ as $k\to\infty$.
Lemma 2.5.
Let $E(S)=E(S,\mathcal{H})$ be an event depending on a square $S\subseteq\mathbb{C}$ and our cell configuration $\mathcal{H}$. Suppose $E$ is translation invariant, in the sense that for any $z\in\mathbb{C}$, $E(S,\mathcal{H})$ happens if and only if $E(S-z,\mathcal{H}-z)$ happens. Suppose that a.s. $E(S_{k})$ occurs for each large enough $k\in\mathbb{N}$. Then for each $\ell\in\mathbb{N}$, it is a.s. the case that for large enough $k\in\mathbb{N}$, $E(S)$ occurs for each square $S$ in $\mathcal{S}_{k,\ell}$.
Proof.
Consider the event $E_{1}(S)$ where $E(S^{\prime})$ occurs for each dyadic descendant $S^{\prime}$ of $S$ with side length at least $2^{-2\ell}|S|$. Then by [GMS22, Lemma 2.8], a.s. $E_{1}(S_{k})$ occurs for each large enough $k\in\mathbb{N}$. The claim is then straightforward by applying  [GMS22, Lemma 2.8] once more to the event $E_{1}(S)$. ∎
Lemma 2.6.
Almost surely, for each $\varepsilon\in(0,1)$ and $\ell>0$, it holds that for large enough $k\in\mathbb{N}$ that
$\mathrm{diam}(H)\leq\varepsilon|S|\ \ \text{for all }H\in\mathcal{H}(S)\text{ and $S\in\mathcal{S}_{k,\ell}$}.$ (2.4)
Proof.
By [GMS22, Lemma 2.9], where we set the conductance here to be equal to 1 and (1.3) implies the moment bound in [GMS22, Equation (1.5)], almost surely it holds that for large enough $k\in\mathbb{N}$ such that
$\mathrm{diam}(H)\leq\varepsilon|S_{k}|/2,\ \ \text{for all }H\in\mathcal{H}(S_{k}).$
The claim then follows from Lemma 2.5. ∎
Lemma 2.7.
Suppose for some $p\geq 0$, the cell configuration $\mathcal{H}$ satisfies the moment bound
$\mathbb{E}\Big[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\deg(H_{0})^{p}\Big]<\infty.$ (2.5)
There is a deterministic constant $C>0$ depending only on the law of $\mathcal{H}$ such that the following is true. For each $\ell\in\mathbb{N}$, it is a.s. the case that for large enough $k\in\mathbb{N}$ and each $S\in\mathcal{S}_{k,\ell}$,
$\frac{1}{|S|^{2}}\sum_{H\in\mathcal{H}(S)}\mathrm{diam}(H)^{2}\deg(H)^{p}<C.$ (2.6)
Moreover, almost surely, for each large enough $k$ and each square $S\subseteq[-2^{k},2^{k}]\times[-2^{k},2^{k}]$ of side length at least $2^{k-\ell}$,  (2.6) holds.
Proof.
Following the same argument as in [GMS22, Lemma 2.10], applying Lemma 2.4 with
$F(\mathcal{H},\mathcal{D})=\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\deg(H_{0})^{p}$
and Lemma 2.6, it follows that for some deterministic $C>0$, almost surely for each large enough $k\in\mathbb{N}$,
$\frac{1}{|S_{k}|^{2}}\sum_{H\in\mathcal{H}(S_{k})}\mathrm{diam}(H)^{2}\deg(H)^{p}<C/16.$ (2.7)
By Lemma 2.5, almost surely for large enough $k$, for each $S\in\mathcal{S}_{k,\ell+2}$,  (2.7) holds for $S$ instead of $S_{k}$. This verifies the first part of the claim. The second part of the claim follows since each square $S\subseteq[-2^{k},2^{k}]\times[-2^{k},2^{k}]$ of side length at least $2^{k-\ell}$ can be covered using squares in $\mathcal{S}_{k,\ell+2}$ with side length $2^{-\ell-2}|S_{k}|$ such that their total area is less than $16|S|$. ∎
3 Convergence of cell configurations under the circle packing
In this section we prove Theorem 1.7. Throughout this section, we assume that $\mathcal{H}$ is a random cell configuration that satisfies the constraints in Theorem 1.7. In Section 3.1, we recap the Ring Lemma from [RS87] and prove a variant (Lemma 3.3), which roughly states that in circle packings, a disk cannot simultaneously be tangent to two large disks unless one of the two large disks have high degree. In Section 3.2, we use Lemma 3.3 and the invariance principle from [GMS22] to prove that, the random walk on $\mathcal{H}$ with Dubejko weights converges to Brownian motion. In Section 3.3, we prove that the random walk on the circle packing $\mathcal{P}$ for $\mathcal{H}$ with Dubejko weights converges to Brownian motion. Both of the convergence are in quenched sense, which allow us to conclude that $\mathcal{H}$ is close to $\mathcal{P}$ in Section 3.4.
3.1 The Ring Lemma and a three-circle variant
One fundamental result in circle packing is the Ring Lemma [RS87]. Let $(D_{0};D_{1},...,D_{d})$ be disks with disjoint interiors. We say $(D_{0};D_{1},...,D_{d})$ form a flower with $d$ petals if $D_{1},...,D_{d}$ are all tangent to $D_{0}$ and separate $D_{0}$ from infinity, and for j=1,,dj=1,...,d, $D_{j}$ and $D_{j+1(\mathrm{mod}\ d)}$ are tangent. The following is the well known Ring Lemma by Rodin and Sullivan [RS87].
Lemma 3.1 (Ring Lemma).
Suppose $d\geq 3$ and $(D_{0};D_{1},...,D_{d})$ form a flower with $d$ petals. Let $r_{0},r_{1},...,r_{d}$ be the radii of $D_{0},D_{1},...,D_{d}$. Then there exists some absolute constant $c>0$ such that for each j=1,,dj=1,...,d, $r_{j}/r_{0}\geq\exp(-cd)$.
We will also use the following result known as Descartes’ theorem.
Lemma 3.2.
Let $D_{1},D_{2},D_{3}$ be mutually tangent disks with disjoint interiors and radii $r_{1},r_{2},r_{3}$. There exists exactly two circles that are tangent to all of $D_{1},D_{2},D_{3}$, and the radius $r$ of these two circles is determined by
$\frac{1}{r}=\frac{1}{r_{1}}+\frac{1}{r_{2}}+\frac{1}{r_{3}}\pm 2\sqrt{\frac{1}{r_{1}}+\frac{1}{r_{2}}+\frac{1}{r_{3}}}.$ (3.1)
If we take the plus (resp., minus) sign, then the circle is contained within the bounded (resp., unbounded) component of $\mathbb{C}\backslash(D_{1}\cup D_{2}\cup D_{3})$.
The bound in Lemma 3.1 is sharp. Let $D_{0}=\mathbb{D}$, $D_{1}=\{\mathrm{Im}\ z=1\}$, $D_{2}=\{\mathrm{Im}\ z=-1\}$ (where we view parallel lines as circles tangent at infinity), $D_{3}=B(2;1)$, and we recursively define $D_{n+2}$ by the disk which is (externally) tangent to both $D_{0},D_{n}$ and $D_{n+1}$. Then one can check by the Descartes’ theorem (see (3.1) above) that the ratio $r_{d}/r_{0}$ is equal to $\frac{1}{F_{2d-3}-1}$, where $F_{1}=1$, $F_{2}=1$, and $F_{n}$ is the $n$-th Fibonacci number. It is shown in [HAN88] that the bound described as above is optimal.
We will use the following variant of the Ring Lemma in Section 3.2, which states that in the flower $(D_{0};D_{1},...,D_{d})$, $D_{2}$ cannot have too small radii compared to both $D_{0}$ and $D_{1}$. This will eventually allow us to verify only polynomial bounds of degree (rather than exponential bounds) when applying (our variant of) the invariance principle in [GMS22].
Lemma $3.$
For any flower $(D_{0};D_{1},...,D_{d})$ with $d$ petals and radii $(r_{0};r_{1},...,r_{d})$,
$r_{2}\geq 0.01d^{-2}\min\{r_{0},r_{1}\}.$ (3.2)
Refer to caption Refer to caption
Figure 3: Left: A flower with 8 petals. Right: An illustration for the setup in the proof of Lemma 3.3 for $d=4$.
Proof.
Assume on the contrary that $r_{2}\leq 0.01d^{-2}\min\{r_{0},r_{1}\}$. We start with the case where $D_{2},...,D_{d}$ are all tangent to $D_{1}$. Suppose $D_{j}$ and $D_{0}$ are tangent at the point $a_{j}$, and $D_{1},...,D_{d}$ are ordered counterclockwise. See Figure 3 for an illustration. Using $r_{2}\leq 0.01d^{-2}\min\{r_{0},r_{1}\}$, one can easily check that the counterclockwise arc from $a_{1}$ to $a_{2}$ has length at most $\pi r_{0}/6$. Then for j=2,,dj=2,...,d, since $D_{j+1}$ is tangent to $D_{j},D_{0},D_{1}$ while $D_{j},D_{0},D_{1}$ are tangent to each other, and $D_{j+1}$ is not in the bounded component of $\mathbb{C}\backslash(D_{j+1}\cup D_{0}\cup D_{1})$, using Descartes’ Theorem, it follows by (3.1)
$\frac{1}{r_{j+1}}=\frac{1}{r_{j}}+\frac{1}{r_{0}}+\frac{1}{r_{1}}-2\sqrt{(\frac{1}{r_{0}}+\frac{1}{r_{1}})\frac{1}{r_{j}}+\frac{1}{r_{0}r_{1}}}.$ (3.3)
Let $j_{0}$ be the smallest $j\geq 2$ such that $r_{j}>\min\{r_{0},r_{1}\}/(100d)$. If no such $j_{0}$ exists, then we set $j_{0}=d$. Since $r_{2}\leq 0.01d^{-2}\min\{r_{0},r_{1}\}$, $3\leq j_{0}\leq d$. On the other hand, for $j=2,...,j_{0}-1$, from (3.3),
$\begin{split}\frac{1}{r_{j+1}}\geq\frac{1}{r_{j}}+\frac{1}{r_{0}}+\frac{1}{r_{1}}-2\sqrt{2(\frac{1}{r_{0}}+\frac{1}{r_{1}})\frac{1}{r_{j}}}\geq\frac{1}{r_{j}}+4\big(\frac{1}{r_{0}}+\frac{1}{r_{1}}\big)-4\sqrt{(\frac{1}{r_{0}}+\frac{1}{r_{1}})\frac{1}{r_{j}}}=\bigg(\sqrt{\frac{1}{r_{j}}}-2\sqrt{\frac{1}{r_{0}}+\frac{1}{r_{1}}}\bigg)^{2},\end{split}$ (3.4)
where we have used $r_{j}\leq\min\{r_{0},r_{1}\}/(100d)$. In particular,
$\sqrt{\frac{1}{r_{j+1}}}\geq\sqrt{\frac{1}{r_{j}}}-2\sqrt{\frac{1}{r_{0}}+\frac{1}{r_{1}}},\ \ \text{and }\sqrt{\frac{1}{r_{j_{0}}}}\geq\sqrt{\frac{1}{r_{2}}}-2d\sqrt{\frac{1}{r_{0}}+\frac{1}{r_{1}}}\geq 6d\min\{r_{0},r_{1}\}^{-1/2},$ (3.5)
If $j_{0}<d$, then this implies $r_{j_{0}}\leq\min\{r_{0},r_{1}\}/(36d^{2})$, which contradicts with the definition of $j_{0}$ since $d\geq 3$. Therefore $j_{0}=d$ and for each j=2,,dj=2,...,d, $r_{j}\leq\min\{r_{0},r_{1}\}/(36d^{2})$, and $\operatorname{dist}(a_{2},a_{d})\leq 2(r_{2}+...+r_{d})\leq r_{0}/(18d)$. In particular, the counterclockwise arc from $a_{2}$ to $a_{d}$ has length at most $\pi r_{0}/3$, and the counterclockwise arc from $a_{1}$ to $a_{d}$ is at most $\pi r_{0}/2$, $D_{1}\cup D_{2}\cup...\cup D_{d}$ cannot separate $D_{0}$ from infinity and $(D_{0};D_{1},...,D_{d})$ is not a flower.
For the general case, for j=1,,dj=1,...,d, suppose $D_{0}$ and $D_{j}$ is tangent at the point $a_{j}$. We perform a Möbius transform $\varphi$ on $\mathbb{C}$ sending $a_{1}$ to infinity. Then $\partial D_{0},\partial D_{1}$ are mapped to parallel straight lines $\ell_{0},\ell_{1}$, where the images $\varphi(D_{2}),...,\varphi(D_{d})$ are all disks tangent to $\ell_{0}$. Since $\varphi(D_{2}),...,\varphi(D_{d})$ all have at most one intersection point with $\ell_{1}$, the length of the line segment from $\varphi(a_{2})$ to $\varphi(a_{d})$ is maximized only when $\varphi(D_{2}),...,\varphi(D_{d})$ are all tangent to $\ell_{1}$ as well. By transforming back and comparing with the previous computation, we see that $D_{1}\cup D_{2}\cup...\cup D_{d}$ cannot separate $D_{0}$ from infinity. This is again a contradiction and we conclude the proof. ∎
We end this section with the following corollary of the Ring Lemma. For a simple infinite triangulation $\mathcal{T}$ with root vertex $v_{0}$, let $B_{m}^{\mathcal{T}}(v_{0})$ be the ball of radius $m$ centered at $v_{0}$ in the sense of graph distance. Let $\mathcal{P}$ be some circle packing of $\mathcal{T}$, where the disk corresponding to the root vertex $v_{0}$ is the unit disk. Also let $\mathcal{R}^{m}_{\mathcal{T}}$ and $\mathcal{X}^{m}_{\mathcal{T}}$ be the vectors of radii and centers, respectively, of the disks in $\mathcal{P}$ corresponding to the vertices in $B_{m}^{\mathcal{T}}(v_{0})$.
Proposition 3.4.
For simple infinite triangulations $\mathcal{T}$ and $(\mathcal{T}^{n})_{n\geq 1}$ with a common root vertex $v_{0}$, suppose that for any fixed $m\geq 1$, there exists $N\geq 1$ such that for any $n\geq N$, $B_{m}^{\mathcal{T}^{n}}(v_{0})$ agrees with $B_{m}^{\mathcal{T}}(v_{0})$ as planar maps rooted at $v_{0}$. Then for any fixed $k\geq 1$, one has the limit
$\lim_{n\to\infty}\mathcal{R}^{k}_{\mathcal{T}^{n}}=\mathcal{R}^{k}_{\mathcal{T}}.$
Proof.
Following the Ring Lemma, for any fixed $m$, $\{(\mathcal{R}^{m}_{\mathcal{T}^{n}},\mathcal{X}^{m}_{\mathcal{T}^{n}})\}_{n\geq 1}$ is a tight sequence. Assume on the contrary that there is a subsequence $\{s_{n}\}$ such that $\lim_{n\to\infty}\mathcal{R}^{k}_{\mathcal{T}^{s_{n}}}\neq\mathcal{R}^{k}_{\mathcal{T}}.$ Then using the argument of taking diagonal subsequences, one can find a further subsequence of $\{s_{n}\}$ (still denoted by $\{s_{n}\}$ in order to simplify notation) such that the limits of $\{(\mathcal{R}^{m}_{\mathcal{T}^{s_{n}}},\mathcal{X}^{m}_{\mathcal{T}^{s_{n}}})\}_{n\geq 1}$ exist for every $m\geq 1$. Moreover, since for every $m$, $B_{m}^{\mathcal{T}^{n}}(v_{0})$ agrees with $B_{m}^{\mathcal{T}}(v_{0})$ for large enough $n$, it is clear that this limit satisfies the circle packing conditions for $\mathcal{T}$. In other words, we have constructed a circle packing for $\mathcal{T}$ with the disk corresponding to $v_{0}$ being the unit disk and a different vector of radii for vertices in $B_{k}^{\mathcal{T}}(v_{0})$. This contradicts with the uniqueness of the circle packing for $\mathcal{T}$ as in Theorem 2.1. ∎
3.2 Convergence of random walk with Dubejko weights on cell configurations
Recall the random walk on circle packing and the Dubejko weights as defined in Section 2.1. In this section, we prove that, for cell configurations $\mathcal{H}$ in the context of Theorem 1.7, if we assign the conductance to be equal to the Dubejko weights, then the random walk on $\mathcal{H}$ with the Dubejko weights converges to Brownian motion modulo time parameterization.
We begin with the convergence of random walks on cell configurations in [GMS22]. Let $\beta_{1}:[0,T_{\beta_{1}}]\to\mathbb{C}$ and $\beta_{2}:[0,T_{\beta_{2}}]\to\mathbb{C}$ be continuous curves defined on possibly different time intervals. We define
$\mathbb{d}^{\rm CMP}(\beta_{1},\beta_{2}):=\inf_{\phi}\sup_{t\in[0,T_{\beta_{1}}]}|\beta_{1}(t)-\beta_{2}(\phi(t))|,$ (3.6)
where the infimum is over all increasing homeomorphisms $\phi:[0,T_{\beta_{1}}]\to[0,T_{\beta_{2}}]$ and “CMP” stands for “curves modulo parameterization”. For curves $\beta_{1}:[0,\infty)\to\mathbb{C}$ and $\beta_{2}:[0,\infty)\to\mathbb{C}$ defined for infinite time and $r>0$, we let $T_{1,r}$ (resp. $T_{2,r}$) be the first exit time of $\beta_{1}$ (resp. $\beta_{2}$) from the ball $B(0;r)$ (or 0 if the curve starts outside $B(0;r)$). We define
$\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{1},\beta_{2}):=\int_{1}^{\infty}e^{-r}\min\{1,\mathbb{d}^{\rm CMP}(\beta_{1}|_{[0,T_{1,r}]},\beta_{2}|_{[0,T_{2,r}]})\}\,dr,$ (3.7)
so that $\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{n},\beta)\to 0$ if and only if for Lebesgue a.e. $r>1$, $\beta_{n}$ stopped at its first exit time from $B(0;r)$ converges to $\beta$ stopped at its first exit time from $B(0;r)$ with respect to the metric (3.6).
Theorem 3.5 (Theorem 3.10 of [GMS22]).
Let $\mathcal{H}$ be a random cell configuration satisfying the conditions in Definition 1.3 and Definition 1.4. Further assume $\mathcal{H}$ satisfies the moment bound
$\mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi(H_{0})\big)\bigg]<\infty,\ \ \mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi^{*}(H_{0})\big)\bigg]<\infty$ (3.8)
where $\pi(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})$ and $\pi^{*}(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})^{-1}$. Then the simple random walk on $\mathcal{H}$ is a.s. recurrent. Moreover, for $z\in\mathbb{C}$, let $X^{z}$ denote the random walk on $\mathcal{H}$ started from $H_{z}$ with conductance $\mathfrak{c}$. For $j\in\mathbb{N}_{0}$, let $\widehat{X}_{j}^{z}$ be an arbitrarily chosen point of the cell $X_{j}^{z}$ and extend $\widehat{X}_{j}^{z}$ from $\mathbb{N}_{0}$ to $0,\infty)$ by piecewise linear interpolation. There is a deterministic covariance matrix $\Sigma$ with $\det\Sigma\neq 0$ such that the following is true. For each fixed compact set $A\subseteq\mathbb{C}$, it is a.s. the case that as $\varepsilon\to 0$, the maximum over all $z\in A$ of the Prokhorov distance between the conditional law of $\varepsilon\widehat{X}^{z/\varepsilon}$ given $\mathcal{H}$ and the law of Brownian motion started from $z$ with covariance matrix $\Sigma$, with respect to the topology on curves induced in (3.7), tends to 0.
Corollary 3.6.
[…]
Proof.
Write $\mathbf{R}_{\theta_{0}}$ for the matrix of rotating by angle $\theta_{0}$. One can directly check by computation that for a 2D Brownian motion $(X,Y)^{\mathrm{T}}$ with covariance matrix $\Sigma$, if $\mathbf{R}_{\theta_{0}}(X,Y)^{\mathrm{T}}$ is also a 2D Brownian motion with the same covariance matrix $\Sigma$, then $\Sigma$ must be a scalar times the identity matrix. This immediately implies Corollary 3.6. ∎
Corollary 3.7.
Let $\mathcal{H}$ be the cell configuration as in Theorem 1.7. Then $\mathcal{H}$ is a.s. circle packing parabolic.
Proof.
The condition (1.3) implies the condition (3.8) for cell configurations with unit conductance, and therefore by Theorem 3.5, the simple random walk on $\mathcal{H}$ is recurrent. Therefore combined with Theorem 2.1, we immediately infer that $\mathcal{H}$ is a.s. circle packing parabolic. ∎
The rest of the section is devoted to the proof of the following result.
Proposition 3.8.
[…]
Proof.
[…]
The only place in [GMS22] where the second equation in the condition (3.8) is used is [GMS22, Lemma 2.17]. We will first prove a slightly weaker version of this lemma (with condition (3.8) replaced by (1.3)) and then we check that the rest of the proof of [GMS22, Theorem 3.10] still carries through with the weaker version of the lemma. Throughout the proof, $c$ stands for some deterministic constant which may vary line by line.
Fix $k\in\mathbb{N}$, $r\in[0,1]$. Let $\mathcal{D}$ be the random dyadic square independent of $\mathcal{H}$ defined as in Definition 2.2, $S_{k}$ be the square in $\mathcal{D}$ containing 0 with side length in $2^{k-1},2^{k})$, and let $P_{k}(r)$ be the horizontal line segment connecting the left and right boundaries of the square $S_{k}$ whose distance from the lower boundary of $S_{k}$ is $r|S_{k}|$. We prove that for each large enough $k\in\mathbb{N}$ and each function $f:\mathcal{H}(S_{k})\to\mathbb{C}$,
$\int_{0}^{1}\sum_{\{H,H^{\prime}\}\in\mathcal{E}\mathcal{H}(P_{k}(r))}|f(H)-f(H^{\prime})|dr\leq c\Big(\sum_{\{H,H^{\prime}\}\in\mathcal{H}(\widetilde{S}_{k})}\mathfrak{c}(H,H^{\prime})|f(H)-f(H^{\prime})|^{2}\Big)^{1/2},$ (3.9)
where $\widetilde{S}_{k}$ is the union of the four dyadic parents of $S_{k}$. Note that this is identical to the statement of [GMS22, Lemma 2.17] except that we have $\widetilde{S}_{k}$ instead of $S_{k}$ on the right side.
Before proving (3.9) let us explain why this estimate is sufficient to conclude the proof of the theorem. The only place in the proof of [GMS22, Theorem 1.16] where [GMS22, Lemma 2.17] is used is in the proof of [GMS22, Equation (2.44)]. The latter equation is proven exactly as before.
For $\{H,H^{\prime}\}\in\mathcal{E}\mathcal{H}(S_{k})$, let $L_{k}(H,H^{\prime})$ be the Lebesgue measure of the set of $r\in[0,1]$ for which $H$ and $H^{\prime}$ both belong to $\mathcal{H}(P_{k}(r))$. Then, as argued in [GMS22, Lemma 2.17],
$\int_{0}^{1}\sum_{\{H,H^{\prime}\}\in\mathcal{E}\mathcal{H}(P_{k}(r))}|f(H)-f(H^{\prime})|dr=\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}L_{k}(H,H^{\prime})|f(H)-f(H^{\prime})|.$ (3.10)
and
$L_{k}(H,H^{\prime})\leq\frac{2}{|S_{k}|}\min\{\mathrm{diam}(H),\mathrm{diam}(H^{\prime})\}.$ (3.11)
By Lemma 2.7, there is a deterministic constant $c_{0}>0$ such that a.s. for large enough $k\in\mathbb{N}$,
$\frac{1}{|S_{k}|^{2}}\sum_{H\in\mathcal{H}(S_{k})}\mathrm{diam}(H)^{2}\deg(H)^{4}\leq c_{0}.$ (3.12)
Figure 4: Construction of the path $P_{H,H^{\prime}}$ in the proof of Proposition 3.8 when $r_{H^{\prime}}\leq 0.01d^{-1}r_{H}$. Here for j=2,,6j=2,...,6, $r_{H_{j}}<0.01d^{-1}r_{H}$ while $0.01d^{-1}r_{H}\leq r_{H_{1}}<r_{H}$.
Now for two cells $H\sim H^{\prime}$, we define as follows a set of edges $P_{H,H^{\prime}}$ forming a path from $H$ to $H^{\prime}$. We fix a circle packing $\mathcal{P}$ for $\mathcal{H}$. For $H\in\mathcal{H}$, we write $D_{H}$ for the corresponding circle in the packing and $r_{H}$ for the radius of $D_{H}$. We assume that $r_{H}\geq r_{H^{\prime}}$; the other case is dealt with by setting $P_{H,H^{\prime}}=P_{H^{\prime},H}$. We write $d=\deg(H)$. If $r_{H^{\prime}}\geq 0.01d^{-1}r_{H}$, then we define $P_{H,H^{\prime}}$ to be the single edge connecting $H$ and $H^{\prime}$. Then following Lemma 3.3, for any $H^{\prime\prime}$ adjacent to both $H$ and $H^{\prime}$, one must have $r_{H^{\prime\prime}}\geq 10^{-4}d^{-3}r_{H}$, and therefore by the Dubejko weight formula (2.1) $\mathfrak{c}(H,H^{\prime})\geq 10^{-4}d^{-2}$. If $r_{H^{\prime}}<0.01d^{-1}r_{H}$, we define the set $P_{H,H^{\prime}}$ as follows. First observe that if we draw straight lines connecting the centers of disks which are tangent to each other and $D_{H}$, then we form a closed curve around $D_{H}$ with length $2\sum_{H^{\prime\prime}\sim H}r_{H^{\prime\prime}}\geq 2\pi r_{H}$, and there must exist some $H^{\prime\prime}\sim H$ such that $r_{H^{\prime\prime}}\geq\pi r_{H}/d>0.01d^{-1}r_{H}$. Let $D_{H_{1}}$ be the first such disk when viewed clockwise from $D_{H^{\prime}}$, and we let $D_{H_{1}}$, $D_{H_{2}}$,…, $D_{H_{n}}$ be the disks tangent to $D_{H}$ and ordered counterclockwise where we write $H_{n}=H^{\prime}$. If $r_{H_{1}}\geq r_{H}$, define $P_{H,H^{\prime}}=\{\{H,H_{2}\},\{H_{2},H_{3}\},...,\{H_{n-1},H_{n}\}\}$; otherwise set $P_{H,H^{\prime}}=\{\{H,H_{1}\},\{H_{1},H_{2}\},...,\{H_{n-1},H_{n}\}\}$. See also Figure 4 for an illustration. Then by definition, $r_{H_{2}},...,r_{H_{n}}<0.01d^{-1}r_{H}$ and $r_{H_{1}}\geq 0.01d^{-1}r_{H}$. By Lemma 3.3, $r_{H_{2}}>0.01d^{-2}\min\{r_{H},r_{H_{1}}\}$. If $r_{H_{1}}\geq r_{H}$, then $r_{H_{2}}\geq 0.01d^{-2}r_{H}$, and by the Dubejko weight formula (2.1), $\mathfrak{c}(H,H_{2})\geq 10^{-4}d^{-1}$. If $r_{H_{1}}<r_{H}$, then using $r_{H_{1}}\geq 0.01d^{-1}r_{H}$, we have $\mathfrak{c}(H,H_{1})\geq 10^{-4}d^{-2}$. In this case, $0.01d^{-2}r_{H_{1}}\leq r_{H_{2}}\leq r_{H_{1}}<r_{H}$, and $\mathfrak{c}(H_{1},H_{2})\geq 10^{-4}d^{-1}$. For j=2,,n1j=2,...,n-1, $r_{H_{j}}<r_{H}$, and by Lemma 3.3, we have $0.01d^{-2}<r_{H_{j+1}}/r_{H_{j}}<100d^{2}$. Then again using the formula (2.1), one has $\mathfrak{c}(H_{j},H_{j+1})>10^{-4}d^{-1}$. In other words, for each edge $e\in P_{H,H^{\prime}}$, if $H$ is an endpoint of $e$, then $\mathfrak{c}(e)\geq 10^{-4}d^{-2}$, and if $H$ is not an endpoint of $e$, then $\mathfrak{c}(e)\geq 10^{-4}d^{-1}$. In particular, for any $\{H,H^{\prime}\}\in\mathcal{E}\mathcal{H}$ with $r_{H}\geq r_{H^{\prime}}$, using $n\leq\deg(H)=d$, we have
$\sum_{e\in P_{H,H^{\prime}}}\mathfrak{c}(e)^{-1}\leq 2\cdot 10^{4}\deg(H)^{2}.$ (3.13)
For an edge $e=\{H,H^{\prime}\}$, we set $|\nabla f(e)|=|f(H^{\prime})-f(H)|$. Then from the triangle inequality, $|f(H)-f(H^{\prime})|\leq\sum_{e\in P_{H,H^{\prime}}}|\nabla f(e)|$. Now by (3.10) and (3.11),
$\begin{split}\int_{0}^{1}&\sum_{\{H,H^{\prime}\}\in\mathcal{E}\mathcal{H}(P_{k}(r))}|f(H)-f(H^{\prime})|dr\leq\frac{2}{|S_{k}|}\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\min\{\mathrm{diam}(H),\mathrm{diam}(H^{\prime})\}|f(H)-f(H^{\prime})|\\ &\leq\frac{4}{|S_{k}|}\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\sum_{e\in P_{H,H^{\prime}}}\mathds{1}_{r_{H}\geq r_{H^{\prime}}}\min\{\mathrm{diam}(H),\mathrm{diam}(H^{\prime})\}|\nabla f(e)|\\ &\leq\frac{4}{|S_{k}|}\Big(\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\sum_{e\in P_{H,H^{\prime}}}\mathds{1}_{r_{H}\geq r_{H^{\prime}}}|\nabla f(e)|^{2}\mathfrak{c}(e)\frac{1}{\deg(H)}\Big)^{1/2}\\ &\times\Big(\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\sum_{e\in P_{H,H^{\prime}}}\mathds{1}_{r_{H}\geq r_{H^{\prime}}}\mathrm{diam}(H)^{2}\deg(H)\mathfrak{c}(e)^{-1}\Big)^{1/2}.\end{split}$ (3.14)
We will now conclude the proof of (3.9) by bounding the right side of (3.14), starting with the first factor. For each edge $e$, there exist at most two different $H$ such that $e\in P_{H,H^{\prime}}$ for some $H^{\prime}$ with $r_{H}\geq r_{H^{\prime}}$ and $H$ is not an endpoint of; indeed, by our definition of $P_{H,H^{\prime}}$ the disks of two end-points of $e$ and $D_{H}$ must all pairwise touch. There are also at most two $H$ such that $e\in P_{H,H^{\prime}}$ for some $H^{\prime}$ with $r_{H}\geq r_{H^{\prime}}$ and $H$ is an endpoint of $e$. In either case, the corresponding number of $H^{\prime}$ is at most $\deg(H)$. Moreover, it follows from Lemma 2.6 that for large enough $k$, if $\{H,H^{\prime}\}\in\mathcal{H}(S_{k})$ and $H^{\prime\prime}$ is an end-point of an edge in $P_{H,H^{\prime}}$ then $H^{\prime\prime}\in\mathcal{H}(\widetilde{S}_{k})$. Therefore
$\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\sum_{e\in P_{H,H^{\prime}}}\mathds{1}_{R_{H}\geq R_{H^{\prime}}}|\nabla f(e)|^{2}\mathfrak{c}(e)\frac{1}{\deg(H)}\leq 4\sum_{e\in\mathcal{E}\mathcal{H}(\widetilde{S}_{k})}|\nabla f(e)|^{2}\mathfrak{c}(e).$ (3.15)
On the other hand, by (3.13), we have
$\begin{split}\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\sum_{e\in P_{H,H^{\prime}}}&\mathds{1}_{R_{H}\geq R_{H^{\prime}}}\mathrm{diam}(H)^{2}\deg(H)\mathfrak{c}(e)^{-1}\\ &\leq 2\cdot 10^{4}\sum_{\{H,H^{\prime}\}\in\mathcal{H}(S_{k})}\mathrm{diam}(H)^{2}\deg(H)^{3}\leq 2\cdot 10^{4}\sum_{H\in\mathcal{H}(S_{k})}\mathrm{diam}(H)^{2}\deg(H)^{4}.\end{split}$ (3.16)
Thus (3.9) follows by plugging (3.12), (3.15), (3.16) into (3.14). ∎
3.3 Convergence of random walk with Dubejko weights on circle packing
Consider a circle packing $\mathcal{P}$ for $\mathcal{H}$ such that the disk for the cell $H_{0}$ is the unit disk, where we recall that $\mathcal{H}$ is circle packing parabolic by Corollary 3.7. For $z\in\mathbb{C}$, consider the random walk $(\widehat{X}_{j}^{z,\rm circle})_{j\in\mathbb{N}_{0}}$ moving along centers of disks in $\mathcal{P}$ with Dubejko weights, starting from the disk with closest distance to $z$. We extend $\widehat{X}_{j}^{z,\rm circle}$ from $\mathbb{N}_{0}$ to $0,\infty)$ by piecewise linear interpolation. The aim of this section is to prove the following.
Proposition 3.9.
For each compact set $A\subseteq\mathbb{C}$, it is a.s. the case that as $\varepsilon\to 0$, the maximum over all $z\in A$ of the Prokhorov distance between the conditional law of $\varepsilon\widehat{X}^{z/\varepsilon,\mathrm{circle}}$ given $\mathcal{P}$ and the law of standard Brownian motion started from $z$, with respect to the topology on curves induced in (3.7), tends to 0.
One critical input for the proof of Proposition 3.9 is the following, which is an immediate consequence of [BG24, Theorem A].
Proposition 3.10 ([BG24]).
Let $(\mathcal{P}^{n})_{n\geq 1}$ be a sequence of circle packings for simple infinite triangulations $(\mathcal{T}^{n})_{n\geq 1}$. Suppose for each $r>0$, the maximal diameter of disks in $\mathcal{P}^{n}$ intersecting $B(0;r)$ converges to 0. Let $\widehat{X}^{z,n,\rm circle}$ be the linearly interpolated random walk on $\mathcal{P}^{n}$ defined as above. Then for each compact set $A\subseteq\mathbb{C}$, it is a.s. the case that as $n\to\infty$, the maximum over all $z\in A$ of the Prokhorov distance between the law of $\widehat{X}^{z,n,\mathrm{circle}}$ and the law of standard Brownian motion started from $z$, with respect to the topology on curves induced in (3.7), tends to 0.
Proposition 3.9 is now immediate from Proposition 3.10 and the following.
Proposition 3.11.
Let $r>0$ be fixed. Then the maximal diameter of the circles in $\varepsilon\mathcal{P}$ intersecting $B(0;r)$ converges to 0 almost surely as $\varepsilon\to 0$.
The rest of this subsection is devoted to the proof of Proposition 3.11. In Section 3.3.1, we prove Lemma 3.12, which relate the maximal diameter of the disks in circle packings with the vertex extremal length of graphs as in [GJN19]. In Section 3.3.2 we prove Proposition 3.11 using Lemma 3.12 and Lemma 2.7. We refer to Remark 4.11 for an alternative proof strategy for Proposition 3.11.
3.3.1 Vertex extremal length and macroscopic circles
Consider a planar map $\mathcal{M}$ with vertex set $\mathcal{V}\mathcal{M}$. Given a function $m:\mathcal{VG}\to[0,\infty)$ and a path $\ell$, define the length $\mathrm{len}_{m}(\ell)=\sum_{v\in\ell}m(v)$. For a family $\Gamma$ of paths, define $\mathrm{len}_{m}(\Gamma)=\inf_{\ell\in\Gamma}\mathrm{len}_{m}(\ell)$. The vertex extremal length of $\Gamma$, as introduced by Cannon [CAN94], is defined as
$\mathrm{VEL}(\Gamma)=\sup_{m:\mathcal{VG}\to[0,\infty)}\frac{\mathrm{len}_{m}(\Gamma)^{2}}{\mathrm{area}(m)},$
where $\mathrm{area}(m)=\sum_{v\in\mathcal{V}\mathcal{M}}m(v)^{2}$.
As studied in [GJN19], vertex extremal length has close relation with macroscopic circles in the circle packing. We will need the following result.
Lemma 3.12.
Consider a simple infinite triangulation $\mathcal{T}$. Let $v_{0}$ be a vertex. Consider a circle packing $\mathcal{P}$ of $\mathcal{T}$ with the circle representing the vertex $v_{0}$ centered at origin. For any vertex $v$ different from $v_{0}$, let $\Gamma_{v,v_{0}}$ be the collection of finite paths of $\mathcal{T}$ starting and ending at $v$ which makes a simple loop around $v_{0}$. Let $D_{v}$ be the circle representing $v$, $r_{v}$ be the radius of $D_{v}$, and $z_{v}$ be the center of $D_{v}$. Then
$\mathrm{VEL}(\Gamma_{v,v_{0}})\leq 4|z_{v}|r_{v}^{-1}.$
Proof.
The proof is identical to that of [GJN19, Theorem 1.1, equation (1)], which proves a similar result for circle packings in a disk. We include the proof here for completeness. For $|z_{v}|-r_{v}\leq t\leq|z_{v}|+r_{v}$, let $\gamma_{t}$ be the loop in $\mathcal{T}$ formed by the vertices whose disk in $\mathcal{P}$ intersects $\{|z|=t\}$. Indeed $\gamma_{t}\in\Gamma_{v,v_{0}}$. Now for any $m:\mathcal{VG}\to[0,\infty)$,
$\begin{split}\int_{|z_{v}|-r_{v}}^{|z_{v}|+r_{v}}\mathrm{len}_{m}(\gamma_{t})dt&=\sum_{w\in\mathcal{V}\mathcal{T}}\int_{|z_{v}|-r_{v}}^{|z_{v}|+r_{v}}\mathds{1}_{w\in\gamma_{t}}m(w)\,dt=\sum_{w\in\mathcal{V}\mathcal{T}}m(w)\int_{|z_{v}|-r_{v}}^{|z_{v}|+r_{v}}\mathds{1}_{w\in\gamma_{t}}\,dt.\end{split}$ (3.17)
For each $w\in\mathcal{V}\mathcal{T}$, let $z_{w}$ and $r_{w}$ be the center and radius of $D_{w}$, and $\theta_{w}=\arg z_{w}$. Let $\underline{a_{w}}=\max\{|z_{w}|-r_{w},|z_{v}|-r_{v}\}$, and $\overline{a}_{w}=\min\{|z_{w}|+r_{w},|z_{v}|+r_{v}\}$. Then
$\int_{|z_{v}|-r_{v}}^{|z_{v}|+r_{v}}\mathds{1}_{w\in\gamma_{t}}\,dt=\max\{\overline{a}_{w}-\underline{a_{w}},0\}.$ (3.18)
For each $w$ with $\overline{a}_{w}-\underline{a_{2}}>0$, we draw the disk $D_{w}^{\prime}$ centered at $(\overline{a}_{w}+\underline{a_{w}})e^{i\theta_{w}}/2$ with radius $(\overline{a}_{w}-\underline{a_{w}})/2$. Then the disks $D_{w}^{\prime}$ are contained within $\{|z_{v}|-r_{v}\leq|z|\leq|z_{v}|+r_{v}\}\cap D_{w}$ and hence have disjoint interiors. See also [GJN19, Figure 3] for an illustration. In particular,
$\sum_{w\in\mathcal{V}\mathcal{T}}\pi\big(\max\{\overline{a}_{w}-\underline{a_{w}},0\}\big)^{2}/4\leq\pi\big((|z_{v}|+r_{v})^{2}-(|z_{v}|-r_{v})^{2}\big)=4\pi|z_{v}|r_{v}.$ (3.19)
Now combining (3.17),  (3.18) and (3.19), by Cauchy-Schwarz inequality,
$\begin{split}\int_{|z_{v}|-r_{v}}^{|z_{v}|+r_{v}}\mathrm{len}_{m}(\gamma_{t})dt\leq\big(\sum_{w\in\mathcal{V}\mathcal{T}}\big(\max\{\overline{a}_{w}-\underline{a_{w}},0\}\big)^{2}\big)^{1/2}\big(\sum_{w\in\mathcal{V}\mathcal{T}}m(w)^{2})^{1/2}\leq(16|z_{v}|r_{v})^{1/2}\big(\sum_{w\in\mathcal{V}\mathcal{T}}m(w)^{2})^{1/2}.\end{split}$ (3.20)
In particular, there exists $t\in(z_{v}-r_{v},z_{v}+r_{v})$, such that $\mathrm{len}_{m}(\gamma_{t})\leq 2|z_{v}|^{1/2}r_{v}^{-1/2}\big(\sum_{w\in\mathcal{V}\mathcal{T}}m(w)^{2})^{1/2}$ and $\mathrm{len}_{m}(\gamma_{t})^{2}/\mathrm{area}(m)\leq 4|z_{v}|r_{v}^{-1}$. Since $m$ is arbitrary, this verifies the claim. ∎
3.3.2 Proof of Proposition 3.11
Recall the notion of vertex extremal length introduced in Section 3.3.1. Let $\widehat{\mathcal{H}}$ be a planar embedding of the associated map of $\mathcal{H}$ that is close to $\mathcal{H}$ in the sense of Definition 1.5, with vertex set $\{z_{H}:H\in\mathcal{H}\}$. For each $H\in\mathcal{H}$, let $\Gamma_{H}$ be the collection of finite paths on $\widehat{\mathcal{H}}$ starting from $z_{H}$ and making a simple loop around $z_{H_{0}}$.
Proposition 3.13.
Let $q_{0}>0$. Then almost surely, there exists at most finitely many $H\in\mathcal{H}$, such that $\mathrm{VEL}(\Gamma_{H})<q_{0}$.
Proof.
Without loss of generality assume $q_{0}>300$ and set $q=q_{0}/72$. We will adapt the proof of [GJN19, Theorem 1.4] together with the input from Section $2.$ For $c>0$ and any square $S$, define the event $E_{1}(c,S)$ where
$\sum_{H\in\mathcal{H}(S)}\mathrm{diam}(H)^{2}\leq c|S|^{2}.$ (3.21)
Let $a\in(0,1)$, and set $E_{k,1}(c,a)$ be the event where $E_{1}(c,S_{k})$ holds for each square $S$ in $\mathcal{D}$ with side length in $(a|S_{k}|,|S_{k}|]$ which intersects $3|S_{k}|,3|S_{k}|]^{2}$. For $\varepsilon\in(0,1)$ and any square $S$, define the event $E_{2,k}(\varepsilon):=\{\mathrm{diam}(H)\leq\varepsilon|S_{k}|,\ \forall H\in\mathcal{H}([-3|S_{k}|,3|S_{k}|]^{2})\}$. Then by Lemmas 2.5,  2.6 and 2.7, for some deterministic constant $c_{0}>0$ and every $a,\varepsilon\in(0,1)$, almost surely, $E_{k,1}(c_{0},a)\cap E_{k,2}(\varepsilon)$ happens for all large enough $k$.
Let $n$ be the smallest integer greater than $c_{0}q$. We set $a=2^{-8n}$ and $\varepsilon=2^{-32n}$. We assume $k$ is large enough and work on the event $E_{k,1}(c_{0},a)\cap E_{k,2}(\varepsilon)$. Consider a cell $H^{*}$ intersecting $2|S_{k}|,2|S_{k}|]^{2}\backslash[-|S_{k}|,|S_{k}|]^{2}$. Let $S^{*}_{0}$ be a square in $\mathcal{D}$ with side length $2^{-8n}|S_{k}|$ intersecting $H^{*}$. For $j=1,2,$..., we recursively define $S_{j}^{*}$ to be the square concentric to $S_{j-1}^{*}$, and having side length $3|S_{j-1}^{*}|$. Then it is clear that $|S_{n}^{*}|<2^{-n}|S_{k}|$ and $\operatorname{dist}(0,S_{n}^{*})\geq|S_{k}|/2$. We define a mass function $m^{*}$ as follows. For each j=2,,nj=2,...,n and $H\subseteq S_{j}^{*}\backslash S_{j-1}^{*}$, we set $m^{*}(H)=|S_{j}^{*}|^{-1}\mathrm{diam}(H)$, and we set $m^{*}(H)=0$ for all other $H$. Then for each $j$, $S_{j}^{*}$ is contained within the union of 4 squares in $\mathcal{D}$ with side length at most $2|S_{j}^{*}|$, and therefore by the definition of the event $E_{k,1}(c_{0},a)$, we have
$\sum_{H\subseteq S_{j}^{*}\backslash S_{j-1}^{*}}m^{*}(H)^{2}=|S_{j}^{*}|^{-2}\sum_{H\subseteq S_{j}^{*}\backslash S_{j-1}^{*}}\mathrm{diam}(H)^{2}\leq 16c_{0},$ (3.22)
and therefore $\mathrm{area}({m^{*}})\leq 16c_{0}(n-1)$. On the other hand, using the definition of $E_{k,2}(\varepsilon)$, $H^{*}\subseteq S_{1}^{*}$ and $H_{0}\cap S_{n}^{*}=\emptyset$. For each path $\ell\in\Gamma_{H^{*}}$, using $H\cap H^{\prime}\neq\emptyset$ for $H\sim H^{\prime}$ for all large enough $k$, $\cup_{H\in\ell}H$ must cross all the regions $S_{j}^{*}\backslash S_{j-1}^{*}$, which implies
$\sum_{H\subseteq\ell\cap\mathcal{H}({S_{j}^{*}\backslash S_{j-1}^{*}})}m^{*}(H)=|S_{j}^{*}|^{-1}\sum_{H\subseteq\ell\cap\mathcal{H}({S_{j}^{*}\backslash S_{j-1}^{*}})}\mathrm{diam}(H)\geq|S_{j}^{*}|^{-1}(\frac{1}{3}|S_{j}^{*}|-2\cdot 2^{-32n}|S_{k}|)\geq\frac{1}{6},$ (3.23)
and $\mathrm{len}_{m^{*}}(\ell)\geq\frac{1}{6}(n-1)$. Therefore by (3.22) and  (3.23),
$\mathrm{VEL}(\Gamma_{H^{*}})\geq\frac{(n-1)}{36c_{0}}\geq\frac{q}{72}=q_{0}.$
Since $H^{*}$ is any arbitrary cell intersecting $2|S_{k}|,2|S_{k}|]^{2}\backslash[-|S_{k}|,|S_{k}|]^{2}$ and almost surely $E_{k,1}(c_{0},a)\cap E_{k,2}(\varepsilon)$ occurs for all large enough $k$, we conclude our proof. ∎
Proof of Proposition 3.11.
Without loss of generality assume $r=1$. Assume on the contrary that there exists $\delta_{0}>0$ such that, with probability at least $\delta_{0}$, there exists $\varepsilon_{n}\to 0$ and disks $D_{H_{n}}$ in the packing $\varepsilon_{n}\mathcal{P}$ intersecting $\mathbb{D}$ with diameter at least $\delta_{0}$. Let $z_{H_{n}}$ and $r_{H_{n}}>\delta_{0}$ be the center and radius, respectively, of $D_{H_{n}}$. Since $D_{H_{n}}$ intersects $\mathbb{D}$, $|z_{H_{n}}|-r_{H_{n}}<1$, and by Lemma 3.12,
$\mathrm{VEL}(\Gamma_{H_{n}})\leq 4|z_{H_{n}}|r_{H_{n}}^{-1}\leq 4(r_{H_{n}}+1)r_{H_{n}}^{-1}\leq 4(1+\delta_{0}^{-1}).$
Furthermore, since for any given disk in $\mathcal{P}$, its diameter converges to 0 in the packing $\varepsilon\mathcal{P}$ as $\varepsilon\to 0$, we see that the vertices in $\{v_{1},v_{2},$...\} are infinite. In other words, with probability at least $\delta_{0}$, there exists infinitely many cells $\{H_{n}\}$ in $\mathcal{H}$ such that $\mathrm{VEL}(\Gamma_{H_{n}})\leq 4(1+\delta_{0}^{-1})$. This contradicts with Proposition 3.13. ∎
3.4 Proof of Theorem 1.7
Our proof of Theorem 1.7 is mostly based on the convergence of the same random walk on planar maps with different embeddings to the Brownian motion as in Proposition 3.8 and Proposition 3.9, as well as the consideration of wrapping around events for 2D Brownian motion. In the rest of this section, we use the following notation. For a continuous-time walk $(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a bounded set $A\subseteq\mathbb{C}$ where $z\notin A$, we write $\mathbb{P}^{z}$ for the law of $(X_{t})_{t\geq 0}$ started from $z$, and $\tau_{A}$ for the first time $T$ when the trace of $(X_{t})_{0\leq t\leq T}$ separates AA from infinity. We write $\tau_{w}$ instead of $\tau_{\{w\}}$ if $A=\{w\}$.
Proposition 3.14.
Let $f:\mathbb{C}\to\mathbb{C}$ be a homeomorphism and $(W_{t})_{t\geq 0}$ be the standard Brownian motion in $\mathbb{C}$ starting from 0. Suppose $f(0)=0,f(1)=1$, and for every $z\in\mathbb{C}$, $\big(f(z+W_{t})\big)_{t\geq 0}$ has the law of a standard Brownian motion starting from $f(z)$ modulo time change, then either $f$ is the identity function, or $f(z)=\overline{z}$ for every $z\in\mathbb{C}$.
Proof.
Step 1. $f$ maps the line $\{\mathrm{R}e\,z=\frac{1}{2}\}$ to itself. By symmetry, we have $\mathbb{P}^{z}[\tau_{0}<\tau_{1}]=\mathbb{P}^{z}[\tau_{0}>\tau_{1}]$ for $\{\mathrm{R}e\,z=\frac{1}{2}\}$. Moreover, if $\mathrm{Re}\,z_{0}<\frac{1}{2}$, then with positive probability the standard Brownian motion started from $z_{0}$ separates 0 from infinity before hitting $\{\mathrm{R}e\,z=\frac{1}{2}\}$, and by strong Markov property, conditioned on the previous event not to happen, $\{\tau_{0}<\tau_{1}\}$ and $\{\tau_{0}>\tau_{1}\}$ have equal probability, which implies that $\mathbb{P}^{z}[\tau_{0}<\tau_{1}]>\mathbb{P}^{z}[\tau_{0}>\tau_{1}]$ (resp. $\mathbb{P}^{z}[\tau_{0}<\tau_{1}]<\mathbb{P}^{z}[\tau_{0}>\tau_{1}]$) if the starting point $z$ is on the left (resp. right) side of $\{\mathrm{R}e\,z=\frac{1}{2}\}$. Thus $f$ maps $\{\mathrm{R}e\,z=\frac{1}{2}\}$ to itself.
Step 2. For every $x>0$, $|f(\frac{1}{2}+xi)|=|f(\frac{1}{2}-xi)|$. This follows from the same argument in the previous case by looking at the stopping times $\tau_{\frac{1}{2}+xi}$ and $\tau_{\frac{1}{2}-xi}$ for the Brownian motion started from 0.
Step 3. For $x\in\mathbb{R}$, $f(\frac{1}{2}+xi)=\frac{1}{2}+xi$, or for all $x\in\mathbb{R}$, $f(\frac{1}{2}+xi)=\frac{1}{2}-xi$. For $x,y\in\mathbb{R}$, we work on the event $E_{x,y}^{\prime}:=\{\tau_{[\frac{1}{2}+xi,\frac{1}{2}+yi]}\leq\tau_{1}\}$, where $\frac{1}{2}+xi,\frac{1}{2}+yi]$ is the line segment joining $\frac{1}{2}+xi$ with $\frac{1}{2}+yi$. Suppose for $x\in\mathbb{R}$, $f(\frac{1}{2}+xi)=\frac{1}{2}+g(x)i$. Then $g$ sends the line segment between $\frac{1}{2}+xi$ and $\frac{1}{2}+yi$ to the line segment between $\frac{1}{2}+g(x)i$ and $\frac{1}{2}+g(y)i$. By using this, $f(0)=0$, $f(1)=1$, and that $f$ sends a Brownian motion to a Brownian motion modulo time change, we have $\mathbb{P}^{0}[E^{\prime}_{x,y}]=\mathbb{P}^{0}[E^{\prime}_{g(x),g(y)}]$. Moreover, $\mathbb{P}^{0}[E^{\prime}_{-x,x}]$ is strictly decreasing as $x>0$. By comparing the $\mathbb{P}^{0}[E^{\prime}_{-x,x}]$ for $(W_{t})_{t\geq 0}$ with $\mathbb{P}^{0}[E^{\prime}_{-g(x),g(x)}]$ for $(f(W_{t}))_{t\geq 0}$, we conclude that $f(\frac{1}{2}+xi)=\frac{1}{2}\pm xi$. Furthermore, if there exists $x,y>0$ such that $f(x)=\frac{1}{2}+ix$ and $f(y)=\frac{1}{2}-iy$, then one would have $\mathbb{P}^{0}[E^{\prime}_{x,y}]<\mathbb{P}^{0}[E^{\prime}_{g(x),g(y)}]$, which is a contradiction. Using a similar argument for other cases we see that the conclusion holds.
Step 4. Conclusion. Without loss of generality, assume that $f(\frac{1}{2}+xi)=\frac{1}{2}+xi$ for all $x\in\mathbb{R}$. Applying the same argument in the previous steps with i,$i$,-i replacing $0,1$, we see that $f(x)=x$ when $x\in\mathbb{R}$ or f(x)=xf(x)=-x for all $x\in\mathbb{R}$. Since $f(1)=1$, we must have $f(x)=x$ for every $x\in\mathbb{R}$. By another application of the argument in Step 1, where we consider the wrapping around $\frac{1}{2}+(y-1)i$ and $\frac{1}{2}+(y+1)i$ for Brownian motions started on the line $\{{\mathrm{Im}}\,z=y\}$, we see that $f(x+iy)$ must be on the line $\{{\rm{Im}}\,z=y\}$. Similarly $f(x+iy)$ must be on the line $\{{\mathrm{Re}}\,z=x\}$. Thus $f(x+iy)=x+iy$ for each $x,y\in\mathbb{R}$ and we conclude the proof. ∎
Proof of Theorem 1.7.
[…]
Let $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ be the linearly interpolated random walk on $\widetilde{\varepsilon}\mathcal{P}$ described in Proposition 3.9, and let $\widehat{X}^{z,\varepsilon}$ be interpolated random walk on $\varepsilon\mathbf{A}\mathcal{H}$ as in Theorem 3.5 using the points $\{\mathbf{A}z_{H}:H\in\mathcal{H}\}$, except that we interpolate along the edges $\{\mathbf{A}\gamma_{H,H^{\prime}}:H\sim H^{\prime}\}$. Note that $f^{\varepsilon}(\widehat{X}^{z,\varepsilon})$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ can be coupled to have the same trace. Indeed, this coupling exists since $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ can be viewed as two embeddings of the associated map for $\mathcal{H}$ and have identical conductance, and we work on this coupling. Then by Proposition 3.8 and Proposition 3.9, as $\varepsilon\to 0$ the conditional law of $\widehat{X}^{z,\varepsilon}$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ given $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ converges to a constant multiple of planar Brownian motion with respect to the topology on curves induced in (3.7). Throughout the proof, we write $\mathbb{P}_{\rm BM}$ for the law of standard Brownian motion, $\mathbb{P}_{\varepsilon}$ for the conditional law of $\widehat{X}^{\varepsilon}$ given $\varepsilon\mathcal{H}$, and $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}$ for the conditional law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ given $\widetilde{\varepsilon}\mathcal{P}$. Also let $K$ be a fixed compact subset of $\mathbb{C}$ containing 0 and 1, and $d_{K}=\mathrm{diam}(K)$. We will verify that $\{f^{\varepsilon}\}$ is uniformly bounded and equicontinuous on compacts.
Uniform boundedness on compact sets. For $z\neq 0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity. Let $\gamma_{K}$ be a deterministic curve formed by the union of the line segment from 0 to $(2d_{K}+3)i$ and a curve from $(2d_{K}+3)i$ to $2+(d_{K}+2)i$, such that it is disjoint from $B(0;d_{K}+1)$ and separates $B(0;{d_{K}+1})$ from infinity. Then for any $\delta>0$, with positive probability, the planar Brownian motion started from 0 stays within the $\delta/2$-neighborhood of $\gamma_{K}$ and form a loop around $K\cup\{1\}\subset B(0;d_{K}+1)$ during time $[0,2]$. In particular, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}\leq\tau_{1}]$ is uniformly positive for $z\in K$ and small enough $\varepsilon$. Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, for each $z\in K$, there exists $\varepsilon^{\prime}>0$ such that $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}]>3\varepsilon^{\prime}$ for $\varepsilon$ sufficiently small. On the other hand, for each $\varepsilon^{\prime}>0$, there exists $M,\delta_{1}>0$, such that the probability that a Brownian motion from 0 separates $B(1;\delta_{1})$ from infinity before leaving $B(0;M)$ is at least $1-\varepsilon^{\prime}$. Using the convergence in law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ to standard Brownian motion, for $\varepsilon$ sufficiently small, with probability at least $1-2\varepsilon^{\prime}$, the trace of $\widehat{X}^{0,\rm circle,\widetilde{\varepsilon}}$ separates 1 from infinity before leaving $B(0;2M)$. In other words, for $\varepsilon$ sufficiently small, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}\leq\tau_{1}]\leq 2\varepsilon^{\prime}$ for $|z|>2M$. Thus $|f^{\varepsilon}(z)|\leq 2M$ for all $z\in K$ and small enough $\varepsilon$, which gives the uniform boundedness for $f^{\varepsilon}$ on compacts.
Equicontinuity near 0 and 1. Let $\delta>0$. For $z\neq 0,1$, we work on the event $\{\tau_{z}=\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity at the same time. Using the same $\gamma_{K}$ and Brownian motion argument as in the previous step, one can see that there exists $\delta^{\prime}>0$ such that for $\varepsilon$ small enough $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}=\tau_{1}]>\delta^{\prime}$ when $z\in K$ and $|z|>\delta$. On the other hand, there exists $\delta^{\prime\prime}>0$ such that the probability that a planar Brownian motion separates $B(0;2\delta^{\prime\prime})$ from infinity before leaving $B(0;1/2)$ is at least $1-\delta^{\prime}/2$, and thus using the convergence in law of $\widehat{X}^{\varepsilon}$ to Brownian motion in local uniform topology, for $|z|<\delta^{\prime\prime}$ and all sufficiently small $\varepsilon$, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}=\tau_{1}]<\delta^{\prime}$. Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, we conclude that there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)|<\delta$. This verifies the equicontinuity near 0. The equicontinuity near 1 can be verified similarly by considering walks started at 1.
Equicontinuity at other points in $K$. Fix $z\in K$ with $z\neq 0,1$. For $z^{\prime}\neq z,0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}\}$, i.e., when the path separates 1 from infinity, it separates $z$ from infinity but does not separate $z^{\prime}$ from infinity. From the previous cases, $\{f^{\varepsilon}(z)\}$ is uniformly bounded away from 0, 1 and $\infty$. Fix $0<\delta<\frac{1}{2}\min_{\varepsilon>0}\{1,|f^{\varepsilon}(z)|,|f^{\varepsilon}(z)-1|\}$. Let $\gamma_{f^{\varepsilon}(z)}$ be a finite path starting from 0 that first separates $B(f^{\varepsilon}(z);\delta/4)$ and then separates $B(1;\delta/4)$ from infinity, but does not separate any points in $\mathbb{C}\backslash(B(f^{\varepsilon}(z);{\delta/2})\cup B(1;\delta/2))$ from infinity. Then using a similar argument in previous cases, where the Brownian motion has positive probability to stay sufficiently close to $\gamma_{f^{\varepsilon}(z)}$, and the random walks converges in law to Brownian motion in local uniform topology, we see that for each $\delta>0$ there exists $\delta^{\prime},\delta^{\prime\prime}>0$ such that for sufficiently small $\varepsilon$, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}<\tau_{z^{\prime}}]>\delta^{\prime}$ if $|z^{\prime}-f^{\varepsilon}(z)|>\delta$ whereas $\mathbb{P}_{\varepsilon}[\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}]<\delta^{\prime}$ if $|z^{\prime}-z|<\delta^{\prime\prime}$. In particular, there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z-z^{\prime}|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)-f^{\varepsilon}(z^{\prime})|<\delta$.
This finishes the proof for the tightness of $\{f^{\varepsilon}\}|_{z\in K}$ for each compact set $K$. By an identical argument, $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight for each compact set $K$. By the Arzelà-Ascoli theorem, for any subsequence of $\varepsilon$, there exists a further subsequence such that $f^{\varepsilon}$ converges uniformly on compacts to a continuous function $f$ on $\mathbb{C}$ such that $f(0)=0,f(1)=1$. By using that $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight and taking a further subsequence we may assume that $f$ is a homeomorphism. Then if $(W_{t})_{t\geq 0}$ is a Brownian motion from $z$, $f(W_{t})_{t\geq 0}$ has the law of a planar Brownian motion from $f(z)$ modulo time change. Then by Proposition 3.14, $f$ must be either the identity function or the reflection. This completes the proof. ∎
4 Convergence of cell configurations under the uniformization map
In this section we prove Theorem 1.8. Throughout this section, we assume that $\mathcal{H}$ is a random cell configuration that satisfies the constraints in Theorem 1.8. In Section 4.1, we collect several regularity estimates for the Riemann surface $M(\mathcal{H})$ under a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$, and prove that $M(\mathcal{H})$ is a.s. parabolic. In Sections 4.2 and 4.3, we use a continuum variant of the discrete argument from [GMS22, Section 2.3, Section 2.4], along with results from Section 4.1, to prove that there is a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ such that on a large scale, it is close to the function mapping vertices on $M(\mathcal{H})$ to their corresponding cells in $\mathcal{H}$. In Section 4.4, we conclude the proof of Theorem 1.8 by showing that the function $\phi_{\infty}\circ\varphi^{-1}$ must be a scalar times a deterministic linear transform $\mathbf{A}$ and a rotation. This argument is different from [GMS22] and gives a general and elementary method for proving that a particular class of random harmonic functions are of this form.
4.1 Regularity estimates for the Riemann surface
In this section, we prove several regularity estimates for the Riemann surface $M(\mathcal{H})$ under a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ that will be applied in the proof of Theorem 1.8. In Lemmas 4.34.4, and 4.5, we prove that if we have a microscopic set in $\mathcal{H}$ (resp. $\varphi(M(\mathcal{H}))$), then the corresponding set in $\varphi(M(\mathcal{H}))$ (resp. $\mathcal{H}$) will also be microscopic. Let $D_{k}$ be the union of $\varphi(P_{H})$ of the cells $H\in\mathcal{H}$ intersecting $2^{k},2^{k}]^{2}$. In Lemma 4.6, we prove that the diameter of $D_{k}$ has a growth rate faster than $2^{k/l}$ for some deterministic $l$. In Lemma 4.8, we prove that for large $k$, $D_{k}$ contains a disk with diameter at least $c\mathrm{diam}(D_{k})$. These lead to Lemma 4.9, where we prove that $M(\mathcal{H})$ is a.s. parabolic. Finally in Lemma 4.10, we prove that the function mapping points in $\varphi(M(\mathcal{H}))$ to their corresponding cells in $\mathcal{H}$ has a polynomial growth, which will be applied in Section 4.4 in the proof of Theorem 1.8.
The key idea in these proofs is as follows. For each cell $H\in\mathcal{H}$ and its vertex $v_{H}$ on $M(\mathcal{H})$, we associate them with a semi-flower $P_{H}\subseteq M(H)$ as in Figure 6. By Lemma 4.2 below, the geometry of a semi-flower under a uniformization map $\varphi$ can be bounded in terms of its degree. In [HS96, Equation (2.8)], He and Schramm induced a length-area argument in the context of circle packing, which originates from the Length-Area Lemma in [RS87]. Roughly speaking, one can control the diameter of a union of circles by estimating the area. Using a variant of this argument in the setting of the cell configuration $\mathcal{H}$ and semi-flowers, we can prove various bounds using ergodic averages of the degree of the cell configurations as in Lemma 2.7, which leads to the regularity results in this subsection.
Figure 5: An illustration of the definition of $M(S;a)$ and $M(S)$. For simplicity here we assume that two cells are adjacent if they have nonempty intersection. The cells in cyan are those intersecting $\mathring{S}^{a}$, and their neighbors are in yellow. We also include the cells that are separated from infinity from the yellow and cyan cells, which are colored in green. The surface $M(S;a)\subseteq M(\mathcal{H})$ are formed by vertices for these cells on $M(\mathcal{H})$ as well as the edges connecting these vertices, and the surface $M(S)$ is the largest surface of the form $M(S;a)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$.
For each square $S$, we define a simply connected surface $M(S)\subseteq M(\mathcal{H})$ as follows. For every $a>0$, let $\mathring{S}^{a}$ be the square concentric to $S$ with side length $a$. We work on the event that $\mathcal{H}(\mathring{S}^{a})$ is connected; otherwise we set $M(S;a)=\emptyset$. Let $M_{0}(S;a)\subseteq M(\mathcal{H})$ be the surface formed by the union of the (closed) equilateral triangles of $M(\mathcal{H})$ with at least one of the vertices in $\mathcal{H}(\mathring{S}^{a})$. Note that when viewed as a graph, $M_{0}(S;a)$ is 2-connected since $\mathcal{H}(\mathring{S}^{a})$ is connected. In fact, one can verify that, for any infinite plane triangulation $\mathcal{T}$ and a connected subgraph $\mathcal{G}_{1}\subseteq\mathcal{G}$, if we let $\mathcal{G}_{2}$ be the subgraph $\mathcal{G}$ formed by all the vertices in $\mathcal{G}_{1}$ and their neighbors, as well as all the edges connecting these vertices, then $\mathcal{G}_{2}$ is 2-connected. We further set $M(S;a)$ to be the union of $M_{0}(S;a)$ and the bounded connected components of $M(\mathcal{H})\backslash M_{0}(S;a)$. Then $M(S;a)$ has disk topology. Let $M(S)$ be the largest surface of the form $M(S;a)$ for some $a\in(0,|S|)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$. Note that these objects are well defined since $\mathcal{H}$ is locally finite. See Figure 5 for an illustration.
Lemma 4.1.
Almost surely, for each given $\varepsilon>0$ and $a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $M(S;(1-\varepsilon)|S|)\subseteq M(S)$.
Proof.
By the connectivity in Definition 1.4, for each $b\in(0,1)$, almost surely for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $\mathcal{H}(\mathring{S}^{b|S|})$ is a connected subgraph of $\mathcal{H}$. Let $H\in\mathcal{H}$ be such that its vertex $v_{H}$ on $M(\mathcal{H})$ is in $M(S;(1-\varepsilon)|S|)$. Then there are two cases: (i) $H$ is adjacent to some vertex $H^{\prime}\in\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$ or (ii) there exists a loop consisting of vertices for $H_{1},...,H_{n}$ on $M(\mathcal{H})$ surrounding $v_{H}$, such that each $H_{j}$ are adjacent to some cell in $\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$. In the setting (i), by Lemma 2.6, if we assume that $k$ is large, then $H\subseteq S$. For the case (ii), by Lemma 2.6, for large $k$, $H_{1},...,H_{n}\in\mathcal{H}(\mathring{S}^{(1-3\varepsilon/4)|S|})$. Letting $\{z_{H}:H\in\mathcal{H}\}$ and $\{e_{H,H^{\prime}}:H\sim H^{\prime}\}$ be the planar embedding of the map $\mathcal{H}$ as in the almost planarity in Definition 1.5, for large enough $k$, $e_{H_{1},H_{2}},...,e_{H_{n},H_{1}}\subseteq\mathring{S}^{(1-\varepsilon/2)|S|}$ and thus $z_{H}\in\mathring{S}^{(1-\varepsilon/2)|S|}$. By the same condition and Lemma 2.6 it follows that $H\subseteq S$. This finishes the proof. ∎
Let $\mathcal{T}$ be a disk triangulation or an infinite plane triangulation. Let $v$ be a vertex of $\mathcal{T}$ with degree $n$. The flower $F_{v}\subseteq M(\mathcal{T})$ is the union of the vertex $v$, the interior of the edges with an endpoint being $v$, and the interior of the equilateral triangles in $M(\mathcal{T})$ that meet at $v$, so $F_{v}$ has open disk topology. Define the semi-flower $P_{v}$ be the collection of points on $M(\mathcal{T})$ for which $v$ is the closest vertex, i.e., $P_{v}=\{x\in M(\mathcal{T}):d_{M}(x,v^{\prime})\geq d_{M}(x,v),\forall v^{\prime}\in\mathcal{V}\mathcal{T}\}$. Then $P_{v}\subseteq F_{v}$ is an $n$-gon (resp. $(n+2)$-gon) if $v$ is an interior (resp. boundary) vertex of degree $n$, and $\cup_{v\in\mathcal{V}\mathcal{T}}P_{v}=M(\mathcal{T})$.
In [GR13], the authors used the term half-flower to describe the shape $\frac{1}{2}F_{v}$, which is slightly different from our semi-flower. They worked on the interstices between the half-flowers, which always have bounded degree and bounded geometry (i.e., bounded ratio between outradius and inradius), to prove the parabolicity of Riemann surfaces associated with random triangulations obtained as the Benjamini-Schramm limit of disk triangulations. In Lemma 4.2 below, we prove that the geometry of semi-flowers can be controlled in terms of the square of the degrees, and we will combine this with the averaging over degrees as in Lemma 2.7 to prove the other results in this subsection.
Figure 6: An illustration of the semi-flower $P_{v}$ (green), its image under the conformal map $\varphi_{v}:F_{v}\to\mathbb{D}$ and a conformal map $\varphi:M(\mathcal{T})\to D\subseteq\mathbb{C}$.
Lemma 4.2.
There exists a universal constant $c_{0}>0$ such that the following is true. Let $\mathcal{T}$ be a disk triangulation or infinite plane triangulation and $\varphi$ be a conformal map from $M(\mathcal{T})$ to some domain $D\subseteq\mathbb{C}$. Let $v$ be an interior vertex of $\mathcal{T}$. Then
$B\Big(\varphi(v);c_{0}\deg(v)^{-2}\mathrm{diam}(\varphi(P_{v}))\Big)\subseteq\varphi(P_{v}).$ (4.1)
If $\mathcal{T}$ is a disk triangulation, $D=\mathbb{D}$, $v$ is a boundary vertex of $\mathcal{T}$, then
$B\Big(\varphi(v);c_{0}\deg(v)^{-2}\mathrm{diam}(\varphi(P_{v}))\Big)\cap\mathbb{D}\subseteq\varphi(P_{v}).$ (4.2)
Proof.
Let $n=\deg(v)$. We start with (4.1). Consider the conformal map $\varphi_{0}$ from the equilateral triangle $\Delta$ with vertices $0,1,e^{\pi i/3}$ to $\widetilde{\Delta}:=\{re^{i\theta}:0<r<1,0<\theta<\pi/3\}$ fixing $0,1,e^{\pi i/3}$. Let $\Delta_{0}=\{z\in\Delta:\operatorname{dist}(z,0)\leq\min\{\operatorname{dist}(z,1),\operatorname{dist}(z,e^{i\pi/3}\}\}$. Then there exists constants $a_{1},a_{2}\in(0,1)$ such that $B(0,a_{1})\cap\widetilde{\Delta}\subseteq\varphi_{0}(\Delta_{0})\subseteq B(0,a_{2})\cap\widetilde{\Delta}$. Let $\varphi_{v}:F_{v}\to\mathbb{D}$ be a conformal map sending $v$ to 0. Then using the map $z\mapsto z^{6/n}$ and symmetry, we have $B(0,a_{1}^{6/n})\subseteq\varphi_{v}(P_{v})\subseteq B(0,a_{2}^{6/n})$. Now $\varphi\circ\varphi_{v}^{-1}$ is conformal on $\mathbb{D}$, and by the Koebe distortion theorem, we have
$B\Big(\varphi(v),\frac{a_{1}^{6/n}}{(1+a_{1}^{6/n})^{2}}|(\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|\Big)\subseteq\varphi(P_{v})\subseteq B\Big(\varphi(v),\frac{a_{2}^{6/n}}{(1-a_{2}^{6/n})^{2}}|(\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|\Big).$ (4.3)
One can check that $\frac{a_{2}^{6/n}}{(1-a_{2}^{6/n})^{2}}\big/\frac{a_{1}^{6/n}}{(1+a_{1}^{6/n})^{2}}\leq cn^{2}$ for some constant $c$ depending only on $a_{1},a_{2}$, and this verifies (4.1).
For (4.2), since $M(\mathcal{T})$ is simply connected, there are precisely two edges $e$ and $e^{\prime}$ on $M(\mathcal{T})$ with $v$ being one of endpoints which are adjacent to only one equilateral triangle of $M(\mathcal{T})$. Consider the conformal map from $\xi:\mathbb{D}\to\mathbb{H}$ sending $(0,\varphi(v))$ to $(i,0)$, and a conformal map $\varphi_{v}:F_{v}\to\mathbb{D}\cap\mathbb{H}$ sending the edges $e$ and $e^{\prime}$ to [1,0][-1,0] and $[0,1]$, respectively. Then by symmetry, $B(0,a_{1}^{3/(n-1)})\cap\mathbb{H}\subseteq\varphi_{v}(P_{v})\subseteq B(0,a_{2}^{3/(n-1)})$, and $\xi\circ\varphi\circ\varphi_{v}^{-1}$ is conformal on $\mathbb{D}\cap\mathbb{H}$. We can extend $\xi\circ\varphi\circ\varphi_{v}^{-1}$ to be conformal on $\mathbb{D}$ by Schwarz reflection. Thus by the Koebe distortion theorem,
$B\Big({0},\frac{a_{1}^{3/(n-1)}}{(1+a_{1}^{3/(n-1)})^{2}}|(\xi\circ\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|\Big)\cap\mathbb{H}\subseteq(\xi\circ\varphi)(P_{v})\subseteq B\Big(0,\frac{a_{2}^{3/(n-1)}}{(1-a_{2}^{3/(n-1)})^{2}}|(\xi\circ\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|\Big).$ (4.4)
The claim then follows by applying the map $\xi^{-1}$ to (4.4) and $\frac{a_{2}^{3/(n-1)}}{(1-a_{2}^{3/(n-1)})^{2}}\big/\frac{a_{1}^{3/(n-1)}}{(1+a_{1}^{3/(n-1)})^{2}}\leq cn^{2}$ for some constant $c>0$; if $\frac{a_{2}^{3/(n-1)}}{(1-a_{2}^{3/(n-1)})^{2}}|(\xi\circ\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|\leq 0.01$, then the claim follows since $(\xi^{-1})^{\prime}$ is analytic near 0, and if $\frac{a_{2}^{3/(n-1)}}{(1-a_{2}^{3/(n-1)})^{2}}|(\xi\circ\varphi\circ\varphi_{v}^{-1})^{\prime}(0)|>0.01$, we further use that $\big(\xi^{-1}\circ(\xi\circ\varphi)\big)(P_{v})\subseteq\mathbb{D}$. ∎
For each $k>0$, let $\varphi_{k}$ be the uniformization map sending $M(S_{k})$ to $|S_{k}|\mathbb{D}$ such that the vertex on $M(S_{k})$ for the cell containing the center of $S_{k}$ is sent to 0. Recall the definition of the semiflower in Lemma 4.2. For each $H\in\mathcal{H}$, we write $P_{H}$ for the semiflower for the vertex $H$ on the surface $M(\mathcal{H})$.
Lemma 4.3.
There exists a deterministic constant $c$ such that the following true. Almost surely, for each $\delta\in(0,0.01)$, for large enough $k$, for each square $S\subseteq S_{k}$ with side length $\delta|S_{k}|$,
$\mathrm{diam}\Big(\bigcup_{H\in\mathcal{H}(S)}\varphi_{k}(P_{H})\Big)\leq c(-\log\delta)^{-1/2}|S_{k}|,$ (4.5)
where we use the convention $\varphi_{k}(P_{H})=\emptyset$ if $P_{H}\cap M(S_{k})=\emptyset$. Furthermore, for any $\delta>0$, almost surely for large enough $k$, for each $H\in\mathcal{H}(S_{k})$, $\mathrm{diam}(\varphi_{k}(P_{H}))<\delta|S_{k}|$.
Figure 7: An illustration of the proof of Lemma 4.3, where the left panel is a part of our cell configuration and the right panel is the corresponding image of the part of the surface $M(\mathcal{H})$ under a conformal map. The squares concentric to $S$ of side lengths $r_{0}$ and $r_{1}$ are colored in blue on the left panel. In (4.12), we prove that one can find some $\rho\in[r_{0},r_{1}]$, such that under a conformal map the union of the semi-flowers for the cells intersecting the square concentric to $S$ of side length $\rho$ (shaded in light blue on the right panel) has diameter less than $c(-\log\delta)^{-1/2}|S_{k}|$. Further using the line connectivity property and almost planarity in Definition 1.4 and Definition 1.5 gives the claim.
Proof.
We adapt the argument from [HS96, Equation (2.8)], which studies maps between two circle packings. We sum over the diameter of $\varphi(P_{H})$ for $H$ intersecting $\partial\mathring{S}^{\rho}$, integrate over $\rho$, apply a Cauchy-Schwarz inequality, and the terms in this Cauchy-Schwarz inequality can be described in terms of area. One key difference is that the image $\varphi(P_{H})$ is no longer exactly a round disk, and there is no direct way to bound the sum over $\mathrm{diam}(\varphi(P_{H}))^{2}$. To circumvent this issue, we use the information from Lemma 4.2 that $\varphi(P_{H})$ contains a disk with radius $c\deg(H)^{-2}\mathrm{diam}(\varphi(P_{H}))$, and add an additional $\deg(H)^{-2}$ term in the Cauchy-Schwarz inequality (4.8) below. Then the sum over $\mathrm{diam}(\varphi(P_{H}))^{2}\deg(H)^{-4}$ can be bounded by the total area of a collection of disjoint disks, and the sum over $\deg(H)^{4}\mathrm{diam}(H)^{4}$ can be controlled by the ergodic average from Lemma 2.7.
Let $r_{0}=2\delta|S_{k}|$ and $r_{1}=2^{\ell}r_{0}$ such that $\ell\in\mathbb{N}$ and $r_{1}\in[|S_{k}|/8,|S_{k}|/4)$. For $\rho\in[r_{0},r_{1}]$, set
$L(\rho)=\sum_{H\in\mathcal{H}(\partial\mathring{S}^{\rho})}\mathrm{diam}(\varphi_{k}(P_{H})).$ (4.6)
See also Figure 7 for an illustration. Then
$inf\{L(\rho):r_{0}\leq\rho\leq r_{1}\}\log(r_{1}/r_{0})\leq\int_{r_{0}}^{r_{1}}\frac{L(\rho)}{\rho}\,d\rho=\sum_{H\in\mathcal{H}(\mathring{S}^{r_{1}}\backslash\mathring{S}^{r_{0}})}\mathrm{diam}(\varphi_{k}(P_{H}))\int_{r_{0}}^{r_{1}}\frac{1}{\rho}\mathds{1}_{H\in\mathcal{H}(\partial\mathring{S}^{\rho})}\,d\rho.$ (4.7)
[…]
$\int_{r_{0}}^{r_{1}}\frac{1}{\rho}\mathds{1}_{H\in\mathcal{H}(\partial\mathring{S}^{\rho})}\,d\rho\leq\frac{\mathrm{diam}(H)}{a_{H}}.$
Therefore by the Cauchy-Schwarz inequality, the right side of (4.7) is less than or equal to
[…]
By Lemma 4.2, for some constant $c_{0}$, each set $\varphi_{k}(P_{H})$ contains a disk of radius $c_{0}\mathrm{diam}(\varphi_{k}(P_{H}))\deg(H)^{-2}$, and since the sets $\{\varphi_{k}(H):H\in\mathcal{H}\}$ have disjoint interiors and are contained within $|S_{k}|\mathbb{D}$,
[…]
where $C$ is the deterministic constant in (2.6). Summing over $j=0,...,\ell-1$,
[…]
Combining (4.9) and (4.11), we conclude from (4.8) that for some deterministic constant $c_{2}>0$, almost surely, for all large enough $k$,
$\inf\{L(\rho):r_{0}\leq\rho\leq r_{1}\}\leq c_{2}\ell^{-1/2}|S_{k}|.$ (4.12)
In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L(\rho)\leq 4c_{2}(-\log\delta)^{-1/2}|S_{k}|$.
If $\mathring{S}^{\rho}\subseteq\mathring{S}_{k}^{7|S_{k}|/8}$, using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ intersecting $\partial\mathring{S}^{\rho}$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells, and in the planar embedding $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ in Condition 4. form a loop around $\mathring{S}^{1.5\delta|S_{k}|}$, and thus encloses all the points $z_{H}$ for $H\in{S}$. This implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ is separated from infinity by $\cup_{j=1}^{n}\varphi(P_{H_{j}})$, which together with (4.12) implies (4.5).
If $\mathring{S}^{\rho}\nsubseteq\mathring{S}_{k}^{7{|S_{k}|}/8}$, since $\rho<r_{1}\leq|S_{k}|/4$, $\mathring{S}^{\rho}$ is disjoint from the cell with closest distance to the center of $S_{k}$. Using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ either intersecting $\partial\mathring{S}^{\rho}$ or with their vertex on $M(\mathcal{H})$ on the boundary of the surface $M(S_{k})$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells and the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ enclose a loop around all the points $z_{H}$ where $H\in\mathcal{H}(S)$ and the vertex for $H$ is on $M(S_{k})$. This together with (4.12) implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ can be separated from 0 by $\partial(|S_{k}|\mathbb{D})$ and some curves of total length $4c_{2}(\log\delta)^{-1/2}|S_{k}|$. This finishes the proof of (4.5).
In order to get the second assertion of the lemma, let $\widetilde{\delta}\in(0,0.01)$ be such that $c(-\log\widetilde{\delta})^{-1/2}<\delta$ and let $S\subseteq S_{k}$ be a square of side length $\widetilde{\delta}|S_{k}|$ such that $H\cap S\neq\emptyset$. Then apply (4.5) with $\widetilde{\delta}$ instead of $\delta$. ∎
The following is the analog of Lemma 4.3 in the reverse direction.
Lemma $4.$
There exists a deterministic constant $c>0$ such that the following is true. Almost surely, for each $\delta\in(0,0.01)$, $\lambda\in(0,1-16\delta)$ and $\varepsilon>0$, for large enough $k$, for $z\in B(0;\lambda|S_{k}|)$ and $r=\delta|S_{k}|$,
$\mathrm{diam}\Big(\bigcup_{H:\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset}H\Big)\leq c(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|.$ (4.13)
Proof.
Fix $z\in\mathbb{C}$. Let $r_{0}=2\delta|S_{k}|$ and $r_{1}=2^{\ell}r_{0}$ such that $\ell\in\mathbb{N}$ and $r_{1}\in[(1-\lambda)|S_{k}|/8,(1-\lambda)|S_{k}|/4)$. For $\rho\in[r_{0},r_{1}]$, we write
$L^{*}(\rho)=\sum_{H:\varphi_{k}(P_{H})\cap\partial B(z;\rho)\neq\emptyset}\mathrm{diam}(H).$
$\begin{split}\inf\{L^{*}(\rho):&r_{0}\leq\rho\leq r_{1}\}\log(r_{1}/r_{0})\leq\int_{r_{0}}^{r_{1}}\frac{L^{*}(\rho)}{\rho}\,d\rho\\ &=\sum_{H:\varphi_{k}(P_{H})\cap(B(z;r_{1})\backslash B(z;r_{0}))\neq\emptyset}\mathrm{diam}(H)\int\frac{1}{\rho}\mathds{1}_{\varphi_{k}(P_{H})\cap\partial B(z;\rho)\neq\emptyset}\,d\rho.\end{split}$ (4.14)
Let $a_{H}^{*}=\inf\{\rho>0:\varphi_{k}(P_{H})\cap\partial B(z;\rho)\neq\emptyset\}$, then
$\int\frac{1}{\rho}\mathds{1}_{\varphi_{k}(P_{H})\cap\partial B(z;\rho)\neq\emptyset}\,d\rho\leq\frac{\mathrm{diam}(\varphi_{k}(H))}{a_{H}^{*}}.$
Therefore by the Cauchy-Schwarz inequality, the right side of (4.14) is less than or equal to
$\Big(\sum_{H:\varphi_{k}(P_{H})\cap(B(z;r_{1})\backslash B(z;r_{0}))\neq\emptyset}\mathrm{diam}(H)^{2}\deg(H)^{4}\Big)^{1/2}\Big(\sum_{H:\varphi_{k}(P_{H})\cap(B(z;r_{1})\backslash B(z;r_{0}))\neq\emptyset}\frac{\mathrm{diam}(\varphi_{k}(H))^{2}}{(a_{H}^{*})^{2}\deg(H)^{4}}\Big)^{1/2}.$ (4.15)
By Lemma 2.7, for some deterministic constant $C>0$ and large enough $k$,
$\sum_{H:\varphi_{k}(P_{H})\cap(B(z;r_{1})\backslash B(z;r_{0}))\neq\emptyset}\mathrm{diam}(H)^{2}\deg(H)^{4}\leq\sum_{H\in\mathcal{H}(S_{k})}\mathrm{diam}(H)^{2}\deg(H)^{4}\leq C|S_{k}|^{2}.$ (4.16)
By the last claim of Lemma 4.3, a.s. for large enough $k$, for $j=0,1,...,\ell-1$ and $H$ such that $\varphi_{k}(P_{H})\cap(B(z;2^{j+1}r_{0})\backslash B(z;2^{j}r_{0}))\neq\emptyset$, $a^{*}_{H}\geq 2^{j-1}r_{0}$ and $\varphi_{k}(P_{H})\subseteq B(z;2^{j+2}r_{0})$. Thus by an application of Lemma 4.2, for some constant $c_{0}$,
$\sum_{H:\varphi_{k}(P_{H})\cap(B(z;2^{j+1}r_{0})\backslash B(z;2^{j}r_{0}))\neq\emptyset}\frac{\mathrm{diam}(\varphi_{k}(H))^{2}}{(a_{H}^{*})^{2}\deg(H)^{4}}\leq c_{0}^{-2}(2^{j-1}r_{0})^{-2}\mathrm{area}(B(z;2^{j+2}r_{0}))=64{\pi}c_{0}^{-2}.$ (4.17)
Summing over $j=0,...,\ell-1$,
$\sum_{H:\varphi_{k}(P_{H})\cap(B(z;r_{1})\backslash B(z;r_{0}))\neq\emptyset}\frac{\mathrm{diam}(\varphi_{k}(H))^{2}}{(a_{H}^{*})^{2}\deg(H)^{4}}\leq 64c_{0}^{-2}\ell.$ (4.18)
Combining (4.16) and (4.18), we conclude from (4.15) that for some constant $c_{2}$, for large enough $k$,
$\inf\{L^{*}(\rho):r_{0}\leq\rho\leq r_{1}\}\leq c_{2}(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|.$ (4.19)
In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L^{*}(\rho)\leq 4c_{2}(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|$. Recall the points $\{z_{H}:H\in\mathcal{H}\}$ and the curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in the planar embedding from the almost planarity condition in Definition 1.5. Note that by our choice of $r_{1}$, $B(z;\rho)\subseteq|S_{k}|\mathbb{D}$. One can find edges $e_{1},...,e_{n}$ forming a loop in $\mathcal{H}$ such that $\gamma_{e_{1}},...,\gamma_{e_{n}}$ make a loop surrounding all the vertices $z_{H}$ with $\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset$, and the union of theses edges has diameter less than $20c_{2}(-\log\delta)^{-1/2}|S_{k}|$. Applying  (1.2) once more gives (4.13). ∎
We will also need the version of Lemmas 4.3 and 4.4 for $2^{k},2^{k}]^{2}$ instead of $S_{k}$. With slight abuse of notation, we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$. For each $k>0$, let $\widehat{\varphi}_{k}$ be the uniformization map sending $M([-2^{k},2^{k}]^{2})\subseteq M(\mathcal{H})$ to $2^{k}\mathbb{D}$ and sending $0\in M(\mathcal{H})$ to 0.
Lemma 4.5.
There exists a deterministic constant $c$ such that the following true. Almost surely, for each $\delta\in(0,0.01)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ with side length $2^{k}\delta$,
$\mathrm{diam}\Big(\bigcup_{H\in\mathcal{H}(S)}\widehat{\varphi}_{k}(P_{H})\Big)\leq c(-\log\delta)^{-1/2}2^{k}.$ (4.20)
Conversely, for each $\lambda\in(0,1-\delta)$ and $z\in B(0;\lambda 2^{k})$, almost surely for large enough $k$,
$\mathrm{diam}\Big(\bigcup_{H:\widehat{\varphi}_{k}(P_{H})\cap B(z;r)\neq\emptyset}H\Big)\leq c(\log(1-\lambda)-\log\delta)^{-1/2}2^{k}.$ (4.21)
Proof.
The proof is identical to that of Lemma 4.3 and Lemma 4.4 using Lemma 2.6 and Lemma 2.7. ∎
Lemma 4.6.
There exist deterministic constants $\alpha,c_{1}\in(0,1/2)$ such that the following is true. Almost surely, for large enough $k$,
$B(0;c_{1}2^{k})\subseteq\widehat{\varphi}_{k}(M([-\alpha 2^{k},\alpha 2^{k}]^{2}))\subseteq B(0;2^{k-1}).$ (4.22)
Proof.
Let $c>0$ be the constant in Lemma 4.5. We pick $\alpha\in(0,0.01)$ such that $c(-\log\alpha)^{-1/2}<1/4$. Then by (4.20) for $S=[-\alpha 2^{k},\alpha 2^{k}]^{2}$,
$\bigcup_{H\in\mathcal{H}([-\alpha 2^{k},\alpha 2^{k}]^{2})}\widehat{\varphi}_{k}(P_{H})\subseteq B(0;2^{k-1}).$
On the other hand, if we pick $c_{1}\in(0,0.01)$ such that $c(-\log c_{1})^{1/2}<\alpha/16$, then by (4.21) for $B(z;r)=B(0;c_{1}2^{k})$,
$\bigcup_{H:\widehat{\varphi}_{k}(P_{H})\cap B(0,c_{1}2^{k})\neq\emptyset}H\subseteq[-\alpha 2^{k-2},\alpha 2^{k-2}]^{2}.$
This along with Lemma 4.1 verifies (4.22). ∎
With slight abuse of language, we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$. Consider a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ sending 0 to 0. Let $\alpha$ be the constant as in Lemma 4.6. We define
$d_{k}:=\mathrm{diam}\big(\varphi(M([-\alpha 2^{k},\alpha 2^{k}]^{2}))\big).$ (4.23)
Lemma 4.7.
Let $a\in(0,1)$. There exists a deterministic $\ell>0$ such that almost surely, for large enough $k$, $d_{k-\ell}<ad_{k}$.
Proof.
The proof is almost identical to that of Lemma 4.3. Fix $\ell\in\mathbb{N}$, $r_{0}=\alpha 2^{k-\ell+1}$, $r_{1}=\alpha 2^{k-2}$, and for $\rho\in[r_{0},r_{1}]$, set
$L(\rho)=\sum_{H\in\mathcal{H}(\partial[-\rho,\rho]^{2})}\mathrm{diam}(\varphi(P_{H})).$ (4.24)
Then
$\begin{split}&\inf\{L(\rho):r_{0}\leq\rho\leq r_{1}\}\log(r_{1}/r_{0})\leq\int_{r_{0}}^{r_{1}}\frac{L(\rho)}{\rho}\,d\rho=\sum_{H\in\mathcal{H}([-r_{1},r_{1}]^{2}\backslash[-r_{0},r_{0}]^{2})}\mathrm{diam}(\varphi(P_{H}))\frac{\mathrm{diam}(H)}{a_{H}}\\ &\leq\Big(\sum_{H\in\mathcal{H}([-r_{1},r_{1}]^{2}\backslash[-r_{0},r_{0}]^{2})}\mathrm{diam}(\varphi(P_{H}))^{2}\deg(H)^{-4}\Big)^{1/2}\Big(\sum_{H\in\mathcal{H}([-r_{1},r_{1}]^{2}\backslash[-r_{0},r_{0}]^{2})}\frac{\mathrm{diam}(H)^{2}}{a_{H}^{2}}\deg(H)^{4}\Big)^{1/2}\end{split}$ (4.25)
where $a_{H}=\inf\{\rho>0:H\in\mathcal{H}(\partial[-\rho,\rho]^{2})\}$. By Lemma 4.1 and the definition of $d_{k}$, $\bigcup_{H\in\mathcal{H}([-r_{1},r_{1}]^{2})}\varphi(P_{H})\subseteq B(0;d_{k})$. By Lemma 4.2, for some constant $c_{0}$,
$\sum_{H\in\mathcal{H}([-r_{1},r_{1}]^{2}\backslash[-r_{0},r_{0}]^{2})}\mathrm{diam}(\varphi(P_{H}))^{2}\deg(H)^{-4}\leq\pi c_{0}^{-2}d_{k}^{2}.$ (4.26)
By a similar argument as in (4.11), for some deterministic constant $C>0$,
$\sum_{H\in\mathcal{H}([-r_{1},r_{1}]^{2}\backslash[-r_{0},r_{0}]^{2})}\frac{\mathrm{diam}(H)^{2}}{a_{H}^{2}}\deg(H)^{4}\leq 64C\ell.$ (4.27)
Thus by (4.25),  (4.26) and (4.27), for some deterministic constant $c_{2}>0$, almost surely for large enough $k$,
$\inf\{L(\rho):r_{0}\leq\rho\leq r_{1}\}\leq c_{2}\ell^{-1/2}d_{k}.$ (4.28)
We pick $\ell$ such that $c_{2}\ell^{-1/2}<a/4$. The lemma then follows using the same argument as in the last paragraph of the proof of Lemma 4.3. ∎
Lemma 4.8.
There exists deterministic constants $\alpha,c_{2}\in(0,1)$ and $\ell>0$ such that the following is true. Almost surely, for large enough $k$,
$\varphi(M([-\alpha 2^{k-\ell},\alpha 2^{k-\ell}]^{2}))\subseteq B\big(0;c_{2}d_{k}\big)\subseteq\varphi(M([-\alpha 2^{k},\alpha 2^{k}]^{2})).$ (4.29)
Proof.
The second half of (4.29) is immediate by applying the Koebe distortion theorem to the conformal map $\varphi\circ\widehat{\varphi}_{k}^{-1}$ along with the first half of (4.22). The first half of (4.29) then follows from Lemma 4.7 with $a=c_{2}$. ∎
Lemma 4.9.
Almost surely, $\lim_{k\to\infty}d_{k}=\infty$. Furthermore, $\varphi(M(\mathcal{H}))=\mathbb{C}$ and thus $M(\mathcal{H})$ is parabolic.
Proof.
The first claim is immediate from Lemma 4.7. The second claim follows since, by Lemma 4.8, for some constant $c_{2}>0$, for all large enough $k$, $B(0;c_{2}d_{k})\subseteq\varphi(M(\mathcal{H}))$. ∎
Lemma 4.10.
There is a deterministic constant $\beta\geq 1$ such that the following is true. Almost surely, for some random constant $\widetilde{C}$ depending on $\mathcal{H}$ and $\varphi$, for every large enough $r$,
$\{H:\varphi(P_{H})\cap B(0;r)\neq\emptyset\}\subseteq B(0;\widetilde{C}r^{\beta}).$ (4.30)
Proof.
Let $\ell$ be the constant in Lemma 4.7 for $a=1/2$. Then there exists some random constant $\widetilde{C}_{0}$, such that $d_{k}\geq\widetilde{C}_{0}2^{k/\ell}$. By Lemma 4.8, for some constant $c_{2}>0$ and $k$ large enough, if $\varphi(P_{H})\cap B(0;c_{2}\widetilde{C}_{0}2^{k/\ell})\neq\emptyset$, then $H\in\mathcal{H}([-\alpha 2^{k},\alpha 2^{k}]^{2})$ or $H$ is adjacent to some $H^{\prime}\in\mathcal{H}([-\alpha 2^{k},\alpha 2^{k}]^{2})$. In particular, by Lemma 2.6, $H\subseteq[-2^{k},2^{k}]^{2}$. Now for each $r>0$ large enough, choose $k$ such that $c_{2}\widetilde{C}_{0}2^{k/\ell}<r<c_{2}\widetilde{C}_{0}2^{(k+1)/\ell}$. If $\varphi(P_{H})\cap B(0;r)\neq\emptyset$, then $H\subseteq[-2^{k+1},2^{k+1}]^{2}\subseteq B(0;4(c_{2}\widetilde{C}_{0})^{-\ell}r^{\ell})$. This verifies the claim for $\beta=\ell$. ∎
Remark 4.11.
We note that the proofs of Lemmas 4.3-4.7 can be adapted to the setting of circle packings. Therefore, by repeating the arguments in this subsection, it might be possible to give an alternative proof of Proposition 3.11 without using the vertex extremal length in [GJN19].
4.2 Existence of the harmonic corrector
Let $\mathcal{H}$ be a random cell configuration satisfying all the conditions in Theorem 1.8 and let $M(\mathcal{H})$ be the Riemann surface constructed from the map associated with $\mathcal{H}$. In this section we will define a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ as the limit of functions $\phi_{m}:M(\mathcal{H})\to\mathbb{C}$ for $m\in\mathbb{N}$, which are defined as certain harmonic extensions of an appropriately chosen function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. This idea is similar to [GMS22, Section 2.3] and the key difference is that we work with (continuum) harmonic functions on $M(\mathcal{H})$ instead of discrete harmonic functions on the graph $\mathcal{H}$.
The function $\phi_{0}$ sends each vertex on $M(\mathcal{H})$ to a point in the associated cell of $\mathcal{H}$ and is otherwise piecewise linear. More precisely, the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ is defined as follows. Given a realization of $\mathcal{H}$, for each cell $H\in\mathcal{H}$, we independently sample a point $\widetilde{c}(H)$ according to the probability measure proportional to the Lebsegue measure on $H$. For each triangular face $\Delta$ in $M(\mathcal{H})$ formed by the cells $H_{1},H_{2},H_{3}\in\mathcal{H}$, consider the triangle $\Delta^{\prime}$ in $\mathbb{C}$ formed by connecting $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ using straight line segments, and define $\phi_{0}|_{\Delta}:\Delta\to\Delta^{\prime}$ to be the unique linear function sending the vertices in $\Delta$ to their counterpart in $\Delta^{\prime}$. This defines a continuous and piecewise linear function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. See Figure 8 for an illustration.
Figure 8: An illustration for the definition of the map $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$, which linearly send each triangle on the left panel to the corresponding triangle on the right panel. Note that the $\phi_{0}$-images of different equilateral triangles of $M(\mathcal{H})$ might overlap with each other.
Recall the definition $\widehat{S}_{m}^{z}$ in (2.2) and define the collection of squares
[…]
If $z^{\prime}$ is in the interior of $\widehat{S}_{m}^{z}$, then $\widehat{S}_{m}^{z}=\widehat{S}_{m}^{z^{\prime}}$. It follows that the squares in $\mathfrak{S}_{m}$ intersect only along their boundaries, and they together cover all of $\mathbb{C}$. In particular, the preimage $\{\phi_{0}^{-1}(S):S\in\mathfrak{S}_{m}\}$ cover $M(\mathcal{H})$. Note that $\phi_{0}^{-1}(S)$ might not necessarily be connected, and their boundaries consist of line segments.
For $m>0$, we define the function $\phi_{m}{:M(\mathcal{H})\to\mathbb{C}}$ as follows. For each $S\in\mathfrak{S}_{m}$ and $\phi_{0}^{-1}(S)$, we define $\phi_{m}$ to be equal to $\phi_{0}$ on $\partial\phi_{0}^{-1}(S)$, and define $\phi_{m}$ to be the harmonic extension of $\phi_{0}$ in the interior of $\phi_{0}^{-1}(S)$. Equivalently, one can take a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ and work on the ordinary harmonic extension of $\phi_{0}\circ\varphi|_{\partial(\varphi\circ\phi_{0}^{-1})(S)}$ in $(\varphi\circ\phi_{0}^{-1})(S)$.
The main goal of this subsection is to prove the following. We do not know at this point whether $\phi_{\infty}(x)$ is a measurable function of $\mathcal{H}$ and $\mathcal{D}$ due to the randomness of the points $\widetilde{c}(H)$. In the following proposition we work with $\mathbb{E}\big[\phi_{\infty}(x)|\mathcal{H},\mathcal{D}\big]$, where the randomness that we average over in the expectation is coming from the randomness of the points $\widetilde{c}(H)$. Recall that we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$.
Proposition 4.12.
There exists a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ such that $\mathbb{E}\big[\phi_{\infty}(x)|\mathcal{H},\mathcal{D}\big]<\infty$ a.s. for each $x\in M(\mathcal{H})$ and
$\phi_{m}-\phi_{m}(0)\to\phi_{\infty}\ \ \text{in probability as $m\to\infty$},$ (4.31)
with respect to the topology of uniform convergence on compact subsets of $M(\mathcal{H})$, i.e., for each compact set $\mathcal{A}\subseteq M(\mathcal{H})$ (chosen in some measurable manner) and each $\varepsilon>0$,
$\lim_{m\to\infty}\mathbb{P}\big[\sup_{z\in\mathcal{A}}|\phi_{m}(z)-\phi_{m}(0)-\phi_{\infty}(z)|>\varepsilon\big]=0.$
To prove Proposition 4.12, we use the similar outline as in [GMS22, Section 2.3]. However, instead of working on discrete Dirichlet energy and discrete harmonic functions $\mathcal{H}\to\mathbb{C}$ as in [GMS22], we work with continuous harmonic functions and their Dirichlet energy. Our particular choice of $\phi_{0}$ above will play an important role in order for the argument to work; it will be important that the images of vertices on $M(\mathcal{H})$ under $\phi_{0}$ are close to the corresponding cells in $\mathcal{H}$ and that we can bound the Dirichlet energy per area for $\phi_{0}$ in terms of $\mathrm{diam}(H)^{2}\deg(H)$ via Lemma 2.7 (see Lemma 4.17 below). Different from the setup of [GMS22], where the function $\phi_{0}$ is defined on $\mathcal{H}$, our function $\phi_{0}$ and their harmonic extensions $\phi_{m}$ are defined on the surface $M(\mathcal{H})$. First we gather some topological properties of the map $\phi_{0}$. Then we show that $\phi_{0}$ (and hence $\phi_{m}$) has uniformly bounded Dirichlet energy on $\phi_{0}^{-1}(S_{n})$ when rescaled by $\frac{1}{|S_{n}|^{2}}$. Then we use the orthogonality of the increments $\phi_{m}-\phi_{m-1}$ with respect to the Dirichlet inner product (see the proof of Lemma 4.19) to show that for large enough $m$, the expected “Dirichlet energy per unit area” is small uniformly over all $m^{\prime}\geq m$. Together with the standard properties of ordinary harmonic functions, this will allow us to show that the functions $\phi_{m}-\phi_{m}(0)$ are Cauchy in probability, and thereby establish Proposition 4.12.
For a function $f:M(\mathcal{H})\to\mathbb{C}$ and some domain $D\subseteq M(\mathcal{H})$, using the conformal invariance of Dirichlet energy, we set
$\mathrm{Energy}(f;D)=\int_{D}|\nabla f(x)|^{2}dx=\int_{\varphi(D)}|\nabla(f\circ\varphi^{-1})(z)|^{2}dz,$
where $\varphi:M(\mathcal{H})\to\mathbb{C}$ is some conformal map.
We start with the following simple lemma.
Lemma 4.13.
For any linear map $f$ from an equilateral triangle $\Delta$ with unit side length to another triangle $\Delta^{\prime}$ with side length $\ell_{1},\ell_{2},\ell_{3}$,
$\int_{\Delta}|\nabla f(z)|^{2}dz\leq 3(\ell_{1}^{2}+\ell_{2}^{2}+\ell_{3}^{2}).$
Proof.
Without loss of generality, by translation and rotation, assume the vertices of $\Delta^{\prime}$ are positioned at $(0,0),(x_{1},0),(x_{2},y_{2})$. One can check that the matrix for the linear map $f$ is given by
$\begin{pmatrix}x_{1}&-\frac{1}{\sqrt{3}}x_{1}+\frac{2}{\sqrt{3}}x_{2}\\ 0&\frac{2}{\sqrt{3}}y_{2}\end{pmatrix}$.
Then $|\nabla f|^{2}\leq 3(x_{1}^{2}+x_{2}^{2}+y_{2}^{2})\leq 3\ell_{1}^{2}+3\ell_{2}^{2}+3\ell_{3}^{2}$. ∎
Lemma 4.14.
Almost surely, for large enough $k$, the following holds. If $H_{1},H_{2},H_{3}\in\mathcal{H}$ are adjacent to each other, such that the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ intersects $S_{k}$, then $H_{1},H_{2},H_{3}$ are contained within the square $S_{k}^{\prime}$ concentric to $S_{k}$ with side length $1.5|S_{k}|$.
Proof.
This is a direct consequence of Lemma 2.6. ∎
We will also need the following topological properties on $\phi_{0}$.
Lemma 4.15.
We have $\phi_{0}(M(\mathcal{H}))=\mathbb{C}$.
Proof.
Pick $k$ large enough such that by Definition 1.4, the 4 subgraphs of $\mathcal{H}$ formed by the cells intersecting the 4 sides of $S_{k}$ are connected. Since a cell intersecting a vertex of $S_{k}$ is contained in (at least) two of these subgraphs, we get that the set of cells intersecting a side of $S_{k}$ forms a connected graph. By further applying Lemma 2.6, we can find a chain of cells $H_{1}\sim H_{2}\sim\cdots\sim H_{n}\sim H_{1}$ in this connected graph such that if we connect $\widetilde{c}(H_{j})$ and $\widetilde{c}(H_{j+1})$ using straight line segments for j=1,,nj=1,...,n with $n+1$ identified with 1, then this creates a loop $\gamma_{n}$ surrounding the square $\widetilde{S}_{k}$ concentric to $S_{k}$ with side length $|S_{k}|/2$, and the winding number of $\gamma_{n}$ around each point in $\widetilde{S}_{k}$ is equal to 1. If we consider the vertices for $H_{1},...,H_{n}$ on the surface $M(\mathcal{H})$, then one can find a loop $\widetilde{\gamma}_{n}$ on $M(\mathcal{H})$ connecting these vertices with straight line segments such that $\phi_{0}(\widetilde{\gamma}_{n})=\gamma_{n}$. Since $M(\mathcal{H})$ is simply connected, we can continuously shrink $\widetilde{\gamma}_{n}$ to a single point $x$. If there exists a point $w\in\widetilde{S}_{k}\backslash\phi_{0}(M(\mathcal{H}))$, then under $\phi_{0}$ image, all of these shrinking loops will have winding number 1 around $w$. On the other hand, for loops sufficiently close to $x$, its image under $\phi_{0}$ has winding number 0 around $w$ since $w\notin\phi_{0}(M(\mathcal{H}))$. This is a contradiction, which implies that almost surely for large enough $k$, $\widetilde{S}_{k}\subseteq\phi_{0}(\mathcal{H})$, which completes the proof. ∎
We call a closed set $D\subseteq M(\mathcal{H})$ a polygon if $\partial D$ is a simple closed curve and can be written as a union of a finite number of line segments on faces of $M(\mathcal{H})$.
Lemma 4.16.
Almost surely, we have
$\{\widetilde{c}(H):H\in\mathcal{H}\}\cap\bigg(\bigcup_{S\in\mathcal{D}}\partial S\bigg)=\emptyset,\qquad\{\widetilde{c}(H):H\in\mathcal{H}\}\cap\bigg(\bigcup_{H_{1},H_{2}\in\mathcal{H},H_{1}\neq H_{2}}(H_{1}\cap H_{2})\bigg)=\emptyset.$ (4.32)
Furthermore, for each $S\in\mathcal{D}$, $\phi_{0}^{-1}(S)$ can be written as a union of a finite number of polygons on $M(\mathcal{H})$. For $S_{1},S_{2}\in\mathcal{D}$ with disjoint interiors, $\phi_{0}^{-1}(S_{1})\cap\phi_{0}^{-1}(S_{2})$ has zero area on $M(\mathcal{H})$. Similarly, for distinct $H_{1},H_{2}\in\mathcal{H}$, $\phi_{0}^{-1}(H_{1})\cap\phi_{0}^{-1}(H_{2})$ has zero area on $M(\mathcal{H})$.
Proof.
The first claim is straightforward since the sets $\partial S$ (for $S\in\mathcal{D}$) and $H_{1}\cap H_{2}$ (for $H_{1},H_{2}\in\mathcal{H}$ distinct) in (4.32) have zero Lebesgue measure and since the points $\widetilde{c}(H)$ are sampled uniformly from sets of positive Lebesgue measure conditioned on the former sets. The other claims follow from (4.32), together with the fact that almost surely, for each $H_{1},H_{2},H_{3}$, the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2})$ and $\widetilde{c}(H_{3})$ is nondeneragate. ∎
Lemma 4.17.
For each $n\geq m\geq 1$, we have
$\mathrm{Energy}(\phi_{m};\phi_{0}^{-1}(\widehat{S}_{n}))\leq\mathrm{Energy}(\phi_{0};\phi_{0}^{-1}(\widehat{S}_{n})).$ (4.33)
Furthermore, there is a deterministic constant $C>0$ such that a.s.
$\limsup_{k\to\infty}\frac{1}{|\widehat{S}_{k}|^{2}}\mathrm{Energy}(\phi_{0};\phi_{0}^{-1}(\widehat{S}_{k}))<C.$ (4.34)
Proof.
The relation (4.33) is an immediate consequence of the fact that harmonic functions minimize Dirichlet energy subject to specified boundary data. Suppose $H_{1},H_{2},H_{3}\in\mathcal{H}$ are adjacent such that the triangle $\Delta_{H_{1},H_{2},H_{3}}^{\prime}$ formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ intersects $\widehat{S}_{k}$. Let $\widehat{S}_{k}^{\prime}$ be the square concentric to $\widehat{S}_{k}$ with side length $1.5|\widehat{S}_{k}|$. Then by Lemma 4.14, $H_{1},H_{2},H_{3}\subseteq\widehat{S}_{k}^{\prime}$ for sufficiently large $k$. Furthermore, if we write $\Delta_{H_{1},H_{2},H_{3}}$ for the triangle in $M(\mathcal{H})$, by Lemma 4.13,
$\begin{split}\mathrm{Energy}(\phi_{0};\Delta_{H_{1},H_{2},H_{3}})&\leq 3\big((\mathrm{diam}(H_{1})+\mathrm{diam}(H_{2}))^{2}+(\mathrm{diam}(H_{1})+\mathrm{diam}(H_{3}))^{2}+(\mathrm{diam}(H_{2})+\mathrm{diam}(H_{3}))^{2}\big)\\ &\leq 12(\mathrm{diam}(H_{1})^{2}+\mathrm{diam}(H_{2})^{2}+\mathrm{diam}(H_{3})^{2}).\end{split}$ (4.35)
Now we sum  (4.35) over all possible choices of $H_{1},H_{2},H_{3}\subseteq\widehat{S}_{k}^{\prime}$ and note that each term $\mathrm{diam}(H)^{2}$ appears at most ${12}\deg(H)$ times in the summation, which gives
$\mathrm{Energy}(\phi_{0};\phi_{0}^{-1}(\widehat{S}_{k}))\leq 12\sum_{H\in\mathcal{H}(\widehat{S}_{k}^{\prime})}\mathrm{diam}(H)^{2}\deg(H).$
The relation (4.34) then follows from Lemma 2.7. ∎
Following the notations of [GMS22], for $z\in\mathbb{C}$, we write $H_{z}$ for the a.s. unique cell in $\mathcal{H}$ containing $z$, and for $m_{1}<m_{2}$, we set
$\mathrm{SE}_{m_{1},m_{2}}^{z}=\frac{\int_{\phi_{0}^{-1}(H_{z}\cap\widehat{S}_{m_{2}}^{z})}|\nabla\phi_{m_{2}}(w)-\nabla\phi_{m_{1}}(w)|^{2}\,dw}{\mathrm{area}(H_{z}\cap\widehat{S}_{m_{2}}^{z})},$ (4.36)
and we write $\mathrm{SE}_{m_{1},m_{2}}$ for $\mathrm{SE}_{m_{1},m_{2}}^{0}$.
Lemma 4.18.
For each fixed $m_{1}<m_{2}$, almost surely
$\lim_{n\to\infty}\frac{1}{|\widehat{S}_{n}|^{2}}\mathbb{E}\Big[\mathrm{Energy}(\phi_{m_{2}}-\phi_{m_{1}};\phi_{0}^{-1}(\widehat{S}_{n}))\big|\mathcal{H},\mathcal{D}\Big]=\mathbb{E}[\mathrm{SE}_{m_{1},m_{2}}],$ (4.37)
and this expectation is bounded above by a finite constant which does not depend on $m_{1},m_{2}$.
Proof.
Observe that for $n\geq m_{2}$, Lemma 4.16 gives that almost surely,
$\frac{1}{|\widehat{S}_{n}|^{2}}\int_{\widehat{S}_{n}}\mathrm{SE}_{m_{1},m_{2}}^{z}\,dz=\frac{1}{|\widehat{S}_{n}|^{2}}\int_{\phi_{0}^{-1}(\widehat{S}_{n})}|\nabla\phi_{m_{2}}(w)-\nabla\phi_{m_{1}}(w)|^{2}\,dw.$ (4.38)
If we dilate the cell configuration $\mathcal{H}$ and the dyadic system $\mathcal{D}$ by some constant $C$, then $\phi_{0}^{-1}(H_{0}\cap\widehat{S}_{m_{2}})$ is invariant, $\phi_{m}$ is multiplied by $C$, and $\mathrm{area}(H_{0}\cap\widehat{S}_{m_{2}})$ is multiplied by $C^{2}$, and thus $\mathbb{E}[\mathrm{SE}_{m_{1},m_{2}}|\mathcal{H},\mathcal{D}]$ is scale invariant. By Lemma 4.17 and a triangle inequality, there is a deterministic constant $C_{0}$ such that  (4.38) is less than $C_{0}$ for all $n$ large enough. The lemma then follows by taking conditional expectation of (4.38) over $(\mathcal{H},\mathcal{D})$ and setting $F(\mathcal{H},\mathcal{D})=\mathbb{E}[\mathrm{SE}_{m_{1},m_{2}}|\mathcal{H},\mathcal{D}]$ in Lemma 2.4 (observe that $F(\mathcal{H}-z,\mathcal{D}-z)=\mathbb{E}[\mathrm{SE}_{m_{1},m_{2}}^{z}|\mathcal{H},\mathcal{D}]$ for each $z\in\mathbb{C}$ since given $\mathcal{H},\mathcal{D}$, up to translation, the $\phi_{m}$-functions for $(\mathcal{H},\mathcal{D})$ and $(\mathcal{H}-z,\mathcal{D}-z)$ have the same law). ∎
Lemma 4.19.
For each $\varepsilon>0$, there exists $m_{\varepsilon}\in\mathbb{N}$ such that
$\mathbb{E}[\mathrm{SE}_{m_{\varepsilon},m}]<\varepsilon,\ \text{ for all }m\geq m_{\varepsilon}.$ (4.39)
Proof.
For each square $S\in\mathfrak{S}_{m}$, and $m^{\prime}\leq m$, $\phi_{m^{\prime}}-\phi_{m^{\prime}-1}$ vanishes on $\partial\phi_{0}^{-1}(S)$ following the definition of $\phi_{m}$, and $\phi_{m}$ is harmonic in the interior of $\phi_{0}^{-1}(S)$. Thus using integration by parts, almost surely, for each $S\in\mathfrak{S}_{m}$,
$\int_{\phi_{0}^{-1}(S)}\nabla\phi_{m}(w)\cdot(\nabla\phi_{m^{\prime}}(w)-\nabla\phi_{m^{\prime}-1}(w))\,dw=0.$
By a summation over the squares, one can check that for $n\geq m\geq m^{\prime}$, almost surely
$\frac{1}{|\widehat{S}_{n}|^{2}}\mathrm{Energy}(\phi_{m}-\phi_{m^{\prime}};\phi_{0}^{-1}(\widehat{S}_{n}))=\sum_{j=m^{\prime}+1}^{m}\frac{1}{|\widehat{S}_{n}|^{2}}\mathrm{Energy}(\phi_{j}-\phi_{j-1};\phi_{0}^{-1}(\widehat{S}_{n})).$
Now taking $n\to\infty$, by Lemma 4.18, we have
$\mathbb{E}[\mathrm{SE}_{m^{\prime},m}]=\sum_{j=m^{\prime}+1}^{m}\mathbb{E}[\mathrm{SE}_{j-1,j}],$
and the sum is uniformly bounded from above by the last part of Lemma 4.18. In particular, there exists $m_{\varepsilon}$ such that $\sum_{j=m_{\varepsilon}}^{\infty}\mathbb{E}[\mathrm{SE}_{j-1,j}]<\varepsilon$, and thus $\mathbb{E}[\mathrm{SE}_{m_{\varepsilon},m}]<\varepsilon$ for all $m\geq m_{\varepsilon}$. ∎
Proof of Proposition 4.12.
Fix $\ell\in\mathbb{N}$. Let $\varepsilon>0$ and $m_{\varepsilon}$ be as in Lemma 4.19. Without loss of generality assume $\ell<m_{\varepsilon}$. Then almost surely, using Lemma 4.16, for $m\geq m_{\varepsilon}$,
$\begin{split}\int_{\widehat{S}_{\ell}}&\mathrm{SE}_{m_{\varepsilon},m}^{z}\,dz=\sum_{H\in\mathcal{H}(\widehat{S}_{\ell})}\frac{\mathrm{area}(H\cap\widehat{S}_{\ell})}{\mathrm{area}(H\cap\widehat{S}_{m})}\int_{\phi_{0}^{-1}(H\cap\widehat{S}_{m})}|\nabla\phi_{m}(w)-\nabla\phi_{m_{\varepsilon}}(w)|^{2}\,dw\\ &\geq\sum_{H\in\mathcal{H}(\widehat{S}_{\ell})\backslash\mathcal{H}(\partial\widehat{S}_{\ell})}\int_{\phi_{0}^{-1}(H)}|\nabla\phi_{m}(w)-\nabla\phi_{m_{\varepsilon}}(w)|^{2}\,dw=\int_{\phi_{0}^{-1}(\mathcal{H}(\widehat{S}_{\ell})\backslash\mathcal{H}(\partial\widehat{S}_{\ell}))}|\nabla\phi_{m}(w)-\nabla\phi_{m_{\varepsilon}}(w)|^{2}\,dw.\end{split}$ (4.40)
On the other hand, by Lemma 2.3 and the definition of $m_{\varepsilon}$,
$\mathbb{E}\Big[\frac{1}{|\widehat{S}_{\ell}|^{2}}\int_{\widehat{S}_{\ell}}\mathrm{SE}_{m_{\varepsilon},m}^{z}\,dz\Big]=\mathbb{E}[\mathrm{SE}_{m_{\varepsilon},m}]<\varepsilon.$ (4.41)
Let $\varphi:M(\mathcal{H})\to\mathbb{C}$ be a conformal map sending the point 0 on $M(\mathcal{H})$ (which corresponds to the a.s. unique origin-containing cell of $\mathcal{H}$) to 0. Using the conformal invariance of Dirichlet energy, for each $m\geq m_{\varepsilon}$,
$\mathbb{E}\Big[\int_{{\widetilde{D}_{\ell}}}|\nabla(\phi_{m}\circ\varphi^{-1})(z)-\nabla(\phi_{m_{\varepsilon}}\circ\varphi^{-1})(z)|^{2}\,dz\Big]<\varepsilon,$ (4.42)
[…]
$\begin{split}\big|f(z{{}_{0}})-f(0)\big|&=\Big|\int_{0}^{r_{0}}\bigg(\frac{\partial f}{\partial x}(re^{i\theta_{0}})\cos\theta_{0}+\frac{\partial f}{\partial y}(re^{i\theta_{0}})\sin\theta_{0}\bigg)\,{dr}\Big|\\ &=\frac{1}{\pi\delta^{2}}\Big|\int_{0}^{r_{0}}\int_{\delta\mathbb{D}}\big(\frac{\partial f}{\partial x}(re^{i\theta_{0}}+w)\cos\theta_{0}+\frac{\partial f}{\partial y}(re^{i\theta_{0}}+w)\sin\theta_{0}\big)\,dw{\,dr}\Big|\\ &\leq\frac{10}{\delta^{2}}\int_{{\widetilde{D}}_{\ell}}|\nabla f(w)|\,dw\leq\frac{10}{\delta^{2}}\,\mathrm{area}({\widetilde{D}}_{\ell})^{1/2}\big(\int_{{\widetilde{D}}_{\ell}}|\nabla f(w)|^{2}\,dw\big)^{1/2}.\end{split}$ (4.43)
Combined with (4.42) we see that the sequence $\{\phi_{m}\circ\varphi^{-1}-\phi_{m}(0)\}$ is Cauchy in probability with respect to the topology of uniform convergence on $\mathring{D}_{\ell}^{\delta}$. Since $\mathcal{H}$ is locally finite, by varying $\ell$ and $\delta$, one can find a function $\phi_{\infty}$ on $M(\mathcal{H})$ such that $\{\phi_{m}\circ\varphi^{-1}-\phi_{m}(0)\}$ converges to $\phi_{\infty}\circ\varphi^{-1}$ in probability with respect to the topology of uniform convergence on compact subsets of $\varphi(M(\mathcal{H}))$. This verifies (4.31). The function $\phi_{\infty}$ is a.s. harmonic since each $\phi_{m}$ is harmonic on $\phi_{0}^{-1}(\mathcal{H}(\widehat{S}_{m})\backslash\mathcal{H}(\partial S_{m}))$, and the uniform limit of harmonic functions is harmonic. The claim $\mathbb{E}\big[\phi_{\infty}(x)|\mathcal{H},\mathcal{D}\big]<\infty$ for each $x\in M(\mathcal{H})$ also follows from (4.42) and the same argument as above. ∎
4.3 Sublinearity of the corrector
The aim of this subsection is to prove Proposition 4.20 below. Recall the notation $M(S)$ introduced at the beginning of Section 4.1. We start with the fact that the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ is small when $m$ is large (Lemma 4.22), and for fixed $m$, the maximal value of $|\phi_{m}(z)-\phi_{0}(z)|$ on $M(S_{k})$ has order $o_{k}(|S_{k}|)$ as $k\to\infty$ (Lemma 4.23). Then we transfer the bound on the the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ to a pointwise bound for $|\phi_{0}-\phi_{\infty}|$. The structure of the proof is similar as in [GMS22, Section 2.4], but we need to use a different setup when we conclude the proof of Proposition 4.20 where we we work with the uniformization map $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ as in Lemmas 4.3 and 4.4, and we also need to use regularity results from Section 4.1. We refer to the paragraph above the latter proof for a proof outline.
Proposition 4.20.
Almost surely, in the setting of Proposition 4.12,
$\lim_{k\to\infty}\frac{1}{2^{k}}\sup_{x\in M([-2^{k},2^{k}]^{2})}|\phi_{0}(x)-\mathbb{E}[\phi_{\infty}(x)|\mathcal{H},\mathcal{D}]|=0.$ (4.44)
Extending the definition (4.36), we set
$\mathrm{SE}_{m,\infty}^{z}=\frac{\int_{\phi_{0}^{-1}(H_{z})}|\nabla\phi_{m}(x)-\nabla\phi_{\infty}(x)|^{2}\,dx}{\mathrm{area}(H_{z})},$ (4.45)
and define $\mathrm{SE}_{m,\infty}=\mathrm{SE}_{m,\infty}^{0}$.
Lemma 4.21.
For each $m\in\mathbb{N}$, $\mathbb{E}[\mathrm{SE}_{m,\infty}]<\infty$, and
$\lim_{m\to\infty}\mathbb{E}[\mathrm{SE}_{m,\infty}]=0.$ (4.46)
Proof.
For each fixed $m$, by Proposition 4.12, along with the fact that the convergence of harmonic functions $\{f_{n}\}$ to $f$ in local uniform topology implies the convergence of $\{\nabla f_{n}\}$ to $\nabla f$ in local uniform topology, $\mathrm{SE}_{m,m^{\prime}}\to\mathrm{SE}_{m,\infty}$ in probability as $m^{\prime}\to\infty$. By the last sentence of Lemma 4.18 and Fatou’s lemma, we have $\mathbb{E}[\mathrm{SE}_{m,\infty}]<\infty$. By Fatou’s lemma and Lemma 4.19, we have $\mathbb{E}[\mathrm{SE}_{m,\infty}]\leq\varepsilon$ for $m\geq m_{\varepsilon}$, which further implies (4.46). ∎
Lemma 4.22.
Almost surely,
$\lim_{m\to\infty}\limsup_{n\to\infty}\frac{1}{|\widehat{S}_{n}|^{2}}\mathbb{E}\big[\mathrm{Energy}(\phi_{\infty}-\phi_{m};M(\widehat{S}_{n}))|\mathcal{H},\mathcal{D}\big]=0.$ (4.47)
Proof.
By Lemma 4.16,
$\begin{split}\frac{1}{|\widehat{S}_{n}|^{2}}\int_{\widehat{S}_{n}}\mathrm{SE}_{m,\infty}^{z}\,dz&=\frac{1}{|\widehat{S}_{n}|^{2}}\sum_{H\in\mathcal{H}(\widehat{S}_{n})}\frac{\mathrm{area}(H\cap\widehat{S}_{n})}{\mathrm{area}(H)}\int_{\phi_{0}^{-1}(H)}|\nabla\phi_{m}(w)-\nabla\phi_{\infty}(w)|^{2}\,dw\\ \geq\frac{1}{|\widehat{S}_{n}|^{2}}&\int_{\phi_{0}^{-1}(\mathcal{H}(\widehat{S}_{n})\backslash\mathcal{H}(\partial\widehat{S}_{n}))}|\nabla\phi_{m}(w)-\nabla\phi_{\infty}(w)|^{2}\,dw\geq\frac{1}{|\widehat{S}_{n}|^{2}}\mathrm{Energy}(\phi_{\infty}-\phi_{m};M(\widehat{S}_{n})),\end{split}$
where we refer to the beginning of Section 4.1 for the definition of $M(\widehat{S}_{n})$ and where, in the last inequality, we used the fact that if $H_{1},H_{2},H_{3}$ form a triangular face on $M(\widehat{S}_{n})$, then the cells $H_{1},H_{2},H_{3}$ are within the interior of $\widehat{S}_{n}$. Now we set $F(\mathcal{H},\mathcal{D})=\mathbb{E}[\mathrm{SE}_{m,\infty}|\mathcal{H},\mathcal{D}]$. Then using the definition of $\phi_{m}$ and $\phi_{\infty}$, $F(\mathcal{H}-z,\mathcal{D}-z)=\mathbb{E}[\mathrm{SE}_{m,\infty}^{z}|\mathcal{H},\mathcal{D}]$, and $F(C\mathcal{H},C\mathcal{D})=F(\mathcal{H},\mathcal{D})$. Therefore the conclusion follows from Lemma 2.4 and Lemma 4.21. ∎
Lemma 4.23.
For each fixed $m\in\mathbb{N}$, almost surely,
$\lim_{k\to\infty}\frac{1}{|S_{k}|}\sup_{x\in M(S_{k})}|\phi_{m}(x)-\phi_{0}(x)|=0.$ (4.48)
Proof.
By [GMS22, Equation (2.42)] (where we take the conductance $\mathfrak{c}(H,H^{\prime})$ there to be the number of edges between $H$ and $H^{\prime}$, which guarantees that their moment bound is satisfied (3.8)), a.s.
$\lim_{k\to\infty}\frac{1}{|S_{k}|}\max_{\widehat{S}\subseteq S_{k},\widehat{S}\in\mathfrak{S}_{m}}\mathrm{diam}(\widehat{S})=0.$ (4.49)
[…]
We are now ready to prove Proposition 4.20 by transferring the bound on the Dirichlet energy from Lemma 4.22 to a pointwise bound for $|\phi_{0}-\phi_{\infty}|$. The bound on the Dirichlet energy for $\phi_{m}-\phi_{\infty}$ will allow us to control $|\phi_{0}\circ\varphi_{k}^{-1}-\phi_{\infty}\circ\varphi_{k}^{-1}|$ on a dense collection $\mathcal{P}_{k}\subseteq|S_{k}|\mathbb{D}$ of horizontal and vertical line segments. The oscillation of $\phi_{\infty}\circ\varphi_{k}^{-1}$ on $\frac{1}{2}|S_{k}|\mathbb{D}\backslash\mathcal{P}_{k}$ is controlled by the maximum principle for harmonic functions, while we use the continuity result from Lemma 4.4 to bound the oscillation of $\phi_{0}\circ\varphi_{k}^{-1}$ on $\frac{1}{2}|S_{k}|\mathbb{D}\backslash\mathcal{P}_{k}$. Combining these with Lemma 4.3, for some deterministic constant $c_{2}>0$, the oscillation of $|\phi_{0}-\phi_{\infty}|$ on $M(\mathring{S}_{k}^{c_{2}|S_{k}|})$ has order $o_{k}(|S_{k}|)$, and the bound can be further extended to $M([-2^{k},2^{k}]^{2})$ via Lemma 2.5. The first half of the proof is similar as [GMS22], except that we work on continuum Dirichlet energy rather than discrete Dirichlet energy and that the lines $\ell_{j},\widetilde{\ell}_{j}$ are contained in the image of certain conformal maps $\varphi_{k}$ instead of living in the same plane as the cell system. The second half is more different, where the function $\phi_{0}$ in [GMS22] is roughly identity, while we need to use Lemma 4.3 to control the oscillation of $\phi_{0}$ in our setting.
Proof of Proposition 4.20.
Fix $\varepsilon\in(0,10^{-6})$. By Lemma 4.22, we may pick $m_{\varepsilon}>0$ such that
$\limsup_{k\to\infty}\frac{1}{|S_{k}|^{2}}\mathbb{E}\big[\mathrm{Energy}(\phi_{\infty}-\phi_{m_{\varepsilon}};M(S_{k}))|\mathcal{H},\mathcal{D}\big]<e^{-1/\varepsilon^{8}}.$
Let $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ be the uniformization map defined as in Lemma 4.3 and Lemma $4.$ Using the conformal invariance of Dirichlet energy,
$\limsup_{k\to\infty}\frac{1}{|S_{k}|^{2}}\int_{|S_{k}|\mathbb{D}}\mathbb{E}\Big[|\nabla(\phi_{\infty}\circ\varphi_{k}^{-1})(w)-\nabla(\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1})(w)|^{2}\big|\mathcal{H},\mathcal{D}\Big]\,dw<e^{-1/\varepsilon^{8}}.$ (4.50)
From (4.50), for large enough $k$, there exists a collection of horizontal straight lines $\ell_{1},...,\ell_{n}$ such that for each j=1,,nj=1,...,n,
$\frac{1}{|S_{k}|}\int_{\ell_{j}\cap|S_{k}|\mathbb{D}}\mathbb{E}\Big[|\nabla(\phi_{\infty}\circ\varphi_{k}^{-1})(y)-\nabla(\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1})(y)|^{2}\big|\mathcal{H},\mathcal{D}\big]\,dy<\varepsilon^{8},$ (4.51)
and for j=0,1,,nj=0,1,...,n, the distance between $\ell_{j}$ and $\ell_{j+1}$ is less than $e^{-1/\varepsilon^{4}}|S_{k}|$, with $\ell_{0}=\{z:\mathrm{Im}\,z=|S_{k}|\}$ and $\ell_{n+1}=\{z:\mathrm{Im}\,z=-|S_{k}|\}$. Similarly, one can find a collection of vertical straight lines $\widetilde{\ell}_{1},...,\widetilde{\ell}_{\widetilde{n}}$ such that for each j=1,,nj=1,...,n,  (4.51) holds for $\widetilde{\ell}_{j}$ instead of $\ell_{j}$, and the distance $\widetilde{\ell}_{j}$ and $\widetilde{\ell}_{j+1}$ is less than $e^{-1/\varepsilon^{4}}|S_{k}|$, with $\widetilde{\ell}_{0}=\{z:\mathrm{Re}\,z=|S_{k}|\}$ and $\widetilde{\ell}_{\widetilde{n}+1}=\{z:\mathrm{Re}\,z=-|S_{k}|\}$. We assume that these lines are selected in some way that is measurable with respect to $(\mathcal{H},\mathcal{D})$. By the Cauchy-Schwartz inequality, for each $\ell\in\{\ell_{1},...,\ell_{n},\widetilde{\ell}_{1},...,\widetilde{\ell}_{\widetilde{n}}\}$,
$\frac{1}{|S_{k}|}\int_{\ell\cap|S_{k}|\mathbb{D}}\mathbb{E}\Big[|\nabla(\phi_{\infty}\circ\varphi_{k}^{-1})(y)-\nabla(\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1})(y)|\big|\mathcal{H},\mathcal{D}\big]\,dy<4\varepsilon^{4}.$ (4.52)
Integrating this, for each $w,w^{\prime}\in\ell\cap|S_{k}|\mathbb{D}$,
$\begin{split}\frac{1}{|S_{k}|}\mathbb{E}\Big[\big|\big(&\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1}(w)\big)-\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})-\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1}(w^{\prime})\big)\big|\Big|\mathcal{H},\mathcal{D}\Big]<4\varepsilon^{4}.\end{split}$
It follows that if we set $\mathcal{P}_{k}=(\ell_{1}\cup...\cup\ell_{n}\cup\widetilde{\ell}_{1}\cup...\cup\widetilde{\ell}_{\widetilde{n}})\cap|S_{k}|\mathbb{D}$, then for large enough $k$,
$\frac{1}{|S_{k}|}\sup_{w,w^{\prime}\in\mathcal{P}_{k}}\mathbb{E}\Big[\big|\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1}(w)\big)-\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})-\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1}(w^{\prime})\big)\big|\Big|\mathcal{H},\mathcal{D}\Big]<8\varepsilon^{4}<\varepsilon^{2},$ (4.53)
and by (4.48), for large enough $k$,
$\frac{1}{|S_{k}|}\sup_{w,w^{\prime}\in\mathcal{P}_{k}}\mathbb{E}\Big[\big|\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{0}\circ\varphi_{k}^{-1}(w)\big)-\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})-\phi_{0}\circ\varphi_{k}^{-1}(w^{\prime})\big)\big|\Big|\mathcal{H},\mathcal{D}\Big]<4\varepsilon^{2}.$ (4.54)
Consider a connected component $K$ of $\big(\frac{|S_{k}|}{2}\mathbb{D}\big)\backslash\mathcal{P}_{k}$. Then $\mathrm{diam}(K)<{\sqrt{2}}e^{-1/\varepsilon^{4}}|S_{k}|$, and by Lemma 4.4, for some constant $c_{1}$, for $k$ large enough and all choices of $K$, we have
$\mathrm{diam}\bigg(\bigcup_{H:\varphi_{k}(P_{H})\cap K\neq\emptyset}H\bigg)\leq c_{1}\varepsilon^{2}|S_{k}|.$ (4.55)
In particular, for each $w,w^{\prime}\in K$, (4.55) along with Lemma 2.6 implies that a.s. for $k$ large enough,
$\sup_{K}\sup_{w,w^{\prime}\in K}|\phi_{0}\circ\varphi_{k}^{-1}(w)-\phi_{0}\circ\varphi_{k}^{-1}(w^{\prime})|<4c_{1}\varepsilon^{2}|S_{k}|.$ (4.56)
Using the maximum principle for the harmonic function $\phi_{\infty}\circ\varphi_{k}^{-1}$, a.s. for large enough $k$, for each $K$, by (4.54) and (4.56), we have
$\begin{split}&\sup_{w,w^{\prime}\in K}\mathbb{E}\Big[|\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})|\Big|\mathcal{H},\mathcal{D}\Big]=\sup_{w,w^{\prime}\in\partial K}\mathbb{E}\Big[|\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})|\Big|\mathcal{H},\mathcal{D}\Big]\\ &\leq\sup_{w,w^{\prime}\in\partial K}\mathbb{E}\Big[|(\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime}))-(\phi_{0}\circ\varphi_{k}^{-1}(w)-\phi_{0}\circ\varphi_{k}^{-1}(w^{\prime}))|\Big|\mathcal{H},\mathcal{D}\Big]\\ &\qquad+\sup_{w,w^{\prime}\in\partial K}\mathbb{E}\Big[|\phi_{0}\circ\varphi_{k}^{-1}(w)-\phi_{0}\circ\varphi_{k}^{-1}(w^{\prime})|\Big|\mathcal{H},\mathcal{D}\Big]\leq{4(1+c_{1})\varepsilon^{2}|S_{k}|}.\end{split}$ (4.57)
We conclude by (4.54), (4.56),  (4.57) along with a triangle inequality that a.s. for $k$ large enough,
$\frac{1}{|S_{k}|}\sup_{w,w^{\prime}\in\frac{|S_{k}|}{2}\mathbb{D}}\Bigg|\mathbb{E}\Big[\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w)-\phi_{0}\circ\varphi_{k}^{-1}(w)\big)-\big(\phi_{\infty}\circ\varphi_{k}^{-1}(w^{\prime})-\phi_{0}\circ\varphi_{k}^{-1}(w^{\prime})\big)\Big|\mathcal{H},\mathcal{D}\Big]\Bigg|<32(1+c_{1})\varepsilon^{2}.$ (4.58)
By Lemma 4.3, there is some deterministic constant $c_{2}>0$ such that a.s. for $k$ large enough, if $H\cap\mathring{S}_{k}^{c_{2}|S_{k}|}\neq\emptyset$, then $\varphi_{k}(P_{H})\subseteq\frac{|S_{k}|}{2}\mathbb{D}$. Therefore it follows from (4.58) that, a.s. for large enough $k$, if we let $\widetilde{M}(\mathring{S}_{k}^{c_{2}|S_{k}|})$ be the union of the faces of $M(\mathcal{H})$ formed by $H_{1},H_{2},H_{3}$ with $H_{1},H_{2},H_{3}\subseteq\mathring{S}_{k}^{c_{2}|S_{k}|}$, then
$\frac{1}{|S_{k}|}\sup_{x,x^{\prime}\in\widetilde{M}(\mathring{S}_{k}^{c_{2}|S_{k}|})}\Bigg|\mathbb{E}\Big[\big(\phi_{\infty}(x)-\phi_{0}(x)\big)-\big(\phi_{\infty}(x^{\prime})-\phi_{0}(x^{\prime})\big)\Big|\mathcal{H},\mathcal{D}\Big]\Bigg|<32(1+c_{1})\varepsilon^{2}.$ (4.59)
We pick $\ell\in\mathbb{N}$ such that $c_{2}2^{\ell}>16$. Applying Lemma 2.5 to the statement (4.59), a.s. for large enough $k$,  (4.59) holds with $S_{k}$ replaced by any of its dyadic ancestor of side length $2^{\ell}|S_{k}|$, and thus
$\frac{1}{|S_{k}|}\sup_{x,x^{\prime}\in M(\mathring{S}_{k}^{4|S_{k}|})}\Bigg|\mathbb{E}\Big[\big(\phi_{\infty}(x)-\phi_{0}(x)\big)-\big(\phi_{\infty}(x^{\prime})-\phi_{0}(x^{\prime})\big)\Big|\mathcal{H},\mathcal{D}\Big]\Bigg|<32(1+c_{1})2^{\ell}\varepsilon^{2}.$ (4.60)
Now using Lemma 4.1 along with the fact that $2^{k},2^{k}]^{2}\subset\mathring{S}_{k}^{2|S_{k}|(1-\varepsilon)}$ a.s. for large $k$ and some random $\varepsilon>0$, $M([-2^{k},2^{k}]^{2})\subseteq M(\mathring{S}_{k}^{4|S_{k}|})$ a.s. for large $k$. Furthermore, $0\in M(\mathring{S}_{k}^{4|S_{k}|})$ a.s. for large $k$. Since $\phi_{\infty}(0)=0$,  (4.60) implies a.s. for large $k$,
$\frac{1}{2^{k}}\sup_{x\in M([-2^{k},2^{k}]^{2})}\Bigg|\mathbb{E}\Big[\big(\phi_{\infty}(x)-\phi_{0}(x)\big)\Big|\mathcal{H},\mathcal{D}\Big]\Bigg|<(1+c_{1})2^{\ell+10}\varepsilon^{2}.$ (4.61)
Finally, Lemma 2.6 implies that almost surely,
$\lim_{k\to\infty}\frac{1}{2^{k}}\sup_{x\in M([-2^{k},2^{k}]^{2})}\big|\phi_{0}(x)-\mathbb{E}[\phi_{0}(x)|\mathcal{H},\mathcal{D}]\big|=0.$
Combining with (4.61) gives  (4.44). ∎
4.4 Proof of Theorem 1.8
In this section we complete the proof of Theorem 1.8. Recall that by Lemma 4.9, the surface $M(\mathcal{H})$ is a.s. parabolic, and by Proposition 4.20, $\mathbb{E}[\phi_{\infty}|\mathcal{H},\mathcal{D}]$ is close to $\phi_{0}$ on a large scale. Let $\varphi$ be a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ chosen in a way measurable with respect to $\mathcal{H}$ such that $\varphi(0)=0$. We will prove that the harmonic function $\widetilde{\phi}_{\infty}$$\mathbb{C}\to\mathbb{C}$ defined by $\widetilde{\phi}_{\infty}(z):=\mathbb{E}[\phi_{\infty}\circ\varphi^{-1}(z)|\mathcal{H},\mathcal{D}]$ is the composition of a deterministic linear transform $\mathbf{A}$, a rotation and a scaling, which implies that the embedding induced by $\varphi(M(\mathcal{H}))$ is close to $\mathcal{H}$ up to a rotation. This finishes the proof of Theorem 1.8 when combined with Proposition 4.20.
We prove the decomposition of $\widetilde{\phi}_{\infty}$ as follows. First we prove that the harmonic function $\widetilde{\phi}_{\infty}$ must be a polynomial (Lemma 4.24). This follows from Proposition 4.20 as well as the polynomial growth property in Lemma 4.10. Then in Lemmas 4.254.26 and 4.27, we prove that this polynomial must have degree 1, by arguing that if this is not the case, then one can find cells $H_{1},H_{2}$ which are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.3 and Lemma 4.6. Next we use the ergodic theorem for cell configurations (Lemma 2.4) to prove that for $a,b\in\mathbb{R}$, the average Dirichlet energy of $a\,\mathrm{Re}\,\phi_{\infty}+b\,\mathrm{Im}\,\phi_{\infty}$ is deterministic (Lemma 4.29). This implies that for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, if we view $\widetilde{\phi}_{\infty}$ as a $2\times 2$ matrix, then $\langle{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{u}},{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{v}}\rangle$ is deterministic up to constant (Lemma 4.30), and the desired decomposition is a consequence of standard linear algebra (Lemma 4.31). Finally in Lemma 4.33, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.6, then $\mathbf{A}$ must be a rotation.
Before giving the details of this proof, let us remark that our approach in this section is completely different from the counterpart of Theorem 1.8 for the Tutte embedding in [GMS22, GMS21] and circle packings in our Section 3. In the setting of Tutte embedding, there is no well-defined canonical Tutte embedding for infinite maps; instead an invariance principle is proven in [GMS22], where the simple random walk on $\mathcal{H}$ a.s. converges to the Brownian motion (Theorem 3.5). In the finite volume setting convergence of the conformal embedding is an immediate consequence of the invariance principle by the definition of the Tutte embedding [GMS21]. In the setting of circle packings, we proved in Section 3.2 and Section 3.3 that the random walk on $\mathcal{H}$ and the circle packing for $\mathcal{H}$ with Dubejko weights a.s. converges to the Brownian motion, and in Section 3.4 we used this to conclude that the cell configuration $\mathcal{H}$ is close to its circle packing.
We remark that it might be possible to adapt the arguments in [GMS22, Section 3] to show that for a Brownian motion $(W_{t})_{t\geq 0}$ on $M(\mathcal{H})$, $(\phi_{0}(W_{t}))_{t\geq 0}$ is approximately a Brownian motion on $\mathbb{C}$, and, combined with the same arguments from Section 3.4, we would conclude that the a priori embedding $\phi_{0}$ of $M(\mathcal{H})$ is close to a conformal map. On the other hand, our proof in this section is relatively elementary and has independent interest, as it gives another way of proving that harmonic functions are conformal under certain conditions. We emphasize that our argument works under rather mild assumptions; the key inputs are polynomial growth, almost injectivity, and that the average Dirichlet energy over large regions is approximately constant.
Lemma 4.24.
There exists a deterministic constant $\beta>0$, such that almost surely, for some random constant $\widetilde{C}>0$ and every large enough $r$, $|\widetilde{\phi}_{\infty}(z)|\leq 4\widetilde{C}r^{\beta}$ for every $z\in{B(0;r)}$. Furthermore, the function $\widetilde{\phi}_{\infty}$ is a polynomial over $z$ and $\overline{z}$ of degree at most $\beta$.
Proof.
Let $\widetilde{C}$ and $\beta$ be the constants from Lemma 4.10, and assume $r$ is large such that (4.30) holds. Let $z\in B(0;r)$, and suppose $\varphi^{-1}(z)\in M(\mathcal{H})$ is on the triangle formed by the cells $H_{1},H_{2},H_{3}$. Then for some $j=\{1,2,3\}$, $H_{j}\subseteq B(0;\widetilde{C}r^{\beta})$, and by Lemma 2.6, $H_{1},H_{2},H_{3}\subseteq B(0;2\widetilde{C}r^{\beta})$, and by Lemma 4.1, ${\varphi^{-1}(z)}\in M([-3\widetilde{C}r^{\beta},3\widetilde{C}r^{\beta}]^{2})$. Therefore by (4.30), $|\phi_{0}\circ\varphi^{-1}(z)|\leq 2\widetilde{C}r^{\beta}$, and by (4.44), for large enough $r$, $|\phi_{0}\circ\varphi^{-1}(z)-\widetilde{\phi}_{\infty}(z)|\leq\widetilde{C}r^{\beta}$. This verifies the first claim. For the second claim, note that $\widetilde{\phi}_{\infty}$ is a harmonic function defined on the whole plane. By standard properties of harmonic functions, for each $n$, there exists absolute constant $C_{n}$ such that for every harmonic function $f$ defined on $\mathbb{C}$,
$\sup_{z\in B(0;r)}|D^{n}f(z)|\leq C_{n}r^{-n}\sup_{z\in B(0;2r)}|f(z)|,$
where $D^{n}f$ is the $n$th derivative of $f$. Therefore the first claim implies that for all $n>\beta$, $D^{n}\widetilde{\phi}_{\infty}=0$, which implies the second claim. ∎
In order to get the next lemma we argue that the lower-order terms of the polynomial $\widetilde{\phi}_{\infty}$ (i.e., those having exponent less than the degree of the polynomial) are of a smaller order than $2^{k}$ when evaluated on $M([-2^{k},2^{k}])$.
Lemma 4.25.
There exists possibly random constants $\beta\in\mathbb{N}$, $a_{\beta},b_{\beta}\in\mathbb{C}$ with $|a_{\beta}|^{2}+|b_{\beta}|^{2}\neq 0$, such that almost surely
$\lim_{k\to\infty}\frac{1}{2^{k}}\sup_{x\in M([-2^{k},2^{k}]^{2})}|\phi_{0}(x)-a_{\beta}\mathrm{Re}(\varphi(x)^{\beta})-b_{\beta}\mathrm{Im}(\varphi(x)^{\beta})\mathrm{|}=0.$ (4.62)
Proof.
By Lemma 4.24, we may assume that $\widetilde{\phi}_{\infty}$ is a polynomial of degree $\beta\in\mathbb{N}\cup\{0\}$. From the definition of the surface $M([-2^{k},2^{k}]^{2})$ and Lemma 4.1, we infer from (4.44) that $\beta\neq 0$. Note that any harmonic polynomial in $z$ and $\overline{z}$ of degree $\beta$ can be written as a linear combination of $\{1,\mathrm{Re}\,z,...,\mathrm{Re}\,z^{\beta},\mathrm{Im}\,z,...,\mathrm{Im}\,z^{\beta}\}$, and we may assume $\widetilde{\phi}_{\infty}(z)=a_{\beta}\mathrm{Re}\,z^{\beta}+b_{\beta}\mathrm{Im}\,z^{\beta}+p_{\infty}(z)$, where $|a_{\beta}|^{2}+|b_{\beta}|^{2}\neq 0$, and $p_{\infty}(z)$ is a polynomial of $z,\overline{z}$ of degree at most $\beta-1$. In particular, there exists some random $\widetilde{C}>0$ such that $|\widetilde{\phi}_{\infty}(z)-a_{\beta}\mathrm{Re}\,z^{\beta}-b_{\beta}\mathrm{Im}\,z^{\beta}|\leq\widetilde{C}|z|^{\beta-1}$ for $|z|$ large enough, and there exists some $\theta_{0}\in[0,2\pi]$ and $\widetilde{c}>0$ such that $|\widetilde{\phi}_{\infty}(re^{i\theta_{0}})|\geq\widetilde{c}r^{\beta}$ for $r$ large enough. Recall the constant $\alpha$ from Lemma 4.6 and the definition of $d_{k}$ from (4.23). By (4.29), for some constant $c_{2}$, $B(0;c_{2}d_{k})\subseteq\varphi(M([-\alpha 2^{k},\alpha 2^{k}]^{2}))$ for all large enough $k$. Then
$\phi_{0}\circ\varphi^{-1}(B(0;c_{2}d_{k}))\subseteq[-\alpha 2^{k},\alpha 2^{k}]^{2}$ (4.63)
from the definition of $M([-\alpha 2^{k},\alpha 2^{k}]^{2})$, and further by (4.44), $|\widetilde{\phi}_{\infty}(z)|\leq 2^{k+1}\alpha$ for $z\in B(0;c_{2}d_{k})$. Therefore a.s. for large enough $k$, we have
$\widetilde{c}(c_{2}d_{k})^{\beta}\leq 2^{k+1}\alpha.$ (4.64)
On the other hand, a.s. for all $k$ sufficiently large, we have
$|\widetilde{\phi}_{\infty}(z)-a_{\beta}\mathrm{Re}\,z^{\beta}-b_{\beta}\mathrm{Im}\,z^{\beta}|\leq\widetilde{C}d_{k}^{\beta-1},\ \ \text{for all }z\in\overline{B(0;d_{k})}.$ (4.65)
Combining (4.64), (4.65) with (4.44), and using $\varphi(M([-\alpha 2^{k},\alpha 2^{k}]^{2}))\subseteq\overline{B(0;d_{k})}$, we conclude that a.s.,
$\lim_{k\to\infty}\frac{1}{2^{k}}\sup_{x\in M([-\alpha 2^{k},\alpha 2^{k}]^{2})}|\phi_{0}(x)-a_{\beta}\mathrm{Re}(\varphi(x)^{\beta})-b_{\beta}\mathrm{Im}(\varphi(x)^{\beta})\mathrm{|}=0,$
which further implies (4.62). ∎
As discussed above, we get that $\beta=1$ since regularity results for $\varphi$ imply that if two cells are close in $\mathcal{H}$ then their images under $\varphi$ are also close.
Lemma 4.26.
In the setting of Lemma 4.25, almost surely $\beta=1$.
Proof.
Let $\alpha$ and $d_{k}$ be given by Lemma 4.6 and (4.23) as in the proof of Lemma 4.25, and $c_{2}$ be the constant in Lemma 4.8. Using identical arguments as in the proof of Lemma 4.3 and Lemma 4.7, one can show that there exists a deterministic constant $c>0$, such that for each $\delta\in(0,0.01)$, a.s. for each $w\in B(0;\alpha 2^{k-1})$ and large $k$,
$\mathrm{diam}\Big(\bigcup_{H\in\mathcal{H}(w+[-\delta 2^{k},\delta 2^{k}]^{2})}\varphi(P_{H})\Big)\leq c(-\log\delta)^{-1/2}d_{k}.$ (4.66)
Fix $\delta>0$ such that $c(-\log\delta)^{1/2}<0.01c_{2}$. We assume that $\beta>1$ and want to derive a contradiction. Choose $\theta_{\beta}$ such that $\beta\theta_{\beta}\in 2\pi\mathbb{Z}$ and $|e^{i\theta_{\beta}}-1|\geq 1$. By (4.63) and  (4.62), a.s. for all large $k$, for every $z\in B(0;c_{2}d_{k})$, $|\phi_{0}\circ\varphi^{-1}(z)|\leq\alpha 2^{k}$ and $|\phi_{0}\circ\varphi^{-1}(z)-a_{\beta}\mathrm{Re}\,z^{\beta}-b_{\beta}\mathrm{Im}\,z^{\beta}|<\alpha 2^{k-1}$ and therefore $|a_{\beta}\mathrm{Re}\,z^{\beta}+b_{\beta}\mathrm{Im}\,z^{\beta}|\leq\alpha 2^{k+1}$. This implies that, a.s. for all large $k$, $|a_{\beta}\mathrm{Re}\,z^{\beta}+b_{\beta}\mathrm{Im}\,z^{\beta}|\leq\alpha 2^{k-3}$ for $z\in B(0;c_{2}d_{k}/4)$ since $\beta\geq 2$. By another application of (4.62), a.s. for large $k$,
$|\phi_{0}\circ\varphi^{-1}(z)|\leq\alpha 2^{k-2},\ \ \text{for }z\in B(0;c_{2}d_{k}/4).$ (4.67)
Now let $w_{0}=a_{\beta}(c_{2}d_{k}/4)^{\beta}$. Then by (4.62), a.s. for large enough $k$,
$|\phi_{0}\circ\varphi^{-1}(\frac{c_{2}d_{k}}{4})-w_{0}|\leq 0.01\delta 2^{k},\ \ \ |\phi_{0}\circ\varphi^{-1}(\frac{c_{2}d_{k}e^{i\theta_{\beta}}}{4})-w_{0}|\leq 0.01\delta 2^{k}.$ (4.68)
By Lemma 2.6, a.s. for large enough $k$,
$\frac{c_{2}d_{k}}{4},\frac{c_{2}d_{k}e^{i\theta_{\beta}}}{4}\in\bigcup_{H\in\mathcal{H}(w_{0}+[-\delta 2^{k-1},\delta 2^{k-1}]^{2})}\varphi(P_{H}).$ (4.69)
By the line connectivity property as in Definition 1.4, a.s. for large $k$, the set in (4.69) is connected, and thus by (4.66),
$|\frac{c_{2}d_{k}}{4}-\frac{c_{2}d_{k}e^{i\theta_{\beta}}}{4}|\leq c(-\log\delta)^{-1/2}d_{k}<0.01c_{2}d_{k}.$
This contradicts with our assumption that $|e^{i\theta_{\beta}}-1|\geq 1$. ∎
Lemma 4.27.
The function $\widetilde{\phi}_{\infty}$ is a.s. a linear transform of $\mathbb{C}$. Furthermore, there exists a deterministic constant $c\in(0,1)$ such that
$B(0;c\mathrm{diam}(\widetilde{\phi}_{\infty}(\mathbb{D})))\subseteq\widetilde{\phi}_{\infty}(\mathbb{D}).$ (4.70)
Proof.
The claim that $\widetilde{\phi}_{\infty}$ is a linear transform of $\mathbb{C}$ follows from Proposition 4.20, Lemma 4.26 and the fact that $\phi_{\infty}(0)=0$. For the claim (4.70), let $c_{2},\ell$ be the constant from Lemma 4.8. We first show that
$\alpha 2^{k-\ell-1},\alpha 2^{k-\ell-1}]^{2}\subseteq\widetilde{\phi}_{\infty}(B(0;c_{2}d_{k}))\ \ \text{a.s.\ for large }k.$ (4.71)
Assume (4.71) does not hold. Using linearity of $\widetilde{\phi}_{\infty}$ along with Proposition 4.20, there exists some point $w_{0}\in[-0.9\alpha 2^{k-\ell},0.9\alpha 2^{k-1}]^{2}$ such that $w_{0}$ has distance at least $0.1\alpha 2^{k-\ell}$ from $\phi_{0}\circ\varphi^{-1}(B(0;c_{2}d_{k}))$. By Lemma 2.6 and Lemma 4.1, there exists cells $H_{1},H_{2},H_{3}$ with distance at least $0.05\alpha 2^{k-\ell}$ from $\phi_{0}\circ\varphi^{-1}(B(0;c_{2}d_{k}))$ and form a triangular face on $M([-\alpha 2^{k-\ell},\alpha 2^{k-\ell}]^{2})$. By (4.29), the $\varphi$-image of this face is contained within $B(0;c_{2}d_{k})$, and this contradicts with the statement that $H_{1},H_{2},H_{3}$ with distance at least $0.05\alpha 2^{k-\ell}$ from $\phi_{0}\circ\varphi^{-1}(B(0;c_{2}d_{k}))$. This verifies (4.71). By the second half of (4.29) along with Lemma 2.6, a.s. for large enough $k$, $\phi_{0}\circ\varphi^{-1}(B(0;c_{2}d_{k}))\subseteq[-\alpha 2^{k+1},\alpha 2^{k+1}]^{2}$, and by Proposition 4.20,
$\widetilde{\phi}_{\infty}(B(0;c_{2}d_{k}))\subseteq[-\alpha 2^{k+2},\alpha 2^{k+2}]^{2}\ \ \text{a.s.\ for large }k.$ (4.72)
Using linearity of $\widetilde{\phi}_{\infty}$, from (4.71) and (4.72) we see that a.s. for some $b>0$,
$B(0;b)\subseteq\widetilde{\phi}_{\infty}(\mathbb{D})\subseteq B(0;\alpha 2^{\ell+4}b).$ (4.73)
This verifies (4.70) for $c=(\alpha 2^{\ell+4})^{-1}$. ∎
Recall that for a square $S$, $\mathring{S}^{a}$ is the square concentric to $S$ with side length $a$. Below we write
$\nabla\widetilde{\phi}_{\infty}=\begin{pmatrix}\frac{\partial\mathrm{Re}\widetilde{\phi}_{\infty}}{\partial x}&\frac{\partial\mathrm{Im}\widetilde{\phi}_{\infty}}{\partial x}\\ \frac{\partial\mathrm{Re}\widetilde{\phi}_{\infty}}{\partial y}&\frac{\partial\mathrm{Im}\widetilde{\phi}_{\infty}}{\partial y}\end{pmatrix}$,
[…]
Lemma 4.28.
Almost surely, for each $\varepsilon,a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$,
$\widetilde{\phi}_{\infty}^{-1}(\mathring{S}^{(1-\varepsilon)|S|})\subseteq\varphi(M(S))\subseteq\widetilde{\phi}_{\infty}^{-1}(\mathring{S}^{(1+\varepsilon)|S|}).$ (4.74)
In particular,
$\lim_{k\to\infty}\frac{1}{|S_{k}|^{2}}|\det(\nabla\widetilde{\phi}_{\infty})|\,\mathrm{area}(\varphi(M(S_{k})))=1.$ (4.75)
Proof.
By Lemma 2.6 and Lemma 4.1, the Hausdorff distance between $\phi_{0}(\partial M(S))$ and $\partial S$ divided by $2^{k}$ a.s. tends to 0 as $n\to\infty$. Therefore by Proposition 4.20,
$\mathring{S}^{(1-\varepsilon)|S|}\subseteq\widetilde{\phi}_{\infty}\circ\varphi(M(S))\subseteq\mathring{S}^{(1+\varepsilon)|S|},$ (4.76)
which further implies (4.74). The relation (4.75) is immediate from (4.74) since $\widetilde{\phi}_{\infty}$ is linear. ∎
Our goal is to argue that $\widetilde{\phi}_{\infty}$ is the composition of a deterministic linear transform, a rotation and a scaling; once this has been proven, Theorem 1.8 follows via a short argument. Our next step towards this goal is to use the ergodic theorem for cell configurations (Lemma 2.4) to prove that for vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, up to a multiplicative constant, the inner product $\langle{\nabla\widetilde{\phi}_{\infty}\mathbf{u}},{\nabla\widetilde{\phi}_{\infty}\mathbf{v}}\rangle$ is deterministic.
Lemma 4.29.
Let $a,b\in\mathbb{R}$. Almost surely, the limit
$\lim_{k\to\infty}\frac{1}{|S_{k}|^{2}}\int_{M(S_{k})}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx$ (4.77)
exists and equals a deterministic finite number.
[…]
Proof.
For $z\in\mathbb{C}$, let
$\mathrm{SE}_{\infty}^{z,a,b}=\frac{\int_{\phi_{0}^{-1}(H_{z})}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx}{\mathrm{area}(H_{z})},$ (4.78)
and define $F(\mathcal{H},\mathcal{D})=\mathbb{E}[\mathrm{SE}_{\infty}^{0,a,b}|\mathcal{H},\mathcal{D}]$. Then from the construction of $\phi_{\infty}$, one can check that $F(\mathcal{H}-z,\mathcal{D}-z)=\mathbb{E}[\mathrm{SE}_{\infty}^{z,a,b}|\mathcal{H},\mathcal{D}]$ and $F(C\mathcal{H},C\mathcal{D})=F(\mathcal{H},\mathcal{D})$, thus by Lemma 2.4,
$\lim_{k\to\infty}\frac{1}{|S_{k}|^{2}}\int_{S_{k}}\mathbb{E}[\mathrm{SE}_{\infty}^{z,a,b}|\mathcal{H},\mathcal{D}]\,dz=\mathbb{E}[\mathrm{SE}_{\infty}^{0,a,b}].$ (4.79)
In order to conclude the proof we need to argue that the right side of the last indented equation is finite and that the left side does not change if the domain of integration is changed to be consistent with (4.77). Fix $\delta\in(0,1)$, and consider a square $S\subseteq S_{k}$ of side length at least $\delta|S_{k}|$. For each $z$ contained in some $H\in\mathcal{H}(S)$, if $z$ is inside the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ as in the definition of $\phi_{0}$, then for some $j$, $H_{j}\cap S\neq\emptyset$, and using Lemmas 2.6 and 4.1, it follows that a.s. for large enough $k$, $H_{1},H_{2},H_{3}\subseteq\mathring{S}^{1.5|S|}$ and thus $\phi_{0}^{-1}(z)\subseteq M(\mathring{S}^{2|S|})$. Therefore, a.s. for large $n$, for each square $S\subseteq S_{k}$ of side length at least $\delta|S|$,
$\begin{split}\frac{1}{|S|^{2}}\int_{S}&\mathrm{SE}_{\infty}^{z,a,b}dz\leq\frac{1}{|S|^{2}}\sum_{H\in\mathcal{H}(S)}\int_{\phi_{0}^{-1}(H)}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx\\ &\leq\frac{1}{|S|^{2}}\int_{M(\mathring{S}^{2|S|})}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx\\ &=\frac{1}{|S|^{2}}\int_{\varphi(M(\mathring{S}^{2|S|}))}\Big|\nabla\big(a\mathrm{Re}\,\widetilde{\phi}_{\infty}(z)+b\mathrm{Im}\,\widetilde{\phi}_{\infty}(z)\big)\Big|^{2}\,dz\\ &=\frac{1}{|S|^{2}}\Big|\big(a\nabla\mathrm{Re}\,\widetilde{\phi}_{\infty}+b\nabla\mathrm{Im}\,\widetilde{\phi}_{\infty}\big)\Big|^{2}\mathrm{area}(\varphi(M(\mathring{S}^{2|S|})))\\ &\leq 8(a^{2}+b^{2})|\nabla\widetilde{\phi}_{\infty}|^{2}|\det(\nabla\widetilde{\phi}_{\infty})|^{-1},\end{split}$ (4.80)
[…]
Finally,
$\begin{split}\frac{1}{|S_{k}|^{2}}\int_{S_{k}}\mathrm{SE}_{\infty}^{z,a,b}dz&\geq\frac{1}{|S_{k}|^{2}}\sum_{H\in\mathcal{H}(S_{k})\backslash\mathcal{H}(\partial S_{k})}\int_{\phi_{0}^{-1}(H)}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx\\ &\geq\frac{1}{|S_{k}|^{2}}\int_{M(S_{k})}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx,\end{split}$ (4.82)
and using the identical argument in the paragraph after (4.79), a.s. for large enough $k$,
$\begin{split}\frac{1}{|S_{k}|^{2}}\int_{\mathring{S}_{k}^{(1-\delta)|S_{k}|}}\mathrm{SE}_{\infty}^{z,a,b}dz&\leq\frac{1}{|S_{k}|^{2}}\sum_{H\in\mathcal{H}(\mathring{S}_{k}^{(1-\delta)|S_{k}|})}\int_{\phi_{0}^{-1}(H)}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx\\ &\leq\frac{1}{|S_{k}|^{2}}\int_{M(S_{k})}\Big|\nabla\mathbb{E}[a\mathrm{Re}\,\phi_{\infty}(x)+b\mathrm{Im}\,\phi_{\infty}(x)\big|\mathcal{H},\mathcal{D}]\Big|^{2}\,dx.\end{split}$ (4.83)
Therefore the conclusion follows by combining (4.79),  (4.81),  (4.82) and (4.83) and further sending $\delta\to 0$. ∎
Lemma 4.30.
There exists a deterministic symmetric bilinear function $g(\mathbf{u},\mathbf{v})$ for vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, such that almost surely,
$\frac{1}{|\det(\nabla\widetilde{\phi}_{\infty})|}\langle{\nabla\widetilde{\phi}_{\infty}\mathbf{u}},{\nabla\widetilde{\phi}_{\infty}\mathbf{v}}\rangle=g(\mathbf{u},\mathbf{v}).$ (4.84)
Furthermore, for some constant $c>0$, $g(\mathbf{u},\mathbf{u})\geq c|\mathbf{u}|^{2}$ for each $\mathbf{u}\in\mathbb{R}^{2}$.
Proof.
Using conformal invariance of Dirichlet energy, for $\mathbf{u}=(a,b)^{\mathrm{T}}$, we have
[…]
By (4.85), (4.75) and Lemma 4.29, $\frac{|\nabla\widetilde{\phi}_{\infty}\mathbf{u}|^{2}}{|\det(\nabla\widetilde{\phi}_{\infty})|}$ is a.s. deterministic, and using
[…]
we see that (4.84) holds for some deterministic symmetric bilinear function $g(\mathbf{u},\mathbf{v})$.
[…]
Lemma 4.31.
Let $\mathbf{P}\in\mathbb{R}^{2\times 2}$ be a $2\times 2$ random matrix with $\det(\mathbf{P})=1$, such that for every $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, the inner product $\langle{\mathbf{P}\mathbf{u}},{\mathbf{Pv}}\rangle=g(\mathbf{u},\mathbf{v})$ is deterministic. For $\theta\in[0,2\pi]$, we write $\mathbf{R}_{\theta}$ for the matrix for the linear transform $z\mapsto e^{i\theta}z$. Then there exists a deterministic matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and some random $\theta\in[0,2\pi]$, such that
$\mathbf{P}^{\mathrm{T}}=\mathbf{A}\mathbf{R}_{\theta}.$ (4.86)
Proof.
By our condition, there is a deterministic symmetric matrix $\mathbf{B}$ such that $g(\mathbf{u},\mathbf{v})=\mathbf{u}^{\mathrm{T}}\mathbf{B}\mathbf{v}$, and since $\det(\mathbf{P})=1$, $\mathbf{B}$ is positive definite. In particular, the eigenvectors of $\mathbf{B}$ are orthogonal, and we can pick some $\vartheta\in[0,2\pi]$, such that the eigenvectors of $\mathbf{B}$ are given by $(\cos\vartheta,\sin\vartheta)^{\mathrm{T}}$ and $(-\sin\vartheta,\cos\vartheta)^{\mathrm{T}}$. If we write
$\mathbf{Q}:=\mathbf{R}_{\vartheta}=\begin{pmatrix}\cos\vartheta&-\sin\vartheta\\ \sin\vartheta&\cos\vartheta\end{pmatrix}$,
then there is a diagonal matrix $\mathbf{D}$ such that $\mathbf{Q}^{\mathrm{T}}\mathbf{D}\mathbf{Q}=\mathbf{B}$. We set $\mathbf{A}=\mathbf{Q}^{\mathrm{T}}\mathbf{D}^{1/2}$, and pick the positive square root of $\mathbf{D}$ such that $\det(\mathbf{A})>0$. Then $\langle{\mathbf{P}\mathbf{u}},{\mathbf{P}\mathbf{v}}\rangle=\langle{\mathbf{A}^{\mathrm{T}}\mathbf{u}},{\mathbf{A}^{\mathrm{T}}\mathbf{v}}\rangle$ for every $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$ and therefore $\langle{\mathbf{P}(\mathbf{A}^{\mathrm{T}})^{-1}\mathbf{u}},{\mathbf{P}(\mathbf{A}^{\mathrm{T}})^{-1}\mathbf{v}}\rangle=\langle{\mathbf{u}},{\mathbf{v}}\rangle$ for every $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$. This implies that $\mathbf{P}(\mathbf{A}^{\mathrm{T}})^{-1}$ is orthogonal. Since $\det(\mathbf{P}(\mathbf{A}^{\mathrm{T}})^{-1})>0$, we have $\mathbf{P}(\mathbf{A}^{\mathrm{T}})^{-1}=\mathbf{R}_{-\theta}$ for some possibly random $\theta$; this implies (4.86) and $\det(\mathbf{A})=1$. ∎
Lemma 4.32.
There exists a deterministic matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and some random $\theta\in[0,2\pi]$, such that if we write $\mathbf{R}_{\theta}$ for the matrix for the linear transform $z\mapsto e^{i\theta}z$, then a.s.
$(\nabla\widetilde{\phi}_{\infty})^{\mathrm{T}}=|\det(\nabla\widetilde{\phi}_{\infty})|^{1/2}\mathbf{A}\mathbf{R}_{\theta}.$ (4.87)
In other words, up to a scaling, the linear transform $\widetilde{\phi}_{\infty}$ is given by first rotating for some random angle $\theta$, then applying the linear transform $\mathbf{A}$.
Proof.
We first prove $\det(\nabla\widetilde{\phi}_{\infty})>0$. For large $n$, we choose an orientation of $M(S_{n})$ such that for the clockwise loop $\gamma_{n}$ on the boundary of $M(S_{n})$, $\phi_{0}\circ\gamma_{n}$ is clockwise. Then $\phi_{0}\circ\gamma_{n}$ has Hausdorff distance at most $0.01|S_{n}|$ from $\partial S_{n}$ by Lemma 4.1, and $\varphi\circ\gamma_{n}$ is also a clockwise loop. Therefore by Proposition 4.20, a.s. for large enough $n$, the loop $\widetilde{\phi}_{\infty}\circ\varphi\circ\gamma_{n}$ is clockwise as both $\widetilde{\phi}_{\infty}\circ\varphi\circ\gamma_{n}$ and $\phi_{0}\circ\gamma_{n}$ have distance at least $0.1|S_{n}|$ from the center of $S_{n}$ and the two curves can be parameterized such that $|\widetilde{\phi}_{\infty}\circ\varphi\circ\gamma_{n}(t)-\phi_{0}\circ\gamma_{n}(t)|\leq 0.02|S_{n}|$, which implies that $\det(\nabla\widetilde{\phi}_{\infty})>0$. Now Lemma 4.32 follows by applying Lemma 4.31 to $\mathbf{P}=\det(\nabla\widetilde{\phi}_{\infty})^{-1/2}\nabla\widetilde{\phi}_{\infty}$. ∎
Lemma 4.33.
[…]
Proof.
By assumption, $\mathcal{H}$ and $e^{i\theta_{0}}\mathcal{H}$ have the same law, and the corresponding matrices $\mathbf{A}$ in Lemma 4.32 are the same. Therefore by Proposition 4.20 and Lemma 4.32, there exists $\vartheta_{1},\vartheta_{2}$ such that
$\mathbf{R}_{\theta_{0}}\mathbf{A}\mathbf{R}_{\vartheta_{1}}=\mathbf{A}\mathbf{R}_{\vartheta_{2}},$ (4.88)
which is equivalent to $\mathbf{A}^{-1}\mathbf{R}_{-\theta_{0}}=\mathbf{R}_{\vartheta_{1}-\vartheta_{2}}\mathbf{A}^{-1}$. In particular, for each $\mathbf{u}\in\mathbb{R}^{2}$, $|\mathbf{A}^{-1}\mathbf{R}_{-\theta_{0}}\mathbf{u}|=|\mathbf{A}^{-1}\mathbf{u}|$. Recall that in the proof of Lemma 4.32, we have set $\mathbf{A}=\mathbf{Q}^{\mathrm{T}}\mathbf{D}^{1/2}=\mathbf{R}_{-\vartheta}\mathbf{D}^{1/2}$, and $\mathbf{D}$ is a diagonal matrix with nonnegative entries. Therefore $|\mathbf{D}^{-1/2}\mathbf{R}_{\vartheta-\theta_{0}}\mathbf{u}|=|\mathbf{D}^{-1/2}\mathbf{R}_{\vartheta}\mathbf{u}|$ and further $|\mathbf{D}^{-1/2}\mathbf{R}_{-\theta_{0}}\mathbf{u}|=|\mathbf{D}^{-1/2}\mathbf{u}|$ for every $\mathbf{u}\in\mathbb{R}^{2}$. By setting $\mathbf{u}=(1,0)^{\mathrm{T}}$ and $\mathbf{u}=(0,1)^{\mathrm{T}}$ and using $\theta_{0}\neq\pi$, one can check that $\mathbf{D}$ must be a scalar times the identity matrix. Since $\det(\mathbf{A})=1$, $\mathbf{D}$ must be the identity matrix and $\mathbf{A}$ is equal to the rotation matrix $\mathbf{R}_{-\vartheta}$. ∎
Proof of Theorem 1.8.
By Lemma 4.9, $M(\mathcal{H})$ is a.s. parabolic. Recall that $\varphi$ is a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ chosen in a way measurable with respect to $\mathcal{H}$ such that $\varphi(0)=0$. By Proposition 4.20, Lemma 4.27 and Lemma 4.32, there exists a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and some possibly random $\theta$ and $\widetilde{a}>0$, such that
$\lim_{k\to\infty}2^{-k}\sup_{x\in M([-2^{k},2^{k}]^{2})}|\phi_{0}(x)-\widetilde{a}\mathbf{A}^{-1}e^{i\theta}\varphi(x)|=0.$ (4.89)
By Lemma 4.33, if $\mathcal{H}$ satisfies the condition in Definition 1.6, then $\mathbf{A}$ can be taken to be identity matrix. Almost surely for large enough $k$, for each $H\in\mathcal{H}([-2^{k},2^{k}]^{2})$, by Lemma 2.6 and Lemma 4.1, the vertex $v_{H}$ for $H$ on $M(\mathcal{H})$ is on $M([-2^{k+1},2^{k+1}]^{2})$, and thus by (4.89),
$\lim_{k\to\infty}2^{-k}\max_{H\in\mathcal{H}([-2^{k},2^{k}]^{2})}|\phi_{0}(v_{H})-\widetilde{a}\mathbf{A}^{-1}e^{i\theta}\varphi(v_{H})|=0.$ (4.90)
By Lemma 2.6,
$\lim_{k\to\infty}\max_{H\in\mathcal{H}([-2^{k},2^{k}]^{2})}2^{-k}|c(H)-\phi_{0}(v_{H})|=0.$ (4.91)
Thus if we set $\varphi_{0}=\widetilde{a}e^{i\theta}\varphi$, then by (4.90) and (4.91),
$\lim_{k\to\infty}\max_{H\in\mathcal{H}([-2^{k},2^{k}]^{2})}2^{-k}|c(H)-\mathbf{A}^{-1}\varphi_{0}(v_{H})|=0,$ (4.92)
which further implies (1.5). Now a.s. for large enough $k$, for each edge $e\in\mathcal{E}\mathcal{H}(B(0;2^{k}))$, by Lemma 2.6 and Lemma 4.1, the edge $\mathfrak{e}_{e}\subseteq M([-2^{k+1},2^{k+1}]^{2})$. By Lemma 2.6,
$\lim_{k\to\infty}2^{-k}\max_{e\in\mathcal{E}\mathcal{H}(B(0;2^{k}))}\mathrm{diam}(\phi_{0}(\mathfrak{e}_{e}))=0.$ (4.93)
The relation (1.6) then follows from (4.89) and (4.93). ∎
5 Extensions of Theorems 1.7 and 1.8
In this section, we explain several extensions of Theorems 1.7 and 1.8. In Section 5.1, we prove Theorem 5.1, which is a version of Theorems 1.7 and 1.8 with a weaker notion of the line connectivity property in Definition 1.4. In Section 5.2, we prove an extension of Theorem 1.7 for a version of circle packing for general planar maps. In Section 5.3, we prove an extension of Theorem 1.8 for $p$-angulations with $p\in\{4,5,\dots\}$.
5.1 An extension with weaker notion of line connectivity
The following version of Theorem 1.7 and Theorem 1.8, which uses a weaker version of the line connectivity in Definition 1.4, will be used in our future work.
Theorem 5.1.
Suppose $\mathcal{H}$ is a cell configuration that satisfies all the conditions in Theorem 1.7 or Theorem 1.8 except for the line connectivity condition in Definition 1.4. Suppose there exists a cell configuration $\mathcal{H}^{\rm large}$ satisfying the condition in Definition 1.3 and the moment bound (1.3) along with the following properties:
  1. 1.
    The cell configuration $\mathcal{H}^{\rm large}$ is isomorphic to $\mathcal{H}$, in the sense that $\mathcal{H}$ and $\mathcal{H}^{\rm large}$ have isomorphic associated maps; below we let $\phi:\mathcal{H}\to\mathcal{H}^{\rm large}$ denote the isomorphism.
  2. 2.
    The convex hull of $H$ is contained in the convex hull of $\phi(H)$ for any $H\in\mathcal{H}$.
  3. 3.
    Almost surely, for each given $\varepsilon>0$, there exists $R>0$ such that for $r>R$ and each horizontal or vertical line segment $\ell\subseteq B(0;r)$ with length at least $\varepsilon r$, the following holds. For any two cells $H,H^{\prime}\in\mathcal{H}$ intersecting $\ell$, we can find a path $H\sim H_{1}\sim\cdots\sim H_{k}\sim H^{\prime}$, such that $\phi(H_{j})$ intersects $\ell$ for each $j$.
Then the conclusions in Theorem 1.7 or Theorem 1.8 continue to hold.
Theorem 5.1 can be viewed as a generalization of Theorem 1.7 and Theorem 1.8 since one can set $\mathcal{H}^{\rm large}=\mathcal{H}$ in the latter setting.
Proof.
Most of the proof is identical to that of Theorem 1.7 and Theorem 1.8, and we will only explain the differences. Note that the arguments in Section 2.2 do not require the line connectivity property in Definition 1.4 or the almost planarity condition in Defininition 1.5, so the conclusions there continue to hold. In particular, by Lemma 2.6 applied to $\mathcal{H}^{\rm large}$ and condition (ii) in the theorem statement, the almost planarity condition in Definition 1.5 holds for $\mathcal{H}^{\rm large}$ as well.
In the setting of circle packing, the arguments in Section 3 do not directly use the line connectivity property in Definition 1.4. The property is used implicitly as we carry the proof of [GMS22, Theorem 1.16]; in the proof in [GMS22], this property is used only in the proof of [GMS22, Lemma 2.19], in the paragraph right after [GMS22, Eq. (2.46)]. To substitute for that, Conditions (ii) and (iii) guarantee that a.s. for large enough $k$ and horizontal or vertical line segments $\ell\subseteq S_{k}$ of length $S_{k}$, for any function $g:\mathcal{H}^{\rm large}\to\mathbb{R}$ and $H_{1},H_{2}\in\mathcal{H}(\ell)$,
$|g(\phi(H_{1}))-g(\phi(H_{2}))|\leq\sum_{H,H^{\prime}\in\mathcal{H}^{\rm large}(\ell):H\sim H^{\prime}}|g(H)-g(H^{\prime})|.$
The corresponding quantity on the right side can be bounded exactly as [GMS22] since the arguments there only rely on the inputs from Section 2.2, which do not require the line connectivity property.
[…]
$L^{\rm large}(\rho)=\sum_{H^{\rm large}\in\mathcal{H}^{\rm large}(\partial\mathring{S}^{\rho})}\mathrm{diam}(\varphi_{k}(P_{H^{\rm large}})).$
Then  (4.11) holds with $L^{\rm large}(\rho)$ instead of $L(\rho)$ using identical arguments applied to $\mathcal{H}^{\rm large}$, and (4.5) follows from identical arguments using conditions (ii) and (iii). For the proof of Lemma 4.15, using conditions (ii) and (iii) along with Lemma 2.6, one can find cells $H_{1}\sim\cdots H_{n}\sim H_{1}$ with identical properties and the conclusion still holds. The arguments for Lemma 4.26 are identical, except that at the final step we use that by Lemma 4.1 the set on the right hand side of (4.69) is contained within $\varphi(M(w+[-\delta 2^{k},\delta 2^{k}]^{2}))$, which is a connected subset of the set in (4.66). The rest of the proofs in Sections 4.1-4.4 do not require the line connectivity property and can be copied to the setting of Theorem 5.1. ∎
5.2 An Extension of Theorem 1.7 to general planar maps
Let $\mathcal{M}$ be an one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$. Recall from Section 1.4 that one can add a vertex $v_{f}$ to every face $f\in\mathcal{F}\mathcal{M}$, and connect the vertices $v_{f}$ to the vertices incident to $f$, to get a triangulation $\mathcal{T}_{\mathcal{M}}$. The triangulation $\mathcal{T}_{\mathcal{M}}$ will be simple since we required $\mathcal{M}$ to be 2-connected. We define the circle packing of $\mathcal{M}$ to be the circle packing for $\mathcal{T}_{\mathcal{M}}$ minus the disks for the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. In this section we prove a variant of Theorem 1.7 in this setting.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ as above, and let $\mathcal{T}_{\mathcal{M}}$ be the triangulation generated from $\mathcal{M}$. We write $v_{H}\sim f$ if the vertex $v_{H}$ for $H$ in the associated map $\mathcal{M}$ is incident to the face $f\in\mathcal{F}\mathcal{M}$. Let
$\overline{H}_{0}=\bigcup_{f\in\mathcal{F}\mathcal{M}:v_{H_{0}}\sim f}\bigcup_{H\in\mathcal{H}:v_{H}\sim f}H$ (5.1)
be the union of cells whose vertex is on the same face with $v_{H_{0}}$ and let $d_{0}$ be the maximal degree of the faces incident to $v_{H_{0}}$.
Theorem 5.2.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ being an one-ended 2-connected simple planar map. Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5. Further assume
$\mathbb{E}\Big[\frac{\mathrm{diam}(\overline{H}_{0})^{2}}{\mathrm{area}(H_{0})}\big(\deg(H_{0})+d_{0}\big)^{4}\deg(H_{0})\Big]<\infty,$ (5.2)
Then the conclusions in Theorem 1.7 holds for the circle packing of $\mathcal{M}$ defined as above.
[…]
Lemma 5.3.
The cell configuration $\widetilde{\mathcal{H}}$ is translation invariant modulo scaling as in Definition 1.2.
Proof.
Let $U_{j},z_{j}$ and $C_{j}$ be as in Definition 1.2 for the cell configuration $\mathcal{H}$, such that $C_{j}(\mathcal{H}-z_{j})\to\mathcal{H}$ in law with respect to the metric (1.1). Then by our way of construction of $\widetilde{\mathcal{H}}$ using $\mathcal{H}$ where we work with each single cell $H$, we have $C_{j}(\widetilde{\mathcal{H}}-z_{j})\to\widetilde{\mathcal{H}}$ in law with respect to the metric (1.1). This verifies the claim. ∎
Lemma 5.4.
The cell configuration $\widetilde{\mathcal{H}}$ is ergodic modulo scaling as in Definition 1.3.
Proof.
[…]
$F_{k}(\widetilde{\mathcal{H}},\mathcal{D}):=\mathbb{E}\Big[F(\widetilde{\mathcal{H}})\big|\widetilde{\mathcal{H}}(\widehat{S}_{k})\Big].$
[…]
Therefore for any $\varepsilon>0$, as $k\to\infty$,
$\begin{split}\mathbb{P}\big[|F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})-F(\widetilde{\mathcal{H}})|>\varepsilon\big]&=\mathbb{P}\big[|F_{k}(\widetilde{C}(\widetilde{\mathcal{H}}-z_{m}),\widetilde{C}(\mathcal{D}-z_{m}))-F(\widetilde{C}(\widetilde{\mathcal{H}}-z_{m}))|>\varepsilon\big]\\ &=\mathbb{P}\big[|F_{k}(\widetilde{\mathcal{H}},\mathcal{D})-F(\widetilde{\mathcal{H}})|>\varepsilon\big]\to 0,\end{split}$ (5.4)
and the convergence is uniform in $m>0$. In particular, since $0\leq F\leq 1$, by (5.4),
$\begin{split}&\mathbb{P}\bigg[\bigg|\mathbb{E}\Big[F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})-F(\widetilde{\mathcal{H}})\big|\widetilde{\mathcal{H}}(\widehat{S}_{k})\Big]\bigg|\geq\varepsilon\bigg]\to 0,\\ &\mathbb{P}\bigg[\bigg|\mathbb{E}\Big[F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})-F(\widetilde{\mathcal{H}})\big|\mathcal{H}(\widehat{S}_{k})\Big]\bigg|\geq\varepsilon\bigg]\to 0\text{ uniformly in $m$ as $k\to\infty$.}\end{split}$ (5.5)
Now for fixed $k>0$, $F_{k}(\widetilde{\mathcal{H}},\mathcal{D})$ is measurable with respect to $(\widehat{S}_{k},\widetilde{\mathcal{H}}(\widehat{S}_{k}))$, and for any fixed $z$, $F_{k}(\widetilde{\mathcal{H}}-z,\mathcal{D}-z)$ is measurable with respect to $(\widehat{S}_{k}^{z},\widetilde{\mathcal{H}}(\widehat{S}_{k}^{z}))$. For $H\in\mathcal{H}$, let $\Theta_{H}$ denote the random variables we used to partition $H$ in the construction of $\widetilde{\mathcal{H}}$, and for a square $S$, let $\overline{\mathcal{H}}(S)$ be the collection of cells $H$ in $\mathcal{H}$ such that the vertices for $H$ and some cell in $\mathcal{H}(S)$ are on a common face of $\mathcal{M}$. For $z\in\mathbb{C}$, let $A_{z,k}$ be the event such that $\overline{\mathcal{H}}(\widehat{S}^{z}_{k})$ has positive distance from $\overline{\mathcal{H}}(\widehat{S}_{k})$. For deterministic $z$, conditioned on $A_{z,k}$ and $\mathcal{H}(\widehat{S}_{k})$, $F_{k}(\widetilde{\mathcal{H}}-z,\mathcal{D}-z)$ is independent from $\{\Theta_{H}:H\in\mathcal{H}(\widehat{S}_{k})\}$. This implies
$\mathbb{E}\Big[\mathds{1}_{A_{z_{m},k}}F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})\big|\widetilde{\mathcal{H}}(\widehat{S}_{k}),z_{m}\Big]=\mathbb{E}\Big[\mathds{1}_{A_{z_{m},k}}F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})\big|\mathcal{H}(\widehat{S}_{k}),z_{m}\Big].$
Since $z_{m}$ is sampled from the Lebesgue measure on $\widehat{S}_{m}$, we see that for fixed $k>0$, $\mathbb{P}[A_{z_{m},k}]\to 1$ for fixed $k$ as $m\to\infty$, and therefore for fixed $\varepsilon>0$,
$\mathbb{P}\bigg[\Big|\mathbb{E}\Big[F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})\big|\widetilde{\mathcal{H}}(\widehat{S}_{k})\Big]-\mathbb{E}\Big[F_{k}(\widetilde{\mathcal{H}}-z_{m},\mathcal{D}-z_{m})\big|\mathcal{H}(\widehat{S}_{k})\Big]\Big|>\varepsilon\bigg]\to 0\text{ as }m\to\infty.$ (5.6)
By (5.5) and (5.6), sending $m\to\infty$ sufficiently faster than $k\to\infty$, for any $\varepsilon>0$, we have
$\mathbb{P}\Big[\Big|\mathbb{E}[F(\widetilde{\mathcal{H}})|\widetilde{\mathcal{H}}(\widehat{S}_{k})]-\mathbb{E}[F(\widetilde{\mathcal{H}})|\mathcal{H}(\widehat{S}_{k})]\Big|>\varepsilon\Big]\to 0\ \text{as $k\to\infty$.}$
[…]
Lemma $5.$
There are no macroscopic faces of $\mathcal{H}$, in the sense that almost surely for any $\varepsilon>0$, for large enough $k$, for each cell $H_{1},...,H_{j}\in\mathcal{H}$ with $(H_{1}\cup...\cup H_{j})\cap S_{k}\neq\emptyset$ and $v_{H_{1}},...,v_{H_{k}}$ lie on the same face of $\mathcal{M}$, $\mathrm{diam}(H_{1}\cup...\cup H_{j})<\varepsilon|S_{k}|$.
Proof.
The proof is identical to that of [GMS22, Lemma 2.9] where one consider the set $\overline{H}_{0}$ from (5.1) and use the moment bound from (5.2). ∎
Proof of Theorem 5.2.
Let $\widetilde{\mathcal{H}}$ be the cell configuration constructed as above. The ergodicity modulo scaling property for $\widetilde{\mathcal{H}}$ holds by Lemma 5.4. Since each $\widetilde{H}_{H,k}$ contains the boundary of $H$, we see that the line connectivity property from Definition 1.4 holds. The almost planarity condition in Definition 1.5 also holds since one can add vertices and edges directly to the planar map that is close to $\mathcal{H}$ in the sense of Definition 1.5 and there are no macroscopic faces of $\mathcal{H}$ by Lemma $5.$ Furthermore, the moment bound (1.3) for $\widetilde{\mathcal{H}}$ follows from the moment bound (5.2). Therefore the conditions in Theorem 1.7 hold for $\widetilde{\mathcal{H}}$ and the conclusion follows from Theorem 1.7. ∎
5.3 Extension of Theorem 1.8 to $p$-angulations
In Theorem 1.8, we worked with the setting where the associated map $\mathcal{M}$ of the cell configuration $\mathcal{H}$ is a triangulation. In this section, we briefly explain how to extend Theorem 1.8 to the setting where $\mathcal{M}$ is a $p_{0}$-angulation for $p_{0}\geq 3$.
Let $\mathcal{M}$ be a one-ended planar map with no boundary. Assume that $\mathcal{M}$ has no loops, and every face of $\mathcal{M}$ is adjacent to $p_{0}$ vertices. We let $M(\mathcal{M})$ be the Riemann surface constructed in the same way as in the triangulation settings in Section 2.1, where we view each face of $\mathcal{M}$ as a regular $p_{0}$-gon with unit side length.
Theorem 5.6.
Let $p_{0}\geq 3$. Let $\mathcal{H}$ be a random cell configuration such that its associated map is a one-ended $p_{0}$-angulation as above. Let $M(\mathcal{H})$ be the Riemann surface associated with the associated map of $\mathcal{H}$ defined as above. Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound (1.3). Then the conclusions in Theorem 1.8 still holds.
Proof.
The proof is almost identical to the proof of Theorem 1.8. The results in Section 2.2 still holds since we did not assume there that the associated map of $\mathcal{H}$ is a triangulation. To derive the regularity estimates in Section 4.1, one define the flowers and semi-flowers in the same way as in the triangulation case. It is straightforward to check that Lemma 4.2 still holds in this setting, and the rest of the proof in Section 4.1 still carries. To define the function $\phi_{0}$ in Section 4.2, we sample the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$ in the same way. For each regular $p_{0}$-gon on $M(\mathcal{H})$ formed by the cells $H_{1},...,H_{p_{0}}$, we further uniformly at random choose a way to draw straight lines to partition this face into $p_{0}-2$ triangles, and let $\phi_{0}$ be the piecewise linear function sending each such triangle on $M(\mathcal{H})$ to the corresponding triangle formed by the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$. Then one can check that for each face on $M(\mathcal{H})$ formed by $H_{1},...,H_{p_{0}}$, the Dirichlet energy of $\phi_{0}$ on this face is controlled by $C(p_{0})(\mathrm{diam}(H_{1})^{2}+....+\mathrm{diam}(H_{p_{0}})^{2})$ where $C(p_{0})$ is a constant depending only on $p_{0}$. This will allow us to bound the Dirichlet energy per area for $\phi_{0}$, and the rest of the proof is identical. ∎

### 06 May 2026 v2

@@
-Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment
+Circle packing and Riemann uniformization of random planar maps in an ergodic scale-free environment
 Abstract
-We prove that embedded infinite plane triangulations in ergodic scale-free environments are close to their circle packing and Riemann uniformization embedding on a large scale, as long as suitable moment and connectivity conditions are satisfied.
+We prove that embedded infinite planar maps in ergodic scale-free environments are close to their circle packing and Riemann uniformization embedding on a large scale, as long as suitable moment and connectivity conditions are satisfied.
 Ergodic scale-free environments were earlier considered by Gwynne, Miller and Sheffield (2018) in the context of the invariance principles for random walk, and they arise naturally in the study of random planar maps and Liouville quantum gravity.
@@
 This defines a Riemann surface $M$, and one uses the Riemann uniformization theorem to draw the vertices and edges of $\mathcal{M}$ via a conformal map $\varphi:M\to\mathbb{C}$.
-In this paper, we work with random infinite triangulations under the two aforementioned conformal embeddings.
+In this paper, we work with random planar maps with whole-plane topology, i.e., one-ended planar maps without boundary, under the two aforementioned conformal embeddings.
 Circle packing for random triangulations have been studied in e.g.
@@
 [GR13, CUR15].
-Our main result is, for a family of infinite plane triangulations in ergodic scale-free environments, they are close to their circle packing and Riemann uniformization embedding on a large scale.
+Our main result is, for a family of random planar maps with whole-plane topology in ergodic scale-free environments, they are close to their circle packing and Riemann uniformization embedding on a large scale.
 Our result holds under a moment condition on the degrees of the triangulation and the regularity of the environment, along with a condition guaranteeing that the embedded triangulation is sufficiently well-connected.
@@
 In Section 1.2 we introduce ergodic scale-free cell configurations.
-In Section 1.3 we state our main results, Theorems 1.7 and 1.8.
+In Section 1.3 we state our main results, Theorems 1.6 and 1.7.
 In Section 1.4 we discuss some applications and future directions related to this work.
-Finally in Section 1.5, we give a detailed outline of our proof of Theorems 1.7 and 1.8 in the rest of the paper.
+Finally in Section 1.5, we give a detailed outline of our proof of Theorems 1.6 and 1.7 in the rest of the paper.
 1.1 Discrete conformal embeddings
@@
 A planar map $\mathcal{M}$ is an embedding of a planar graph $\mathcal{G}$ modulo orientation-preserving homeomorphisms of $\widehat{\mathbb{C}}$.
-A face is a connected component of the complement of the planar map.
+The face set $\mathcal{F}\mathcal{M}$ is the set of the connected components of the complement of the planar map, and we say that $\mathcal{M}$ has whole-plane topology if all the faces are bounded in $\mathbb{C}$, and the union of faces in $\mathcal{F}\mathcal{M}$ along with their closures is equal to $\mathbb{C}$.
 A triangular face is a face with three edges on its boundary.
@@
 This means in particular that embeddings of the same planar map in different domains induce (or is believed to induce) a map that approximates the conformal map between the two domains.
-For circle packing, this was conjectured by Thurston and solved in [RS87] in the uniformly bounded degree case [^1] The paper only considers the case of the hexagonal lattice, but it is commented in [HS96] that the proof in [RS87] also works for uniformly bounded degrees.
-and by [HS96] in full generality.
+For circle packing, this was conjectured by Thurston and solved in [RS87] in the case of the hexagonal lattice, in [STE90, STE96] for circle packings with uniformly bounded radii ratio for neighboring disks via probabilistic methods, and by [HS96] in full generality.
 For Riemann uniformization embedding, this follows from the definition.
@@
 A second natural way to embed random planar maps is through the Riemann uniformization theorem.
-Roughly speaking, for a triangulation $\mathcal{T}$, one associate each face with an equilateral triangle and glue them isometrically along their edges to get a Riemann surface $M(\mathcal{T})$; see Section 2.1 for details.
-If $\mathcal{T}$ is an infinite plane triangulation, there are again two possible cases, where the Riemann surface $M(\mathcal{T})$ is parabolic (resp.
+Roughly speaking, for a planar map $\mathcal{M}$, one associate each face of degree $p$ with a regular $p$-gon of unit side length and glue them isometrically along their edges to get a Riemann surface $M(\mathcal{M})$; see Section 2.1 for details.
+If $\mathcal{M}$ has whole-plane topology, there are again two possible cases, where the Riemann surface $M(\mathcal{M})$ is parabolic (resp.
 hyperbolic) if it is conformally equivalent to the whole plane (resp.
 the unit disk).
-As commented in [GR13], similar to the circle packing case, this is equivalent to the recurrence/transience of the Brownian motion on the surface $M(\mathcal{T})$.
+As commented in [GR13, AHN+18], similar to the circle packing case, this is equivalent to the recurrence/transience of the Brownian motion on the surface $M(\mathcal{M})$.
 See e.g.
-[GR13, CUR15] for some works in this direction.
+[GR13, CUR15, AHN+18] for some works in this direction.
 There exist a number of discrete conformal embeddings beyond the two discussed above, including the Tutte embedding, the Smith embedding [BSS+40], the Cardy-Smirnov embedding [HS23], and the t-embedding/Coulomb gauge/s-embedding [CHE18, CLR23, KLR+22].
@@
 In this subsection, we introduce cell configurations [GMS22], as well as a number of conditions we need in our main results.
-Definition $1.$
+Definition 1.1.
 A cell configuration on $\mathbb{C}$ consists of the following objects.
@@
     A planar map $\mathcal{M}$, which we call the associated map of $\mathcal{H}$, with vertices $\{v_{H}:H\in\mathcal{H}\}$.
-We require that if $v_{H}$ and $v_{H^{\prime}}$ are connected by an edge of $\mathcal{M}$ then $H\neq H^{\prime}$ and $H\cap H^{\prime}\neq\emptyset$.
+We require that if $v_{H}$ and $v_{H^{\prime}}$ are connected by an edge of $\mathcal{M}$ then $H\cap H^{\prime}\neq\emptyset$.
   3. 3.
@@
 We emphasize that, in the definition of cell configurations, it is allowed that two cells $H$ and $H^{\prime}$ have nonempty intersections while $H$ and $H^{\prime}$ are not connected by any edge in $\mathcal{E}\mathcal{H}$.
+We also allow that the cells in $\mathcal{H}$ are not simply connected and may cross each other.
 See Figure 1 for an illustration.
@@
 In this subsection, we state our main results on the closeness between cell configurations and their discrete conformal embeddings.
-We work on the case where the associated map $\mathcal{M}$ of $\mathcal{H}$ is an infinite plane triangulation.
-We emphasize that the cells in $\mathcal{H}$ do not need to be simply connected, and different cells may cross each other; see Figure 1.
-However, we do need the following almost planarity assumption, which says that $\mathcal{H}$ is close to some planar embedding of $\mathcal{M}$.
-In the rest of the paper, without specific mentioning, we will always assume that all the conductances are equal to 1, i.e., $\mathfrak{c}(e)=1$ for every edge $e\in\mathcal{E}\mathcal{H}$.
+We start with the case where the associated map $\mathcal{M}$ of $\mathcal{H}$ is an infinite plane triangulation.
 Figure 1: An illustration of five cells in a cell configuration $\mathcal{H}$.
@@
 We allow that the green/blue and red/purple cells to touch each other while not being connected by an edge in $\mathcal{E}\mathcal{H}$.
-Definition 1.5.
-Let $\mathcal{H}$ be a cell configuration as in Definition $1.$ We say $\mathcal{H}$ is almost planar, if there exists a collection of points $\{z_{H}:H\in\mathcal{H}\}$ and simple curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in $\mathbb{C}$ forming a planar embedding of the associated map of $\mathcal{H}$, and almost surely as $r\to\infty$,
-$\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}\max_{e}\,\operatorname{dist}(\gamma_{e},H)\to 0,$ (1.2)
-where the second maximum is over edges with $H$ being an endpoint and the distance is the Hausdorff distance.
 Definition 1.5.
@@
 Let $\mathcal{H}$ be a random cell configuration such that its associated map is a simple infinite plane triangulation.
-Suppose $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5.
+Suppose $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4.
 Further assume $\mathcal{H}$ satisfies the moment bound
@@
 circle packing parabolic, and there exists a circle packing $\mathcal{P}$ for $\mathcal{H}$ and a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ such that the following is true.
-For $H\in\mathcal{H}$, let $c(H)=\frac{1}{\mathrm{area}(H)}\int_{H}z\,dz$ be the Euclidean center of $H$, and $o(H)$ be the center of the disk for $H$ in $\mathcal{P}$.
-Then almost surely,
+Almost surely, for any $\varepsilon>0$, there exists $R_{0}>0$ such that for every $r>R_{0}$, the maximal diameter of disks in $\mathcal{P}$ intersecting $B(0;r)$ is less than $\varepsilon r$.
+Furthermore, for $H\in\mathcal{H}$, letting $c(H)=\frac{1}{\mathrm{area}(H)}\int_{H}z\,dz$ be the Euclidean center of $H$, and letting $o(H)$ be the center of the disk for $H$ in $\mathcal{P}$, almost surely,
 $\lim_{r\to\infty}\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}|\mathbf{A}c(H)-o(H)|=0.$ (1.3)
-Furthermore, if $\mathcal{H}$ satisfies the rotation invariance in Definition 1.6, then $\mathbf{A}$ can be taken to be the identity matrix.
+Finally, if $\mathcal{H}$ satisfies the rotation invariance in Definition 1.5, then $\mathbf{A}$ can be taken to be the identity matrix.
 Theorem 1.7.
 Let $\mathcal{H}$ be a random cell configuration such that its associated map is an infinite plane triangulation.
-Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5.
+Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3 and 1.4.
 Further assume $\mathcal{H}$ satisfies the moment bound (1.2).
@@
 Finally, if $\mathcal{H}$ further satisfies the rotation invariance in Definition 1.5, then $\mathbf{A}$ can be taken to be the identity matrix.
-In Section 5.2, we will prove that Theorem 1.7 also holds for general planar maps (non-triangulations) for a particular choice of the circle-packing and under a moment condition, where we recall that circle packings are typically not unique for general planar maps.
-In Section 5.3, we will show that Theorem 1.8 also holds if $\mathcal{M}$ is a $p$-angulation for any $p\geq 3$, where we recall that a $p$-angulation is a planar map where all faces have degree pp.
-We comment that the convergence results of random planar maps under discrete conformal embeddings in [GMS20, GMS21, HS23, BGS25] are all stated for finite planar maps, whereas Theorems 1.7 and 1.8 are stated for infinite plane triangulations.
-In our subsequent work, we will apply Theorems 1.7 and 1.8 to work on the convergence of random planar maps under circle packing and Riemann uniformization embedding in both finite volume and infinite volume setting.
-More precisely, we will use a version of Theorems 1.7 and 1.8 (Theorem 5.1) with a weaker version of the line connectivity in Definition 1.4, which will be proved in Section 5.1.
+For infinite plane triangulations, the circle packing is unique.
+For general planar maps, the circle packing is not necessarily unique; however, one can get uniqueness if one imposes further conditions; see e.g.
+[BS93, SCH92, THU02].
+There are several kinds of conditions that one can impose.
+One of the simplest ways to define a unique circle packing for general simple planar maps is as follows.
+Let $\mathcal{M}$ be a 2-connected simple planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$.
+For each face $f\in\mathcal{M}$, we add a vertex $v_{f}$ and add edges to connect $v_{f}$ to the vertices incident to $f$.
+This turns $\mathcal{M}$ into a simple infinite plane triangulation $\mathcal{T}_{\mathcal{M}}$.
+We can consider the circle packing for $\mathcal{T}_{\mathcal{M}}$ and remove the disks for all the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$.
+This defines a circle packing for $\mathcal{M}$ and it is the unique, up to Möbius transform, circle packing constructed in this way.
+Let $\mathcal{H}$ be a cell configuration such that the associated map $\mathcal{M}$ is 2-connected and has whole-plane topology.
+Let $\mathcal{T}_{\mathcal{M}}$ be the triangulation generated from $\mathcal{M}$ as above.
+We write $v_{H}\sim f$ if the vertex $v_{H}$ for $H$ in the associated map $\mathcal{M}$ is incident to the face $f\in\mathcal{F}\mathcal{M}$.
+Let
+$\overline{H}_{0}=\bigcup_{f\in\mathcal{F}\mathcal{M}:v_{H_{0}}\sim f}\bigcup_{H\in\mathcal{H}:v_{H}\sim f}H$ (1.6)
+be the union of cells whose vertex is on the same face as $v_{H_{0}}$ and let $d_{0}$ be the maximal degree of the faces incident to $v_{H_{0}}$.
+Theorem 1.8.
+Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ being 2-connected and having whole-plane topology.
+Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4.
+Further assume
+$\mathbb{E}\Big[\frac{\mathrm{diam}(\overline{H}_{0})^{2}}{\mathrm{area}(H_{0})}\big(\deg(H_{0})+d_{0}\big)^{4}\deg(H_{0})\Big]<\infty,$ (1.7)
+Then the conclusions in Theorem 1.6 hold for the circle packing of $\mathcal{M}$ defined as above.
+In the setting of the Riemann uniformization, let $\mathcal{M}$ be a planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$.
+For each face $f$ enclosed by a single loop, we simply remove this loop.
+For each face $f\in\mathcal{F}\mathcal{M}$ with 2 edges, we contract the two edges to turn it into a single edge, such that each face of the resulting map has at least 3 edges.
+Let $M(\mathcal{M})$ be the Riemann surface obtained by associating each remaining face of degree $p\geq 3$ with regular $p$-gon of unit side length and glue them isometrically along their edges.
+Theorem 1.9.
+Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ having whole-plane topology.
+Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4.
+Further assume that (1.7) holds.
+Then the conclusions in Theorem 1.7 hold for the Riemann surface $M(\mathcal{M})$ defined as above.
+As discussed in [AHN+18, Section 6], there are several other natural ways to define Riemann surfaces associated with planar maps; for instance, one can assign each face of $\mathcal{M}$ of degree $p$ a disk of circumference $p$, with boundary split into $p$ arcs length one corresponding to the edges, and glue adjacent faces according to arc length along their shared edges.
+We expect that similar conclusions still hold by minor modifications of our proof of Theorem 1.9.
+We comment that the convergence results of random planar maps under discrete conformal embeddings in [GMS20, GMS21, HS23, BGS25] are all stated for finite planar maps, whereas Theorems 1.6-1.9 are stated for infinite plane triangulations.
+In our subsequent work, we will apply Theorems 1.6-1.9 to work on the convergence of random planar maps under circle packing and Riemann uniformization embedding in both finite volume and infinite volume setting.
+More precisely, we will use a version of Theorems 1.6-1.9 (Theorem 5.5) with a weaker version of the line connectivity in Definition 1.4, which will be proved in Section 5.3.
 1.4 Applications and future directions
 In this section, we discuss some applications and future directions for this work.
-We start with the following two simple examples, which provide applications of Theorems 1.7 and 1.8.
+We start with the following two simple examples, which provide applications of Theorems 1.6 and 1.7.
 We believe the results in these examples are new, except that a finite-volume version of Example 1.10 for circle packings was proven in [IM23].
@@
 An important motivation this work is that we, in future work, will apply the results of this paper to planar maps arising in the study of Liouville quantum gravity.
-Another future direction is to generalize our results to planar maps which are not triangulations.
-Let us discuss the latter direction in more detail.
-For infinite plane triangulations, the circle packing is unique.
-For general planar maps, the circle packing is not necessarily unique; however, one can get uniqueness if one imposes further conditions; see e.g.
-[BS93, SCH92, THU02].
-There are several different kinds of conditions that one can impose.
-One of the simplest ways to define a unique circle packing for general simple planar maps is as follows.
-Let $\mathcal{M}$ be a one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$.
-For each face $f\in\mathcal{M}$, we add a vertex $v_{f}$ and add edges to connect $v_{f}$ to the vertices incident to $f$.
-This turns $\mathcal{M}$ into a simple infinite plane triangulation $\mathcal{T}_{\mathcal{M}}$.
-We can consider the circle packing for $\mathcal{T}_{\mathcal{M}}$ and remove the disks for all the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$.
-This defines a circle packing for $\mathcal{M}$ and it is the unique, up to Möbius transform, circle packing constructed in this way.
-In Section 5.2, we will prove a variant of Theorem 1.7 for the version of the circle packing defined here.
-On the other hand, for the Riemann uniformization embedding one has existence and uniqueness of the embedding for any planar map.
-In Section 5.3, we will show that Theorem 1.8 still holds if we assume our planar map is a $p$-angulation for $p=4,5,6,\dots$.
-Problem 1.11.
-Prove Theorem 1.7 for other variants of circle packings for non-triangulations, and prove Theorem 1.8 for non-$p$-angulations.
-Regarding the Riemann uniformization embedding, we expect that general planar maps without loops can be treated by slightly modifying the argument of the current paper, possibly under a suitable probabilistic bound on the face degrees.
-The key step would be to prove that Lemma 4.2 still holds in this context.
-Furthermore, we believe it is also possible to treat maps with loops; in this case, it might additionally be necessary to give a different definition of $\phi_{0}$ in Section 4.2.
 1.5 Outline
-In this section we provide a moderately detailed overview of the content of the rest of the paper, as well as the outline of the proof of Theorems 1.7 and 1.8.
+In this section we provide a moderately detailed overview of the content of the rest of the paper, as well as the outline of the proof of Theorems 1.6 and 1.7.
 In Section 2, we present various preliminaries.
-In Section 2.1, we review various properties of circle packings and the Riemann surface associated with triangulations.
-In Section 2.2, gather some results on the ergodic averages of the ergodic modulo scaling cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2].
+In Section 2.1, we review various properties of circle packings and the Riemann surface associated with planar maps.
+In Section 2.2, we gather some results on the ergodic averages of the ergodic modulo scaling cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2].
 We will repeatedly use the following two facts: the maximum diameter of the cells which intersect a large origin-containing square is typically of smaller order than its side length (Lemma 2.6), and the sum of $\mathrm{diam}(H)^{2}\deg(H)^{4}$ over all the cells $H$ intersecting a large square $S$ which is not too far from origin is bounded from above by a deterministic constant times the area of $S$ (Lemma 2.7).
+As a consequence, similarly as in  [GMS22, Section 2.3], we start from a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ which serves as the (non-injective) a priori embedding of $M(\mathcal{H})$ and prove some of its topological properties.
+Much care needs to be taken with the exact definition of $\phi_{0}$ and we choose to define it by mapping each vertex of $M(\mathcal{H})$ to a uniform point in the corresponding cell of $\mathcal{H}$, and then extend to every equilateral triangle in $M(\mathcal{H})$ via linear interpolation.
+In Section 2.3, we recap the convergence of random walk on $\mathcal{H}$ to Brownian motion from [GMS22], and use it to prove that for each large square $S$, there is some planar embedding of $\mathcal{H}(S)$ that is close to $\mathcal{H}(S)$ (Proposition 2.15).
 In Section 3, we prove Theorem 1.6.
@@
 In Proposition 3.12 we prove that any limiting homeomorphism must be a linear transform.
-By the almost planarity assumption it follows that in order to conclude the proof, it is sufficient to prove that the random walk with Dubejko weights on the cell configuration $\mathcal{H}$ and on the circle packing $\mathcal{P}$ of $\mathcal{H}$ both converge to planar Brownian motion.
+By the almost planarity result from Proposition 2.15 it follows that in order to conclude the proof, it is sufficient to prove that the random walk with Dubejko weights on the cell configuration $\mathcal{H}$ and on the circle packing $\mathcal{P}$ of $\mathcal{H}$ both converge to planar Brownian motion.
 To prove that the random walk on $\mathcal{H}$ with Dubejko weights converges to Brownian motion, which is handled in Section 3.2, a natural first attempt is to try to directly use the convergence of random walks on $\mathcal{H}$ with generic conductance from [GMS22] (Theorem 2.12).
@@
 Proposition 3.6 is a variant of Theorem 2.12 from [GMS22] for the case where the conductances are given by the Dubejko weights and where the moment bound (2.10) has been replaced by (1.2).
-It remains to show that the random walk with Dubejko weights on the circle packing $\mathcal{P}$ converges to Brownian motion, which is carried out in Section $3.$ Following work by Bou-Rabee and Gwynne [BG24] it is sufficient to check that up to scaling, there are no macroscopic disks in $\mathcal{P}$.
+It remains to show that the random walk with Dubejko weights on the circle packing $\mathcal{P}$ converges to Brownian motion, which is carried out in Section 3.3.
+Following work by Bou-Rabee and Gwynne [BG24] it is sufficient to check that up to scaling, there are no macroscopic disks in $\mathcal{P}$.
 In [GJN19], Gurel-Gurevich, Jerison and Nachmias proved that the maximal diameter of the disks in a circle packing can be controlled by the inverse of Cannon’s vertex extremal length [CAN94].
@@
 parabolic, and in Lemma 4.9 we prove that the mapping between $\mathcal{H}$ and $\varphi(M(\mathcal{H}))$ has polynomial growth.
-In Section 4.2 we construct a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$.
-Similarly as in  [GMS22, Section 2.3], we start from a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ which serves as the (non-injective) a priori embedding of $M(\mathcal{H})$.
-Much care needs to be taken with the exact definition of $\phi_{0}$ and we choose to define it by mapping each vertex of $M(\mathcal{H})$ to a uniform point in the corresponding cell of $\mathcal{H}$, and then extend to every equilateral triangle in $M(\mathcal{H})$ via linear interpolation.
+In Section 4.2 we construct a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ using the function $\phi_{0}$ defined in Section 2.2.
 For each $m\in\mathbb{N}$ we let $\phi_{m}$ be equal to $\phi_{0}$ on the boundary of the $\phi_{0}^{-1}$ image of certain large squares, and harmonic elsewhere.
@@
 We first follow the same outline as [GMS22, Section 2.4] to prove that $\phi_{\infty}$ is close to $\phi_{0}$ on a set of line segments, where in our setting the line segments are contained in the image of a suitable subset of $M(\mathcal{H})$ under a conformal map.
-To prove the closeness of $\phi_{\infty}$ to $\phi_{0}$ on the complementary set, the oscillation of $\phi_{\infty}$ in each connected component of the complement is equal to the oscillation of $\phi_{\infty}$ on the boundary of that component thanks to the maximum principle for harmonic functions, whereas the oscillation of $\phi_{0}$ in each connected component can be bounded by the continuity from Lemma $4.$ This shows that $|\phi_{0}-\phi_{\infty}|$ is also small in every connected component of the complementary set, which concludes the proof of Proposition 4.20.
+To prove the closeness of $\phi_{\infty}$ to $\phi_{0}$ on the complementary set, the oscillation of $\phi_{\infty}$ in each connected component of the complement is equal to the oscillation of $\phi_{\infty}$ on the boundary of that component thanks to the maximum principle for harmonic functions, whereas the oscillation of $\phi_{0}$ in each connected component can be bounded by the continuity from Lemma 4.3.
+This shows that $|\phi_{0}-\phi_{\infty}|$ is also small in every connected component of the complementary set, which concludes the proof of Proposition 4.18.
 In Section 4.4, we conclude our proof of Theorem 1.7 by showing that $\phi_{\infty}$ has to be the composition of a deterministic linear transform, a random rotation and a scaling.
@@
 Furthermore, in Lemma 4.25, we prove that this polynomial must have degree 1 because if this is not the case, then one can find two cells that are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.2 and Lemma 4.5.
-In Lemmas 4.29,  4.30 and  4.32, we further use the ergodic theorem for the cell configuration (Lemma 2.4) to show that, this linear map $\phi_{\infty}\circ\varphi^{-1}$ can be written as a composition of a deterministic linear transform $\mathbf{A}$, a random rotation and a scaling, and in Lemma 4.33, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.6, then $\mathbf{A}$ must be a rotation.
+In Lemmas 4.27,  4.28 and  4.30, we further use the ergodic theorem for the cell configuration (Lemma 2.4) to show that, this linear map $\phi_{\infty}\circ\varphi^{-1}$ can be written as a composition of a deterministic linear transform $\mathbf{A}$, a random rotation and a scaling, and in Lemma 4.31, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.5, then $\mathbf{A}$ must be a rotation.
 This finishes the proof of Theorem 1.7.
 Our proof in Section 4.4 works under quite mild assumptions and gives a general method for proving that a random harmonic function is of a particular form (namely, the composition of a deterministic linear transformation and a random rotation and scaling) as long as it has polynomial growth, is almost injective, and the averages of the Dirichlet energy over large regions are approximately a deterministic constant.
-In Section 5.1, we prove Theorem 5.1, a version of Theorem 1.7 and Theorem 1.8 under a weaker version of the line connectivity property.
-In Section 5.2, we prove a version of Theorem 1.7 for general planar maps (non-triangulations); more specifically, we consider the variant of circle packings mentioned in Section 1.4.
-Finally, in Section 5.3 we prove the extension of Theorem 1.8 to $p$-angulations.
+In Section 5, we extend our results to general planar maps by proving Theorem 1.8 and Theorem 1.9.
+In Section 5.1, we prove Theorem 1.8, where we construct a cell configuration $\widetilde{\mathcal{H}}$ in the framework of Theorem 1.6 from $\mathcal{H}$.
+In Section 5.2 we prove Theorem 1.9.
+We first add a vertex on each edge and face of the planar map $\mathcal{M}$, and further add edges to turn $\mathcal{M}$ into an infinite plane triangulation.
+We define a different version of semi-flowers for the vertices and the added vertices in Lemma 4.1 and prove that under conformal maps, the ratio of the outradius and inradius of the semi-flowers can be bounded by the square of the degree, as in Lemma 5.4.
+We also split the cell configuration $\mathcal{H}$ to get a new cell configuration whose associated map is the new infinite triangulation.
+This together with the same arguments in Section 4 gives the proof of Theorem 1.9.
+Finally, in Section 5.3, we prove Theorem 5.5, a version of Theorems 1.6-1.9 under a weaker version of the line connectivity property.
 Acknowledgements.
@@
 2 Preliminaries
-2.1 Triangulations, circle packings and Riemann surfaces
-We follow the papers [BS01, AS03, GR13] for the definition on random plane triangulations.
-A planar map $\mathcal{M}$ is a proper (without edge-crossings) embedding of a connected planar (multi)graph $\mathcal{G}$ in the two-dimensional sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, where we identify two planar maps if one can be obtained from the other via an orientation-preserving homeomorphism of the sphere.
+2.1 Planar maps, circle packings and Riemann surfaces
+We follow the papers [BS01, AS03, GR13] for the definition of planar maps and triangulations.
+A planar map $\mathcal{M}$ is a proper (in particular, without edge-crossings) embedding of a connected planar (multi)graph $\mathcal{G}$ in the two-dimensional sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, where we identify two planar maps if one can be obtained from the other via an orientation-preserving homeomorphism of the sphere which maps vertices (resp.
+edges, faces) to vertices (resp.
+edges, faces).
+For convenience we abbreviate “equivalence class of embedded planar maps” to “planar map”.
 We will always assume that $\mathcal{M}$ is locally finite, i.e., each vertex and each face has finite degree.
+The face set $\mathcal{F}\mathcal{M}$ of $\mathcal{M}$ is the set of connected components of $\widehat{\mathbb{C}}\backslash\mathcal{G}$, and the support $S(\mathcal{M})\subseteq\widehat{\mathbb{C}}$ is defined to be the union of $\mathcal{M}$ and a subset of the faces in $\mathcal{F}\mathcal{M}$.
+In this paper we assume that all the faces in $S(\mathcal{M})$ are bounded in $\mathbb{C}$.
+We call a vertex $v\in\mathcal{V}\mathcal{M}$ an interior (resp.
+boundary) vertex of $\mathcal{M}$ if $v$ is in the interior (resp.
+on the boundary) of the support $S(\mathcal{M})$.
+We say $\mathcal{M}$ has whole-plane topology if the underlying graph is one-ended and $\mathcal{M}$ has no boundary vertices, and we say $\mathcal{M}$ has disk topology, if $\mathcal{M}$ is finite and the support $S(\mathcal{M})$ is homeomorphic to a closed disk.
+A rooted planar map $(\mathcal{M},e)$ is a planar map $\mathcal{M}$ together with an oriented edge $e=(x,y)$ of $\mathcal{M}$, called the root edge.
+The vertex $x$ is the root vertex.
 For a planar map $\mathcal{M}$ on the sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, we call a connected component of $\widehat{\mathbb{C}}\backslash\mathcal{M}$ a triangular face if its boundary is consisted of precisely three edges of $\mathcal{M}$.
-An embedded (planar) triangulation $\mathcal{T}$ consists of a finite or infinite connected planar map $\mathcal{M}$ in the sphere, together with a subset of the triangular faces, and its support $S(\mathcal{T})\subseteq\widehat{\mathbb{C}}$ is defined to be the union of $\mathcal{M}$ and all the triangular faces in $\mathcal{M}$.
-Two embedded triangulations $T$ and $T^{\prime}$ are considered equivalent if there is an orientation preserving homeomorphism between their supports which maps vertices (resp., edges, faces) to vertices (resp., edges faces).
-For convenience we abbreviate “equivalence class of embedded triangulations” to “triangulation”.
+An embedded (planar) triangulation $\mathcal{T}$ is a finite or infinite connected planar map $\mathcal{M}$ in the sphere, such that all the faces in the support $S(\mathcal{M})$ are triangular faces.
 Following the terminology in [AS03, Section 1.2], we call $\mathcal{T}$ a type II triangulation if $\mathcal{T}$ has no loops but possibly multiple edges, and call $\mathcal{T}$ a type III triangulation if $\mathcal{T}$ has no loops and no multiple edges.
-We call a vertex $v\in\mathcal{T}$ an interior (resp.
-boundary) vertex of $\mathcal{T}$ if $v$ is in the interior (resp.
-on the boundary) of the support $S(\mathcal{T})$.
 We call a type II (resp.
 type III) triangulation $\mathcal{T}$ an infinite plane triangulation (resp.
-simple infinite plane triangulation), if the underlying graph is one-ended, and there are no boundary vertices.
+simple infinite plane triangulation), if $\mathcal{T}$ is a planar map with whole-plane topology.
 We call a type II (resp.
 type III) triangulation $\mathcal{T}$ a (finite) disk triangulation (resp.
-simple disk triangulation), if the underlying planar map $\mathcal{M}$ is finite and the support $S(\mathcal{T})$ is homeomorphic to a disk.
-A rooted triangulation $(\mathcal{T},e)$ is a triangulation $\mathcal{T}$ together with an oriented edge $e=(x,y)$ of $\mathcal{T}$, called the root edge.
-The vertex $x$ is the root vertex.
+simple disk triangulation), if $\mathcal{T}$ is a planar map with disk topology.
 […]
@@
 Figure 2: Two adjacent vertices $u,v$.
-The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\big(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\big)$.
+The Dubejko conductance of the edge $uv$ is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\big(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\big)$.
 One can also consider random walks on circle packings.
 Consider the circle packing $\mathcal{P}$ of some simple infinite triangulation $\mathcal{T}$.
-Let $u,v$ be adjacent vertices in $\mathcal{T}$, and $w_{1},w_{2}$ be the vertices on the two faces of $\mathcal{T}$ containing the edge uvuv other than $u$ and $v$.
+Let $u,v$ be adjacent vertices in $\mathcal{T}$, and $w_{1},w_{2}$ be the vertices on the two faces of $\mathcal{T}$ containing the edge $uv$ other than $u$ and $v$.
 Let $o_{u},o_{v},o_{w_{1}},o_{w_{2}}$ be the centers of the circles for $u,v,w_{1},w_{2}$, and $o_{a},o_{b}$ be the centers of circles which are inscribed to the triangles $o_{u}o_{v}o_{w_{1}}$ and $o_{u}o_{v}o_{w_{2}}$, respectively.
-The Dubejko conductance of the edge uvuv is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}$; see Figure 2 for an illustration.
+The Dubejko conductance of the edge $uv$ is defined to be $\mathfrak{c}(u,v)=\frac{|o_{a}-o_{b}|}{|o_{u}-o_{v}|}$; see Figure 2 for an illustration.
 If we write $r_{u},r_{v},r_{w_{1}},r_{w_{2}}$ for the radii of the disks $D_{u},D_{v},D_{w_{1}},D_{w_{2}}$, then
 $\mathfrak{c}(u,v)=\frac{\sqrt{r_{u}r_{v}}}{r_{u}+r_{v}}\cdot\bigg(\sqrt{\frac{r_{w_{1}}}{r_{w_{1}}+r_{u}+r_{v}}}+\sqrt{\frac{r_{w_{2}}}{r_{w_{2}}+r_{u}+r_{v}}}\bigg).$ (2.1)
-One can infer from (2.1) that $\mathfrak{c}(u,v)<\frac{1}{2}$ for any edge uvuv using $2\sqrt{r_{u}r_{v}}\leq r_{u}+r_{v}$.
+One can infer from (2.1) that $\mathfrak{c}(u,v)<\frac{1}{2}$ for any edge $uv$ using $2\sqrt{r_{u}r_{v}}\leq r_{u}+r_{v}$.
 Furthermore, by Dubejko’s theorem (see e.g.
 [STE05, Theorem 18.3]), for a circle packing $\mathcal{P}$ of a simple infinite triangulation, if we consider the random walk $(Z_{n})_{n\geq 0}$ moving between the centers of the circles in $\mathcal{P}$, then $(Z_{n})_{n\geq 0}$ is a martingale.
-Next we follow the description of [GR13, Section 2.2] to define the Riemann surfaces associated with triangulations.
-Let $\mathcal{T}$ be a disk triangulation or an infinite plane triangulation.
-We associate $\mathcal{T}$ a metric space $(M(\mathcal{T}),d_{M})$, by regarding each triangle of $\mathcal{T}$ as an equilateral triangle of side length 1 and defining the distance between two points as the length of the shortest path joining them.
-Equivalently, $M(\mathcal{T})$ is obtained by gluing equilateral triangles according to their adjacency pattern in $\mathcal{T}$.
-The metric space $M(\mathcal{T})$ can be equipped with a compatible Riemann surface structure by defining coordinate charts as follows.
-If $x$ is an interior point of a triangle or interior to an edge on the boundary of the support, then we use this triangle, viewed as a subset of the plane, as the chart, and use the identity map as the projection.
-Similarly, if $x$ is interior to an edge and not on the boundary of the support, then we place two adjacent equilateral triangles in the plane and use the identity as the projection.
-Now let $x$ be an interior vertex of degree $n$, and let $\Delta_{1},...,\Delta_{n}$ be the equilaterals which meet at $x$ in cyclic order.
-We place them in the plane such that for j=1,,nj=1,...,n, $\Delta_{j}$ is the triangle with vertices $0,e^{\pi(j-1)i/3},e^{\pi ji/3}$, and define the coordinate map for $z=re^{i\theta}\in\Delta_{j}$ by $\phi(z)=r^{6/n}e^{6i\theta/n}$, using $\pi(j-1)/3<\theta<\pi j/3$ on $\Delta_{j}$.
-By further setting $\phi(0)=0$, this defines a homeomorphism between a neighborhood of $x$ in $M(\mathcal{T})$ and a neighborhood of 0 in $\mathbb{C}$.
-Moreover, for the atlas chosen as above, the transition map between two overlapping charts are analytic.
-The case where $x$ be an boundary vertex of degree $n$ can be treated similarly using the map $z\mapsto z^{3/(n-1)}$.
-This defines the Riemann surface structure for $M(\mathcal{T})$.
-With slight abuse of language, for a vertex $v\in\mathcal{V}\mathcal{T}$, we will also view $v$ as a point on the Riemann surface $M(\mathcal{T})$.
-Following the uniformization theorem, for an infinite plane triangulation $\mathcal{T}$, the surface $M(\mathcal{T})$ is either conformally equivalent to the complex plane and thus parabolic, or conformally equivalent to the unit disk and thus hyperbolic.
-As pointed out by [GR13], an equivalent characterization is that the Brownian motion on the surface $M(\mathcal{T})$ is either recurrent (in the parabolic case) or transient (in the hyperbolic case).
-A function $f:M(\mathcal{T})\to\mathbb{C}$ is harmonic if and only if $f\circ\varphi^{-1}$ is a harmonic function in usual sense, where $\varphi:M(\mathcal{T})\to\mathbb{C}$ is some conformal map.
+Next we define the Riemann surfaces associated with general planar maps, following the presentation in [AHN+18, Section 6].
+Let $\mathcal{M}$ be a planar map with whole-plane topology.
+For each face $f$ of $\mathcal{M}$ of degree 1, i.e., $f$ is enclosed by a single edge of $\mathcal{M}$, we remove this edge and the face from $\mathcal{M}$.
+For each face $f$ of $\mathcal{M}$ of degree 2, i.e., $f$ is formed by a double edge of $\mathcal{M}$, we collapse this double edge into a single edge.
+Now we assume that each face of $\mathcal{M}$ has degree at least 3.
+We associate to $\mathcal{M}$ a metric space $(M(\mathcal{M}),d_{M})$, by regarding each face of $\mathcal{M}$ of degree $p$ as regular $p$-gon of side length 1 and defining the distance between two points as the length of the shortest path joining them.
+Equivalently, $M(\mathcal{M})$ is obtained by gluing regular $p$-gons of unit side length according to their adjacency pattern in $\mathcal{M}$.
+The metric space $M(\mathcal{M})$ can be equipped with a compatible Riemann surface structure by defining coordinate charts as follows.
+For a face $f$, we use the regular polygon associated with $f$, viewed as a subset of the plane, as the chart, and use the identity map as the projection.
+For each edge $e$ of $\mathcal{M}$, we define an open neighborhood of the interior of $e$ in $M(\mathcal{M})$ by adding to $e$ the two triangles formed by the endpoints of $e$ and the centers of the two faces adjacent to $e$.
+To define a coordinate chart on this neighborhood, we simply place the two triangles next to each other in the plane.
+The reason to take only a subset of the faces and not the entire faces is that this chart is well defined even if both sides of $e$ are incident to the same face.
+For each vertex $v$ of $\mathcal{M}$, we define an open neighborhood of $v$ in $M(\mathcal{M})$ similarly by intersecting the corners of the faces adjacent to $v$ with open discs of radius 1/2 centered at $v$.
+We define a chart on this neighborhood by first laying the corners out sequentially around the origin (with possible overlapping), and then applying the function $z\mapsto z^{2\pi/\theta(v)}$, suitably interpreted to get an injective map into the plane.
+The radius is chosen so that this definition remains valid when multiple corners of the same face are located at the same vertex.
+We will also work with the setting where $\mathcal{M}$ is a disk triangulation; in this case, the boundary vertices of degree $n$ can be treated similarly using the map $z\mapsto z^{3/(n-1)}$.
+Following the uniformization theorem, for a planar map $\mathcal{M}$ with whole-plane topology, the surface $M(\mathcal{M})$ is either conformally equivalent to the complex plane and thus parabolic, or conformally equivalent to the unit disk and thus hyperbolic.
+As pointed out by [GR13, AHN+18], an equivalent characterization is that the Brownian motion on the surface $M(\mathcal{M})$ is either recurrent (in the parabolic case) or transient (in the hyperbolic case).
+A function $f:M(\mathcal{M})\to\mathbb{C}$ is harmonic if and only if $f\circ\varphi^{-1}$ is a harmonic function in usual sense, where $\varphi:M(\mathcal{M})\to\mathbb{C}$ is some conformal map.
 2.2 Ergodic averages of cell configurations using dyadic systems
@@
 As explained in [GMS22, Section 2.1], one can define a topology on the space of dyadic systems by looking at the local Hausdorff distance on the union of the origin-containing squares.
-Definition $2.$
+Definition 2.2.
 Let $s$ be uniformly sampled from $[0,1]$, and conditioned on $s$, we uniformly randomly sample $w$ in $0,2^{s}]\times[0,2^{s}]$.
@@
 ∎
+Figure 3: An illustration of the definition of $M(S;a)$ and $M(S)$.
+For simplicity here we assume that two cells are adjacent if they have nonempty intersection.
+The cells in cyan are those intersecting $\mathring{S}^{a}$, and their neighbors are in yellow.
+We also include the cells that are separated from infinity from the yellow and cyan cells, which are colored in green.
+The surface $M(S;a)\subseteq M(\mathcal{H})$ are formed by vertices for these cells on $M(\mathcal{H})$ as well as the edges connecting these vertices, and the surface $M(S)$ is the largest surface of the form $M(S;a)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$.
+For a square $S$, we define a disk triangulation $\mathcal{M}(S)$ using cells in $\mathcal{H}(S)$ as follows.
+Let $\mathcal{M}$ be the associated map of the cell configuration $\mathcal{H}$.
+For every $a>0$, let $\mathring{S}^{a}$ be the square concentric to $S$ with side length $a$.
+We work on the event that $\mathcal{H}(\mathring{S}^{a})$ is connected; otherwise we set $\mathcal{M}(S;a)=\emptyset$.
+Let $\mathcal{M}_{0}(S;a)\subseteq\mathcal{M}$ be the submap formed by the union of faces of $\mathcal{M}$ with at least one of the vertices in $\mathcal{H}(\mathring{S}^{a})$.
+Note that when viewed as a graph, $\mathcal{M}_{0}(S;a)$ is 2-connected since $\mathcal{H}(\mathring{S}^{a})$ is connected.
+In fact, one can verify that, for any infinite plane triangulation $\mathcal{T}$ and a connected subgraph $\mathcal{G}_{1}\subseteq\mathcal{G}$, if we let $\mathcal{G}_{2}$ be the subgraph $\mathcal{G}$ formed by all the vertices in $\mathcal{G}_{1}$ and their neighbors, as well as all the edges connecting these vertices, then $\mathcal{G}_{2}$ is 2-connected.
+We further let $\mathcal{M}(S;a)$ be the smallest simply connected submap of $\mathcal{M}$ containing $\mathcal{M}_{0}(S;a)$ as a submap.
+Then $\mathcal{M}(S;a)$ is a disk triangulation.
+Let $\mathcal{M}(S)$ be the largest submap of the form $\mathcal{M}(S;a)$ for some $a\in(0,|S|)$ such that for each vertex in $\mathcal{M}(S;a)$, its cell is contained in the interior of $S$.
+Note that these objects are well defined since $\mathcal{H}$ is locally finite.
+We further write $M(S)$ for the Riemann surface associated with $\mathcal{M}(S)$.
+See Figure 3 for an illustration.
+Lemma 2.8.
+Almost surely, for each given $\varepsilon>0$ and $a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $\mathcal{M}(S;(1-\varepsilon)|S|)\subseteq\mathcal{M}(S)$.
+Proof.
+[…]
+Let $\mathcal{H}$ be the cell configuration in the setting of Theorem 1.6 or Theorem 1.7.
+We introduce a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ below that will later be used in Section 2.3 and Section 4.
+The function $\phi_{0}$ sends each vertex on $M(\mathcal{H})$ to a point in the associated cell of $\mathcal{H}$ and is otherwise piecewise linear.
+More precisely, the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ is defined as follows.
+Given a realization of $\mathcal{H}$, for each cell $H\in\mathcal{H}$, we independently sample a point $\widetilde{c}(H)$ according to the probability measure proportional to the Lebsegue measure on $H$.
+For each triangular face $\Delta$ in $M(\mathcal{H})$ formed by the cells $H_{1},H_{2},H_{3}\in\mathcal{H}$, consider the triangle $\Delta^{\prime}$ in $\mathbb{C}$ formed by connecting $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ using straight line segments, and define $\phi_{0}|_{\Delta}:\Delta\to\Delta^{\prime}$ to be the unique linear function sending the vertices in $\Delta$ to their counterpart in $\Delta^{\prime}$.
+This defines a continuous and piecewise linear function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$.
+See Figure 4 for an illustration.
+Figure 4: An illustration for the definition of the map $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$, which linearly send each triangle on the left panel to the corresponding triangle on the right panel.
+Note that the $\phi_{0}$-images of different equilateral triangles of $M(\mathcal{H})$ might overlap with each other.
+We end this subsection with the following topological properties on $\phi_{0}$.
+Lemma 2.9.
+We have $\phi_{0}(M(\mathcal{H}))=\mathbb{C}$.
+Proof.
+Pick $k$ large enough such that by Definition 1.4, the 4 subgraphs of $\mathcal{H}$ formed by the cells intersecting the 4 sides of $S_{k}$ are connected.
+Since a cell intersecting a vertex of $S_{k}$ is contained in (at least) two of these subgraphs, we get that the set of cells intersecting a side of $S_{k}$ forms a connected graph.
+By further applying Lemma 2.6, we can find a chain of cells $H_{1}\sim H_{2}\sim\cdots\sim H_{n}\sim H_{1}$ in this connected graph such that if we connect $\widetilde{c}(H_{j})$ and $\widetilde{c}(H_{j+1})$ using straight line segments for j=1,,nj=1,...,n with $n+1$ identified with 1, then this creates a loop $\gamma_{n}$ surrounding the square $\widetilde{S}_{k}$ concentric to $S_{k}$ with side length $|S_{k}|/2$, and the winding number of $\gamma_{n}$ around each point in $\widetilde{S}_{k}$ is equal to 1.
+If we consider the vertices for $H_{1},...,H_{n}$ on the surface $M(\mathcal{H})$, then one can find a loop $\widetilde{\gamma}_{n}$ on $M(\mathcal{H})$ connecting these vertices with straight line segments such that $\phi_{0}(\widetilde{\gamma}_{n})=\gamma_{n}$.
+Since $M(\mathcal{H})$ is simply connected, we can continuously shrink $\widetilde{\gamma}_{n}$ to a single point $x$.
+If there exists a point $w\in\widetilde{S}_{k}\backslash\phi_{0}(M(\mathcal{H}))$, then under $\phi_{0}$ image, all of these shrinking loops will have winding number 1 around $w$.
+On the other hand, for loops sufficiently close to $x$, its image under $\phi_{0}$ has winding number 0 around $w$ since $w\notin\phi_{0}(M(\mathcal{H}))$.
+This is a contradiction, which implies that almost surely for large enough $k$, $\widetilde{S}_{k}\subseteq\phi_{0}(M(\mathcal{H}))$, which completes the proof.
+∎
+Lemma 2.10.
+For every compact set $K\subseteq\mathbb{C}$, $\phi_{0}^{-1}(K)$ has compact closure in $M(\mathcal{H})$.
+Proof.
+By Lemma 2.6, a.s.
+for all large enough $k$, any cells in $\mathcal{H}([-2^{k},2^{k}]^{2})$ has diameter less than $2^{k-10}$, and the convex hull of the union of any three cells $H_{1}\sim H_{2}\sim H_{3}$ in $\mathcal{H}([-2^{k},2^{k}]^{2}\backslash[-2^{k-2},2^{k-2}]^{2})$ does not intersect $K$.
+This gives the claim.
+∎
+Lemma 2.11.
+Almost surely for large enough $k$, the following is true.
+For any cells $H_{1},...,H_{n}\in\mathcal{H}(\partial[-2^{k},2^{k}]^{2})$ such that $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ form a simple loop $\gamma_{n}$ on $M(\mathcal{H})$ by drawing a straight line between the vertices representing adjacent cells, if $\phi_{0}(\gamma_{n})$ has winding number 1 around 0, then the loop $\gamma_{n}$ encloses $M([-2^{k},2^{k}]^{2})$.
+Proof.
+From the definition, $M([-2^{k},2^{k}]^{2})$ has disk topology and is disjoint from $\gamma_{n}$.
+Suppose $k$ is large enough such that by Lemma 2.10, $\phi_{0}^{-1}(0)\subseteq M([-2^{k-1},2^{k-1}]^{2})$.
+If $\gamma_{n}$ does not enclose $M([-2^{k},2^{k}]^{2})$, then one can continuously shrink $\gamma_{n}$ to a single point $x$ on $M(\mathcal{H})$ such that all the loops are disjoint from $M([-2^{k},2^{k}]^{2})$.
+In particular, under the image of $\phi_{0}$, all of these shrinking loops will have winding number 1 around 0.
+On the other hand, $x\notin M([-2^{k},2^{k}]^{2})$ and $\phi_{0}(x)\neq 0$.
+This is a contradiction.
+∎
+2.3 Convergence of random walks on cell configurations
+In order to prove Theorem 1.6 and Theorem 1.7, we will need that for each large square $S$, $\mathcal{H}(S)$ is close to some planar embedding of the map $\mathcal{M}(S)$.
+To achieve this, in Proposition 2.15 below, we will use the the convergence of random walks on cell configurations in [GMS22] to prove that for large squares $S$, $\mathcal{H}(S)$ is close to the Tutte embedding of $\mathcal{M}(S)$.
+Let $\beta_{1}:[0,T_{\beta_{1}}]\to\mathbb{C}$ and $\beta_{2}:[0,T_{\beta_{2}}]\to\mathbb{C}$ be continuous curves defined on possibly different time intervals. We define
+$\mathbb{d}^{\rm CMP}(\beta_{1},\beta_{2}):=\inf_{\phi}\sup_{t\in[0,T_{\beta_{1}}]}|\beta_{1}(t)-\beta_{2}(\phi(t))|,$ (2.8)
+where the infimum is over all increasing homeomorphisms $\phi:[0,T_{\beta_{1}}]\to[0,T_{\beta_{2}}]$ and “CMP” stands for “curves modulo parameterization”.
+For curves $\beta_{1}:[0,\infty)\to\mathbb{C}$ and $\beta_{2}:[0,\infty)\to\mathbb{C}$ defined for infinite time and $r>0$, we let $T_{1,r}$ (resp.
+$T_{2,r}$) be the first exit time of $\beta_{1}$ (resp.
+$\beta_{2}$) from the ball $B(0;r)$ (or 0 if the curve starts outside $B(0;r)$).
+We define
+$\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{1},\beta_{2}):=\int_{1}^{\infty}e^{-r}\min\{1,\mathbb{d}^{\rm CMP}(\beta_{1}|_{[0,T_{1,r}]},\beta_{2}|_{[0,T_{2,r}]})\}\,dr,$ (2.9)
+so that $\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{n},\beta)\to 0$ if and only if for Lebesgue a.e.
+$r>1$, $\beta_{n}$ stopped at its first exit time from $B(0;r)$ converges to $\beta$ stopped at its first exit time from $B(0;r)$ with respect to the metric (2.8).
+Theorem 2.12 (Theorem 3.10 of [GMS22]).
+Let $\mathcal{H}$ be a random cell configuration satisfying the conditions in Definition 1.3 and Definition 1.4. Further assume $\mathcal{H}$ satisfies the moment bound
+$\mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi(H_{0})\big)\bigg]<\infty,\ \ \mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi^{*}(H_{0})\big)\bigg]<\infty$ (2.10)
+where $\pi(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})$ and $\pi^{*}(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})^{-1}$.
+Then the simple random walk on $\mathcal{H}$ is a.s.
+recurrent.
+Moreover, for $z\in\mathbb{C}$, let $X^{z}$ denote the random walk on $\mathcal{H}$ started from $H_{z}$ with conductance $\mathfrak{c}$.
+For $j\in\mathbb{N}_{0}$, let $\widehat{X}_{j}^{z}$ be an arbitrarily chosen point of the cell $X_{j}^{z}$ and extend $\widehat{X}_{j}^{z}$ from $\mathbb{N}_{0}$ to $0,\infty)$ by piecewise linear interpolation.
+There is a deterministic covariance matrix $\Sigma$ with $\det\Sigma\neq 0$ such that the following is true.
+For each fixed compact set $A\subseteq\mathbb{C}$, it is a.s.
+the case that as $\varepsilon\to 0$, the maximum over all $z\in A$ of the Prokhorov distance between the conditional law of $\varepsilon\widehat{X}^{z/\varepsilon}$ given $\mathcal{H}$ and the law of Brownian motion started from $z$ with covariance matrix $\Sigma$, with respect to the topology on curves induced in (2.9), tends to 0.
+Corollary 2.13.
+[…]
+Proof.
+Write $\mathbf{R}_{\theta_{0}}$ for the matrix of rotating by angle $\theta_{0}$.
+One can directly check by computation that for a 2D Brownian motion $(X,Y)^{\mathrm{T}}$ with covariance matrix $\Sigma$, if $\mathbf{R}_{\theta_{0}}(X,Y)^{\mathrm{T}}$ is also a 2D Brownian motion with the same covariance matrix $\Sigma$, then $\Sigma$ must be a scalar times the identity matrix.
+This immediately implies Corollary 2.13.
+∎
+As studied in [GMS21], one main application of Theorem 2.12 is the convergence of random planar maps under the Tutte embedding.
+Let $\mathcal{M}_{0}$ be a disk triangulation and $v\in\mathcal{V}\mathcal{M}_{0}$ be an interior vertex.
+The Tutte embedding of $\mathcal{M}_{0}$ in $\mathbb{D}$ centered at $v$ is defined as follows.
+Enumerate the boundary vertices of $\mathcal{M}_{0}$ in counterclockwise order as $\{y_{1},...,y_{n}\}$ with $y_{1}$ chosen arbitrarily.
+For j=1,,nj=1,...,n we let $\mathfrak{p}(y_{j})$ be the probability that a simple random walk started from $v$ first hits the boundary at a vertex in the boundary arc $\{y_{1},...,y_{j}\}$.
+The boundary vertices $y_{1},...,y_{n}$ are then mapped in counterclockwise order to the unit circle $\partial\mathbb{D}$ such that $\psi(y_{j}):=e^{2\pi i\mathfrak{p}(y_{j})}$.
+Then the hitting probability of the random walk started from $v$ approximates the uniform measure on the unit circle.
+We map the interior vertices of $\mathcal{M}_{0}$ into the unit disk via the discrete harmonic extension of the boundary values, i.e., for any interior vertex $v$,
+$\psi(v)=\frac{1}{\#\{w\in\mathcal{V}\mathcal{M}_{0}:w\sim v\}}\sum_{w\sim v}\psi(w).$ (2.11)
+Let $\psi:\mathcal{V}\mathcal{M}_{0}\to\mathbb{D}$ be the discrete harmonic function with boundary data as prescribed above.
+For each edge $e$ of $\mathcal{M}_{0}$ connecting $v_{1}$ and $v_{2}$, we also define $\psi(e)$ to be the straight line segment joining $\psi(v_{1})$ and $\psi(v_{2})$.
+This defines the Tutte embedding [^1] Since $\mathcal{M}_{0}$ potentially has multiple edges, the Tutte embedding might not be a planar embedding as defined at the beginning of Section 1.1 since $\psi$ might be non-injective.
+On the other hand, the embedding in (2.12) below will be an actual planar embedding.
+of the map $\mathcal{M}_{0}$ within the unit disk $\overline{\mathbb{D}}$.
+Note that under this drawing, if $e_{1},...,e_{k}$ are edges with the same endpoints (i.e., these edges form a multiple edge), then $\psi(e_{1})=...=\psi(e_{k})$, and vertices enclosed by the edges $e_{1},...,e_{k}$ on the map $\mathcal{M}_{0}$ are mapped to points on the line segment $\psi(e_{1})$.
+Furthermore, the drawing $\psi$ is planar in the following sense.
+Let $e_{1}$ and $e_{2}$ be edges that are not contained within multiple edges of $\mathcal{M}_{0}$, i.e., they are not part of a multiple edge or are enclosed by one.
+If $e_{1}$ and $e_{2}$ have different endpoints, $\psi(e_{1})\cap\psi(e_{2})=\emptyset$, and if $e_{1}$ and $e_{2}$ has one common endpoint $v$, then $\psi(e_{1})\cap\psi(e_{2})=\psi(v)$.
+Next we prove that for large $k$, the embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ is close to some planar embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$.
+We will typically use the Tutte embedding here as the planar embedding.
+However, the embedded boundary of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ only has close Hausdorff distance, not curve distance (2.8), to $\partial[-2^{k},2^{k}]^{2}$, which is not sufficient.
+Below we will construct a map $\mathcal{M}_{k,\ell}$ which contains and is slightly larger than $\mathcal{M}([-2^{k},2^{k}]^{2})$ and work on that map instead.
+Recall the function $\phi_{0}$ defined after Lemma 2.8.
+Lemma 2.14.
+Let $\ell>0$. Almost surely, for all large enough $k$, there exists a submap $\mathcal{M}_{k,\ell}\subseteq\mathcal{M}$ with the following properties:
+  1. (i)
+    $\mathcal{M}_{k,\ell}$ has disk topology, and for each vertex $v$ on its boundary, its cell in $\mathcal{H}$ intersects $\partial[-2^{k},2^{k}]^{2}$;
+  2. (ii)
+    $\mathcal{M}_{k,\ell}$ contains $\mathcal{M}([-2^{k},2^{k}]^{2})$ as a submap;
+  3. (iii)
+    Let $M_{k,\ell}\subseteq M(\mathcal{H})$ be the Riemann surface generated from $\mathcal{M}_{k,\ell}$.
+Then the curve distance (2.8) between $\phi_{0}(\partial M_{k,\ell})$ and $\partial[-2^{k},2^{k}]^{2}$ is less than $2^{k-\ell}$.
+Proof.
+Following Definition 1.4 and Lemma 2.6, let $k$ be large enough such that for any line segment $L\subseteq[-2^{k+1},2^{k+1}]$ of length at least $2^{k-\ell-4}$, $\mathcal{H}(L)$ is connected, and the maximal diameter of cells in $\mathcal{H}([-2^{k+1},2^{k+1}]^{2})$ is less than $2^{k-2\ell-10}$.
+We partition $\partial[-2^{k},2^{k}]^{2}$ into line segments $L_{1},...,L_{2^{\ell+2}}$ of length $2^{k-\ell}$, such that $L_{1}=[-2^{k},-2^{k}+2^{k-\ell}]\times\{-2^{k}\}$, and $L_{1},...,L_{2^{\ell+2}}$ are aligned counterclockwise.
+For $j=1,...,2^{\ell+2}$ with $2^{\ell+2}+1$ identified with 1, let $\mathring{L_{j}}$ be the line segments concentric to $L_{j}$ with length $|L_{j}|/2$, and let $L_{j,j+1}$ be the union of $\mathring{L_{j}},\mathring{L}_{j+1}$ and the middle component of $(L_{j}\cup L_{j+1})\backslash(\mathring{L_{j}}\cup\mathring{L}_{j+1})$.
+We pick a cell $H_{j}\in\mathcal{H}(\mathring{L_{j}})$.
+By the line connectivity property above, we can find a simple path $\widetilde{\gamma}_{k,j}$ on $\mathcal{H}(L_{j,j+1})$ connecting $H_{j}$ and $H_{j+1}$.
+The intersection between $\widetilde{\gamma}_{k,j}$ and $\widetilde{\gamma}_{k,j+1}$ are in $\mathcal{H}(\mathring{L}_{j+1})$, and therefore we can erase all the loops in $\cup_{j}\widetilde{\gamma}_{k,j}$ that are made by vertices in $\mathcal{H}(\mathring{L}_{j+1})$ for each $j$ to obtain a simple loop $\gamma_{k}$.
+Let $\mathcal{M}_{k,\ell}\subseteq\mathcal{M}$ be the planar map formed by $\gamma_{k}$ and the vertices enclosed by $\gamma_{k}$.
+Then properties (i) and (iii) are immediate from construction, while property (ii) is a consequence of Lemma 2.11.
+∎
+We are now ready to prove that the embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ is close to planar.
+Proposition 2.15.
+Let $\mathcal{H}$ be a cell configuration in the context of Theorem 1.6 or Theorem 1.7.
+There exists a collection of points $\{z_{H}^{k}:H\in\mathcal{V}\mathcal{M}([-2^{k},2^{k}]^{2})\}$ and simple curves $\{\gamma_{e}^{k}:e\in\mathcal{E}\mathcal{M}([-2^{k},2^{k}]^{2})\}$ forming a planar embedding of $\mathcal{M}([-2^{k},2^{k}])$, such that almost surely,
+$\frac{1}{2^{k}}\max_{H\in\mathcal{V}\mathcal{M}([-2^{k},2^{k}]^{2})}\max_{e}\operatorname{dist}(\gamma_{e},H)\to 0,$ (2.12)
+where the second maximum is over edges with $H$ being an endpoint and the distance is the Hausdorff distance.
+Proof.
+By setting the conductance to be equal to 1 for all edges, the moment bound (2.6) implies the moment bound (2.10), so the conclusion in Theorem 2.12 holds.
+Without loss of generality assume that the matrix $\Sigma$ there is the identity matrix; otherwise we choose a deterministic matrix $\mathbf{A}$ such that the random walk on $\mathbf{A}\mathcal{H}$ converges to the standard planar Brownian motion in the sense of Theorem 2.12.
+Let $k,\ell>0$, the map $\mathcal{M}_{k,\ell}$ be as in Lemma 2.14, and $L_{1},...,L_{2^{\ell+2}}$ be as in the proof of Lemma 2.14.
+Let $X^{z}$ and $\widehat{X}^{z}$ be the random walks in Theorem 2.12 started from a cell in the interior of $\mathcal{M}_{k,\ell}$, and let $\widehat{H}$ be the cell where the walk $X^{z}$ first hits $\partial\mathcal{M}_{k,\ell}$.
+From our construction, if $\widehat{H}$ is on $\mathcal{H}(L_{1}\cup...\cup L_{j})$, then the walk $\widehat{X}^{z}$ hits the $2^{k-\ell+1}$-neighborhood of $L_{1}\cup...\cup L_{j}$ before hitting $L_{j+2}\cup...\cup L_{2^{\ell+2}-1}$.
+Likewise, if $\widehat{H}$ is on $\mathcal{H}(L_{j+1}\cup...\cup L_{2^{\ell+2}})$, then the walk $\widehat{X}^{z}$ hits the $2^{k-\ell+1}$-neighborhood of $L_{j+1}\cup...\cup L_{2^{\ell+2}}$ before hitting $L_{1}\cup...\cup L_{j-1}$.
+Further applying Theorem 2.12, a.s.
+we send $\ell\to\infty$ sufficiently slower than $k\to\infty$, for each $a\in(0,1)$ and $z\in[-a2^{k},a2^{k}]^{2}$, the conditional probability given $\mathcal{H}$ of the walk $X^{z}$ leaves $\mathcal{M}_{k,\ell}$ at a vertex on $\mathcal{H}(L_{1}\cup...\cup L_{j})$ is close to the harmonic measure on $\partial[-2^{k},2^{k}]^{2}$ as seen from $z$ of $L_{1}\cup...\cup L_{j}$.
+Therefore when embedded into $\mathbb{D}$, by the conformal invariance of Brownian motion, if we take a conformal map $\varphi_{0}:[-1,1]^{2}\to\mathbb{D}$ fixing 0, then up to a rotation, $2^{-k}\varphi_{0}(H)$ is close to the vertex for $H$ under the Tutte embedding of $\mathcal{M}_{k,\ell}$ centered at $H_{0}$.
+We can further consider the image of the Tutte embedding for $\mathcal{M}_{k,\ell}$ under $2^{k}\varphi_{0}^{-1}$ to generate the points and curves $\{\widetilde{z}_{H}^{k}\}$ and curves $\{\widetilde{\gamma}_{e}\}$ where (2.12) holds, and further separate the curves $\widetilde{\gamma}_{e}$ a little for multiple edges and embed the components contained within these multiple edges in an arbitrary way to get an actual planar embedding where (2.12) holds.
 3 Convergence of cell configurations under the circle packing
@@
 This will eventually allow us to verify only polynomial bounds of degree (rather than exponential bounds) when applying (our variant of) the invariance principle in [GMS22].
-Lemma $3.$
+Lemma 3.3.
 For any flower $(D_{0};D_{1},...,D_{d})$ with $d$ petals and radii $(r_{0};r_{1},...,r_{d})$,
@@
 In this section, we prove that, for cell configurations $\mathcal{H}$ in the context of Theorem 1.6, if we assign the conductance to be equal to the Dubejko weights, then the random walk on $\mathcal{H}$ with the Dubejko weights converges to Brownian motion modulo time parameterization.
-We begin with the convergence of random walks on cell configurations in [GMS22].
-Let $\beta_{1}:[0,T_{\beta_{1}}]\to\mathbb{C}$ and $\beta_{2}:[0,T_{\beta_{2}}]\to\mathbb{C}$ be continuous curves defined on possibly different time intervals.
-We define
-$\mathbb{d}^{\rm CMP}(\beta_{1},\beta_{2}):=\inf_{\phi}\sup_{t\in[0,T_{\beta_{1}}]}|\beta_{1}(t)-\beta_{2}(\phi(t))|,$ (3.6)
-where the infimum is over all increasing homeomorphisms $\phi:[0,T_{\beta_{1}}]\to[0,T_{\beta_{2}}]$ and “CMP” stands for “curves modulo parameterization”.
-For curves $\beta_{1}:[0,\infty)\to\mathbb{C}$ and $\beta_{2}:[0,\infty)\to\mathbb{C}$ defined for infinite time and $r>0$, we let $T_{1,r}$ (resp.
-$T_{2,r}$) be the first exit time of $\beta_{1}$ (resp.
-$\beta_{2}$) from the ball $B(0;r)$ (or 0 if the curve starts outside $B(0;r)$).
-We define
-$\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{1},\beta_{2}):=\int_{1}^{\infty}e^{-r}\min\{1,\mathbb{d}^{\rm CMP}(\beta_{1}|_{[0,T_{1,r}]},\beta_{2}|_{[0,T_{2,r}]})\}\,dr,$ (3.7)
-so that $\mathbb{d}^{\rm CMP}_{\rm loc}(\beta_{n},\beta)\to 0$ if and only if for Lebesgue a.e.
-$r>1$, $\beta_{n}$ stopped at its first exit time from $B(0;r)$ converges to $\beta$ stopped at its first exit time from $B(0;r)$ with respect to the metric (3.6).
-Theorem 3.5 (Theorem 3.10 of [GMS22]).
-Let $\mathcal{H}$ be a random cell configuration satisfying the conditions in Definition 1.3 and Definition 1.4. Further assume $\mathcal{H}$ satisfies the moment bound
-$\mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi(H_{0})\big)\bigg]<\infty,\ \ \mathbb{E}\bigg[\frac{\mathrm{diam}(H_{0})^{2}}{\mathrm{area}(H_{0})}\pi^{*}(H_{0})\big)\bigg]<\infty$ (3.8)
-where $\pi(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})$ and $\pi^{*}(H)=\sum_{H^{\prime}\sim H}\mathfrak{c}(H,H^{\prime})^{-1}$.
-Then the simple random walk on $\mathcal{H}$ is a.s.
-recurrent.
-Moreover, for $z\in\mathbb{C}$, let $X^{z}$ denote the random walk on $\mathcal{H}$ started from $H_{z}$ with conductance $\mathfrak{c}$.
-For $j\in\mathbb{N}_{0}$, let $\widehat{X}_{j}^{z}$ be an arbitrarily chosen point of the cell $X_{j}^{z}$ and extend $\widehat{X}_{j}^{z}$ from $\mathbb{N}_{0}$ to $0,\infty)$ by piecewise linear interpolation.
-There is a deterministic covariance matrix $\Sigma$ with $\det\Sigma\neq 0$ such that the following is true.
-For each fixed compact set $A\subseteq\mathbb{C}$, it is a.s.
-the case that as $\varepsilon\to 0$, the maximum over all $z\in A$ of the Prokhorov distance between the conditional law of $\varepsilon\widehat{X}^{z/\varepsilon}$ given $\mathcal{H}$ and the law of Brownian motion started from $z$ with covariance matrix $\Sigma$, with respect to the topology on curves induced in (3.7), tends to 0.
-Corollary 3.6.
-[…]
-Proof.
-Write $\mathbf{R}_{\theta_{0}}$ for the matrix of rotating by angle $\theta_{0}$.
-One can directly check by computation that for a 2D Brownian motion $(X,Y)^{\mathrm{T}}$ with covariance matrix $\Sigma$, if $\mathbf{R}_{\theta_{0}}(X,Y)^{\mathrm{T}}$ is also a 2D Brownian motion with the same covariance matrix $\Sigma$, then $\Sigma$ must be a scalar times the identity matrix.
-This immediately implies Corollary 3.6.
-∎
+We start with the following immediate consequence of Theorem 2.12 and Theorem 2.1.
 Corollary 3.5.
@@
 Recall the notion of vertex extremal length introduced in Section 3.3.1.
-Let $\widehat{\mathcal{H}}$ be a planar embedding of the associated map of $\mathcal{H}$ that is close to $\mathcal{H}$ in the sense of Definition 1.5, with vertex set $\{z_{H}:H\in\mathcal{H}\}$.
-For each $H\in\mathcal{H}$, let $\Gamma_{H}$ be the collection of finite paths on $\widehat{\mathcal{H}}$ starting from $z_{H}$ and making a simple loop around $z_{H_{0}}$.
+For each $H\in\mathcal{H}$, let $\Gamma_{H}$ be the collection of finite paths on $\mathcal{M}$ starting from $H$ and making a simple loop around $H_{0}$}.
 Proposition 3.11.
@@
 Without loss of generality assume $q_{0}>300$ and set $q=q_{0}/72$.
-We will adapt the proof of [GJN19, Theorem 1.4] together with the input from Section $2.$ For $c>0$ and any square $S$, define the event $E_{1}(c,S)$ where
+We will adapt the proof of [GJN19, Theorem 1.4] together with the input from Section 2.2.
+For $c>0$ and any square $S$, define the event $E_{1}(c,S)$ where
 $\sum_{H\in\mathcal{H}(S)}\mathrm{diam}(H)^{2}\leq c|S|^{2}.$ (3.18)
@@
 On the other hand, using the definition of $E_{k,2}(\varepsilon)$, $H^{*}\subseteq S_{1}^{*}$ and $H_{0}\cap S_{n}^{*}=\emptyset$.
-For each path $\ell\in\Gamma_{H^{*}}$, using $H\cap H^{\prime}\neq\emptyset$ for $H\sim H^{\prime}$ for all large enough $k$, $\cup_{H\in\ell}H$ must cross all the regions $S_{j}^{*}\backslash S_{j-1}^{*}$, which implies
+For each path $\ell\in\Gamma_{H^{*}}$, using $H\cap H^{\prime}\neq\emptyset$ for $H\sim H^{\prime}$ for all large enough $k$, together with the almost planarity condition for $\mathcal{M}([-2^{k+3},2^{k+3}]^{2})$ from Proposition 2.15, $\cup_{H\in\ell}H$ must cross all the regions $S_{j}^{*}\backslash S_{j-1}^{*}$, which implies
 $\sum_{H\subseteq\ell\cap\mathcal{H}({S_{j}^{*}\backslash S_{j-1}^{*}})}m^{*}(H)=|S_{j}^{*}|^{-1}\sum_{H\subseteq\ell\cap\mathcal{H}({S_{j}^{*}\backslash S_{j-1}^{*}})}\mathrm{diam}(H)\geq|S_{j}^{*}|^{-1}(\frac{1}{3}|S_{j}^{*}|-2\cdot 2^{-32n}|S_{k}|)\geq\frac{1}{6},$ (3.20)
@@
 In the rest of this section, we use the following notation.
-For a continuous-time walk $(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a bounded set $A\subseteq\mathbb{C}$ where $z\notin A$, we write $\mathbb{P}^{z}$ for the law of $(X_{t})_{t\geq 0}$ started from $z$, and $\tau_{A}$ for the first time $T$ when the trace of $(X_{t})_{0\leq t\leq T}$ separates AA from infinity.
+For a continuous-time walk $(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a bounded set $A\subseteq\mathbb{C}$ where $z\notin A$, we write $\mathbb{P}^{z}$ for the law of $(X_{t})_{t\geq 0}$ started from $z$, and $\tau_{A}$ for the first time $T$ when the trace of $(X_{t})_{0\leq t\leq T}$ separates $A$ from infinity.
 We write $\tau_{w}$ instead of $\tau_{\{w\}}$ if $A=\{w\}$.
+For $\ell>10$, we write $\sigma_{\ell}$ for the first time when the walk exits $2^{\ell-1},2^{\ell-1}]^{2}$.
 Proposition 3.12.
@@
 […]
-Let $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ be the linearly interpolated random walk on $\widetilde{\varepsilon}\mathcal{P}$ described in Proposition 3.9, and let $\widehat{X}^{z,\varepsilon}$ be interpolated random walk on $\varepsilon\mathbf{A}\mathcal{H}$ as in Theorem 3.5 using the points $\{\mathbf{A}z_{H}:H\in\mathcal{H}\}$, except that we interpolate along the edges $\{\mathbf{A}\gamma_{H,H^{\prime}}:H\sim H^{\prime}\}$.
-Note that $f^{\varepsilon}(\widehat{X}^{z,\varepsilon})$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ can be coupled to have the same trace.
+Let $\widehat{X}^{z,\mathrm{circle},k,\ell}$ be the linearly interpolated random walk on $\widetilde{\varepsilon}_{k}\mathcal{P}$ described in Proposition 3.7, and let $\widehat{X}^{z,k,\ell}$ be interpolated random walk on $2^{-k}\mathbf{A}\mathcal{H}$ as in Theorem 2.12, except that for $H\in\mathcal{V}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})$ we use the points $\{\mathbf{A}z_{H}^{k+\ell}\}$ and for $e\in\mathcal{E}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})$ we interpolate along the edges $\{\mathbf{A}\gamma_{e}^{k+\ell}:e\in\mathcal{E}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})\}$.Note that by Lemma 2.8, almost surely for large enough $k$, the traces of $f^{k}(\widehat{X}^{z,k,\ell})$ and $\widehat{X}^{z,\mathrm{circle},k,\ell}$ within $f^{k,\ell}([-2^{\ell-1},2^{\ell-1}]^{2})$ can be coupled to be the same.
 Indeed, this coupling exists since $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ can be viewed as two embeddings of the associated map for $\mathcal{H}$ and have identical conductance, and we work on this coupling.
-Then by Proposition 3.8 and Proposition 3.9, as $\varepsilon\to 0$ the conditional law of $\widehat{X}^{z,\varepsilon}$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ given $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ converges to a constant multiple of planar Brownian motion with respect to the topology on curves induced in (3.7).
-Throughout the proof, we write $\mathbb{P}_{\rm BM}$ for the law of standard Brownian motion, $\mathbb{P}_{\varepsilon}$ for the conditional law of $\widehat{X}^{\varepsilon}$ given $\varepsilon\mathcal{H}$, and $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}$ for the conditional law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ given $\widetilde{\varepsilon}\mathcal{P}$.
-Also let $K$ be a fixed compact subset of $\mathbb{C}$ containing 0 and 1, and $d_{K}=\mathrm{diam}(K)$.
-We will verify that $\{f^{\varepsilon}\}$ is uniformly bounded and equicontinuous on compacts.
-Uniform boundedness on compact sets.
-For $z\neq 0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity.
+Then by Proposition 3.6 and Proposition 3.7, as $k\to\infty$ the conditional law of $\widehat{X}^{z,k,\ell}$ and $\widehat{X}^{z,\mathrm{circle},k,\ell}$ given $\varepsilon_{k}\mathcal{H}$ and $\widetilde{\varepsilon}_{k}\mathcal{P}$ converges to a constant multiple of planar Brownian motion with respect to the topology on curves induced in (2.9).
+Throughout the proof, we write $\mathbb{P}_{\rm BM}$ for the law of standard Brownian motion, $\mathbb{P}_{k,\ell}$ for the conditional law of $\widehat{X}^{k,\ell}$ given $\varepsilon\mathcal{H}$, and $\widetilde{\mathbb{P}}_{k,\ell}$ for the conditional law of $\widehat{X}^{\mathrm{circle},k,\ell}$ given $\widetilde{\varepsilon}_{k}\mathcal{P}$.
+Fix $\widetilde{\ell}>10$, let $K=[-2^{\widetilde{\ell}-5},2^{\widetilde{\ell}-5}]$ and $d_{K}=2^{\widetilde{\ell}-4}$.
+We will verify that $\{f^{k,\ell}{\,:\,k\in\mathbb{N},\ell>\widetilde{\ell}+1}\}$ is uniformly bounded and equicontinuous on $K$ for any fixed $\ell$, and we will later send $\widetilde{\ell}\to\infty$.
+Step 1: Uniform boundedness on $K$.
+For $z\neq 0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}<\sigma_{\ell}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity and has not exited $B(0;2^{\ell-1}$).
 Let $\gamma_{K}$ be a deterministic curve formed by the union of the line segment from 0 to $(2d_{K}+3)i$ and a curve from $(2d_{K}+3)i$ to $2+(d_{K}+2)i$, such that it is disjoint from $B(0;d_{K}+1)$ and separates $B(0;{d_{K}+1})$ from infinity.
 Then for any $\delta>0$, with positive probability, the planar Brownian motion started from 0 stays within the $\delta/2$-neighborhood of $\gamma_{K}$ and form a loop around $K\cup\{1\}\subset B(0;d_{K}+1)$ during time $[0,2]$.
-In particular, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}\leq\tau_{1}]$ is uniformly positive for $z\in K$ and small enough $\varepsilon$.
-Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, for each $z\in K$, there exists $\varepsilon^{\prime}>0$ such that $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}]>3\varepsilon^{\prime}$ for $\varepsilon$ sufficiently small.
+In particular, a.s.
+for fixed $\ell$ and large enough $k$, $\mathbb{P}_{k,\ell}^{0}\big[\{\tau_{z}\leq\tau_{1}\}\cap\{\sup_{0\leq t\leq\tau_{1}}|\widehat{X}_{t}|\leq 2^{\ell-1}\}\big]$ is uniformly positive for $z\in K$.
+Since $f^{\varepsilon}(\widehat{X}^{k,\ell})$ and $\widehat{X}^{\mathrm{circle},k,\ell}$ are coupled to have the same trace until leaving $f^{k,\ell}([-2^{\ell-1},2^{\ell-1}]^{2})$ for each $z\in K$, there exists $\varepsilon^{\prime}>0$ such that $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}\leq\tau_{1}]>3\varepsilon^{\prime}$ for $k$ sufficiently large.
 On the other hand, for each $\varepsilon^{\prime}>0$, there exists $M,\delta_{1}>0$, such that the probability that a Brownian motion from 0 separates $B(1;\delta_{1})$ from infinity before leaving $B(0;M)$ is at least $1-\varepsilon^{\prime}$.
-Using the convergence in law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ to standard Brownian motion, for $\varepsilon$ sufficiently small, with probability at least $1-2\varepsilon^{\prime}$, the trace of $\widehat{X}^{0,\rm circle,\widetilde{\varepsilon}}$ separates 1 from infinity before leaving $B(0;2M)$.
-In other words, for $\varepsilon$ sufficiently small, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}\leq\tau_{1}]\leq 2\varepsilon^{\prime}$ for $|z|>2M$.
-Thus $|f^{\varepsilon}(z)|\leq 2M$ for all $z\in K$ and small enough $\varepsilon$, which gives the uniform boundedness for $f^{\varepsilon}$ on compacts.
-Equicontinuity near 0 and 1.
+Using the convergence in law of $\widehat{X}^{\mathrm{circle},k,\ell}$ to standard Brownian motion, for $k$ sufficiently large, with probability at least $1-2\varepsilon^{\prime}$, the trace of $\widehat{X}^{0,\mathrm{circle},k,\ell}$ separates 1 from infinity before leaving $B(0;2M)$.
+In other words, for $k$ sufficiently large, $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}\leq\tau_{1}]\leq 2\varepsilon^{\prime}$ for $|z|>2M$.
+Thus $|f^{k,\ell}(z)|\leq 2M$ for all $z\in K$, large enough $k$, and $\ell>\widetilde{\ell}+1$.
+By a similar argument, for each compact set $K^{\prime}\subseteq\mathbb{C}$, there exists $\varepsilon^{\prime\prime}>0$ such that for every $\ell$ and large enough $k$, $\inf_{z^{\prime}\in K^{\prime}}\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z^{\prime}}\leq\tau_{1}]>\varepsilon^{\prime\prime}$, and for $\ell$ large enough and every $k$ large enough, $\mathbb{P}_{k,\ell}^{0}[\tau_{1}<\sigma_{\ell-8}]>1-\varepsilon^{\prime\prime}$.
+The latter inequality implies that for large enough $\ell$ and $k$, $\inf_{z^{\prime}\notin f^{k,\ell}(K)}\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z^{\prime}}\leq\tau_{1}]>1-\varepsilon^{\prime\prime}$, which further implies that almost surely, there exists some $\ell_{0}$, such that for all $\ell>\ell_{0}$, there is a $k_{\ell}>0$ such that $K^{\prime}\subseteq f^{k,\ell}(K)$ for all $k>k_{\ell}$.
+Step 2: Equicontinuity near 0.
 Let $\delta>0$.
 For $z\neq 0,1$, we work on the event $\{\tau_{z}=\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity at the same time.
-Using the same $\gamma_{K}$ and Brownian motion argument as in the previous step, one can see that there exists $\delta^{\prime}>0$ such that for $\varepsilon$ small enough $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}=\tau_{1}]>\delta^{\prime}$ when $z\in K$ and $|z|>\delta$.
-On the other hand, there exists $\delta^{\prime\prime}>0$ such that the probability that a planar Brownian motion separates $B(0;2\delta^{\prime\prime})$ from infinity before leaving $B(0;1/2)$ is at least $1-\delta^{\prime}/2$, and thus using the convergence in law of $\widehat{X}^{\varepsilon}$ to Brownian motion in local uniform topology, for $|z|<\delta^{\prime\prime}$ and all sufficiently small $\varepsilon$, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}=\tau_{1}]<\delta^{\prime}$.
-Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, we conclude that there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)|<\delta$.
+Using the same $\gamma_{K}$ and Brownian motion argument as in the previous step, one can see that there exists $\delta^{\prime}>0$ such that for $k$ large enough $\mathbb{P}_{k,\ell}^{0}[\tau_{z}=\tau_{1}<\sigma_{\ell}]>\delta^{\prime}$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}=\tau_{1}]>\delta^{\prime}$ when $z\in K$ and $|z|>\delta$.
+On the other hand, there exists $\delta^{\prime\prime}>0$ such that the probability that a planar Brownian motion separates $B(0;2\delta^{\prime\prime})$ from infinity before leaving $B(0;1/2)$ is at least $1-\delta^{\prime}/2$, and thus using the convergence in law of $\widehat{X}^{\varepsilon}$ to Brownian motion in local uniform topology, for $|z|<\delta^{\prime\prime}$ and all sufficiently large $k$, $\mathbb{P}_{k,\ell}^{0}[\tau_{z}<\tau_{1};\tau_{z}<\sigma_{\ell}]>1-\delta^{\prime}$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}<\tau_{1}]>1-\delta^{\prime}$.
+Since $f^{k,\ell}(\widehat{X}^{k,\ell})$ and $\widehat{X}^{\mathrm{circle},k,\ell}$ are coupled to have the same trace until leaving $f^{k,\ell}(K)$, we have a.s.
+for large $k$, $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}=\tau_{1}]>\delta^{\prime}$ for $|z|>\delta$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}<\tau_{1}]>1-\delta^{\prime}$ for $|z|<\delta^{\prime\prime}$.
+We conclude that there a.s.
+exists $k_{0}$, such that for any $k>k_{0}$, if $|z|<\delta^{\prime\prime}$ then $|f^{k,\ell}(z)|<\delta$, and if $|z|\geq\delta$ then $|f^{k,\ell}(z)|\geq\delta^{\prime\prime}$.
 This verifies the equicontinuity near 0.
-The equicontinuity near 1 can be verified similarly by considering walks started at 1.
-Equicontinuity at other points in $K$.
-Fix $z\in K$ with $z\neq 0,1$.
-For $z^{\prime}\neq z,0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}\}$, i.e., when the path separates 1 from infinity, it separates $z$ from infinity but does not separate $z^{\prime}$ from infinity.
-From the previous cases, $\{f^{\varepsilon}(z)\}$ is uniformly bounded away from 0, 1 and $\infty$.
-Fix $0<\delta<\frac{1}{2}\min_{\varepsilon>0}\{1,|f^{\varepsilon}(z)|,|f^{\varepsilon}(z)-1|\}$.
-Let $\gamma_{f^{\varepsilon}(z)}$ be a finite path starting from 0 that first separates $B(f^{\varepsilon}(z);\delta/4)$ and then separates $B(1;\delta/4)$ from infinity, but does not separate any points in $\mathbb{C}\backslash(B(f^{\varepsilon}(z);{\delta/2})\cup B(1;\delta/2))$ from infinity.
-Then using a similar argument in previous cases, where the Brownian motion has positive probability to stay sufficiently close to $\gamma_{f^{\varepsilon}(z)}$, and the random walks converges in law to Brownian motion in local uniform topology, we see that for each $\delta>0$ there exists $\delta^{\prime},\delta^{\prime\prime}>0$ such that for sufficiently small $\varepsilon$, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}<\tau_{z^{\prime}}]>\delta^{\prime}$ if $|z^{\prime}-f^{\varepsilon}(z)|>\delta$ whereas $\mathbb{P}_{\varepsilon}[\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}]<\delta^{\prime}$ if $|z^{\prime}-z|<\delta^{\prime\prime}$.
-In particular, there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z-z^{\prime}|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)-f^{\varepsilon}(z^{\prime})|<\delta$.
-This finishes the proof for the tightness of $\{f^{\varepsilon}\}|_{z\in K}$ for each compact set $K$.
-By an identical argument, $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight for each compact set $K$.
-By the Arzelà-Ascoli theorem, for any subsequence of $\varepsilon$, there exists a further subsequence such that $f^{\varepsilon}$ converges uniformly on compacts to a continuous function $f$ on $\mathbb{C}$ such that $f(0)=0,f(1)=1$.
-By using that $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight and taking a further subsequence we may assume that $f$ is a homeomorphism.
+Step 3: Equicontinuity at other points in $K$.
+This step is almost identical to the argument in step 2 by considering walks started from a fixed $z\in K$ with $z\neq 0$.
+From the previous cases, for fixed $\ell$, uniformly in $k$, $\{f^{k,\ell}(z)\}$ is bounded away from 0 and $\infty$.
+Pick $0<\delta_{z}<\min\{|z|/2,1\}$, fix $0<\delta<\frac{1}{2}\min\{|f^{k,\ell}(w)|:k\geq 10,|w-z|<\delta_{z}\}$, $M\geq\sup_{k>0}|f^{k,\ell}(K)|$ and $z^{\prime\prime}\in B(0;2M)\backslash B(0;2\delta)$ with $|z^{\prime\prime}-f^{k,\ell}(z)|>\delta$.
+Let $\gamma_{f^{k,\ell}(z),z^{\prime\prime}}$ be a finite path starting from $f^{k,\ell}(z)$ that first separates $B(0;\delta/4)$ and then $B(z^{\prime\prime};\delta/4)$ from infinity, but does not separate any points in $\mathbb{C}\backslash(B(z^{\prime\prime};{\delta/2})\cup B(0;\delta/2))$ from infinity.
+Then using a similar argument in previous cases, where the Brownian motion has positive probability to stay sufficiently close to $\gamma_{f^{k,\ell}(z),z^{\prime\prime}}$, and the random walks converges in law to Brownian motion in local uniform topology, we see that for each $\delta>0$ there exists $\delta^{\prime}>0$ such that a.s.
+for sufficiently large $k$, $\widetilde{\mathbb{P}}_{k,\ell}^{f^{k,\ell}(z)}[\tau_{0}<\tau_{z^{\prime\prime}}]>\delta^{\prime}$ if $|z^{\prime\prime}-f^{k,\ell}(z)|>\delta$ and $2\delta<|z^{\prime\prime}|<2M$.
+On the other hand, for some $\delta^{\prime\prime}>0$, for all $k$ large enough, $\mathbb{P}_{k,\ell}^{z}[\tau_{z^{\prime}}<\tau_{0};\tau_{z^{\prime}}<\sigma_{\ell}]>1-\delta^{\prime}$ for all $|z^{\prime}-z|<\delta^{\prime\prime}$, which further implies that $\widetilde{\mathbb{P}}_{k,\ell}^{f^{k,\ell}(z)}[\tau_{f^{k,\ell}(z^{\prime})}<\tau_{0}]>1-\delta^{\prime}$.
+This implies a.s.
+for large enough $k$, $|f^{k,\ell}(z^{\prime})-f^{k,\ell}(z)|<\delta$ whenever $|z-z^{\prime}|\leq\min\{\delta^{\prime\prime},\delta_{z}\}$ and $z^{\prime}\in K$.
+This finishes the proof for the tightness of $\{\{f^{k,\ell}\}|_{z\in K}:k\in\mathbb{N},\ell>\widetilde{\ell}+1\}$ for each $\widetilde{\ell}>10$ and $K=[-2^{\widetilde{\ell}-5},2^{\widetilde{\ell}-5}]^{2}$.
+By the Arzelà-Ascoli theorem, for any subsequence of $k$, there exists a further subsequence $k_{n},\ell_{n}\to\infty,$ such that for every $\ell>10$, $f^{k_{n},\ell_{n}}$ converges uniformly on $2^{\ell-5},2^{\ell-5}]^{2}$ to a continuous function $f$ on $\mathbb{C}$ such that $f(0)=0,f(1)=1$.
+By the last sentence of Step 1, we have $f(\mathbb{C})=\mathbb{C}$.
+Using an identical argument as Step 2 and Step 3, for each compact set $K^{\prime}$, for each large enough $\ell$, $\{(f^{k,\ell})^{-1}|_{K^{\prime}}\}$ is also tight.
+By taking a further subsequence we may assume that $f$ is a homeomorphism.
 Then if $(W_{t})_{t\geq 0}$ is a Brownian motion from $z$, $f(W_{t})_{t\geq 0}$ has the law of a planar Brownian motion from $f(z)$ modulo time change.
-Then by Proposition 3.14, $f$ must be either the identity function or the reflection.
-This completes the proof.
+Then by Proposition 3.12, $f$ must be either the identity function or the reflection $f(z)=\overline{z}$.
+To rule out the $f(z)=\overline{z}$ case, consider cells $H_{1},...,H_{m}$ in $\mathcal{H}{(\partial[-2^{k},2^{k}]^{2})}$ such that the edges $\gamma_{e}^{k+1}$ connecting the points $z_{H_{j}}^{k+1}$ form a clockwise loop which is close to $\partial[-2^{k},2^{k}]^{2}$.
+Then the corresponding loop formed by the disks for $H_{1},...,H_{m}$ is also clockwise, which implies that we cannot have $f(z)=\overline{z}$.
 ∎
@@
 In this section, we prove several regularity estimates for the Riemann surface $M(\mathcal{H})$ under a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ that will be applied in the proof of Theorem 1.7.
-In Lemmas 4.34.4, and 4.5, we prove that if we have a microscopic set in $\mathcal{H}$ (resp.
+In Lemmas 4.24.3, and 4.4, we prove that if we have a microscopic set in $\mathcal{H}$ (resp.
 $\varphi(M(\mathcal{H}))$), then the corresponding set in $\varphi(M(\mathcal{H}))$ (resp.
@@
 Using a variant of this argument in the setting of the cell configuration $\mathcal{H}$ and semi-flowers, we can prove various bounds using ergodic averages of the degree of the cell configurations as in Lemma 2.7, which leads to the regularity results in this subsection.
-Figure 5: An illustration of the definition of $M(S;a)$ and $M(S)$.
-For simplicity here we assume that two cells are adjacent if they have nonempty intersection.
-The cells in cyan are those intersecting $\mathring{S}^{a}$, and their neighbors are in yellow.
-We also include the cells that are separated from infinity from the yellow and cyan cells, which are colored in green.
-The surface $M(S;a)\subseteq M(\mathcal{H})$ are formed by vertices for these cells on $M(\mathcal{H})$ as well as the edges connecting these vertices, and the surface $M(S)$ is the largest surface of the form $M(S;a)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$.
-For each square $S$, we define a simply connected surface $M(S)\subseteq M(\mathcal{H})$ as follows.
-For every $a>0$, let $\mathring{S}^{a}$ be the square concentric to $S$ with side length $a$.
-We work on the event that $\mathcal{H}(\mathring{S}^{a})$ is connected; otherwise we set $M(S;a)=\emptyset$.
-Let $M_{0}(S;a)\subseteq M(\mathcal{H})$ be the surface formed by the union of the (closed) equilateral triangles of $M(\mathcal{H})$ with at least one of the vertices in $\mathcal{H}(\mathring{S}^{a})$.
-Note that when viewed as a graph, $M_{0}(S;a)$ is 2-connected since $\mathcal{H}(\mathring{S}^{a})$ is connected.
-In fact, one can verify that, for any infinite plane triangulation $\mathcal{T}$ and a connected subgraph $\mathcal{G}_{1}\subseteq\mathcal{G}$, if we let $\mathcal{G}_{2}$ be the subgraph $\mathcal{G}$ formed by all the vertices in $\mathcal{G}_{1}$ and their neighbors, as well as all the edges connecting these vertices, then $\mathcal{G}_{2}$ is 2-connected.
-We further set $M(S;a)$ to be the union of $M_{0}(S;a)$ and the bounded connected components of $M(\mathcal{H})\backslash M_{0}(S;a)$.
-Then $M(S;a)$ has disk topology.
-Let $M(S)$ be the largest surface of the form $M(S;a)$ for some $a\in(0,|S|)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$.
-Note that these objects are well defined since $\mathcal{H}$ is locally finite.
-See Figure 5 for an illustration.
-Lemma 4.1.
-Almost surely, for each given $\varepsilon>0$ and $a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $M(S;(1-\varepsilon)|S|)\subseteq M(S)$.
-Proof.
-By the connectivity in Definition 1.4, for each $b\in(0,1)$, almost surely for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $\mathcal{H}(\mathring{S}^{b|S|})$ is a connected subgraph of $\mathcal{H}$.
-Let $H\in\mathcal{H}$ be such that its vertex $v_{H}$ on $M(\mathcal{H})$ is in $M(S;(1-\varepsilon)|S|)$.
-Then there are two cases: (i) $H$ is adjacent to some vertex $H^{\prime}\in\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$ or (ii) there exists a loop consisting of vertices for $H_{1},...,H_{n}$ on $M(\mathcal{H})$ surrounding $v_{H}$, such that each $H_{j}$ are adjacent to some cell in $\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$.
-In the setting (i), by Lemma 2.6, if we assume that $k$ is large, then $H\subseteq S$.
-For the case (ii), by Lemma 2.6, for large $k$, $H_{1},...,H_{n}\in\mathcal{H}(\mathring{S}^{(1-3\varepsilon/4)|S|})$.
-Letting $\{z_{H}:H\in\mathcal{H}\}$ and $\{e_{H,H^{\prime}}:H\sim H^{\prime}\}$ be the planar embedding of the map $\mathcal{H}$ as in the almost planarity in Definition 1.5, for large enough $k$, $e_{H_{1},H_{2}},...,e_{H_{n},H_{1}}\subseteq\mathring{S}^{(1-\varepsilon/2)|S|}$ and thus $z_{H}\in\mathring{S}^{(1-\varepsilon/2)|S|}$.
-By the same condition and Lemma 2.6 it follows that $H\subseteq S$.
-This finishes the proof.
-∎
 Let $\mathcal{T}$ be a disk triangulation or an infinite plane triangulation.
@@
 ∎
+For a square $S$, recall the disk triangulation $\mathcal{M}(S)\subseteq\mathcal{M}$ defined in Lemma 2.8, and the Riemann surface $M(S)$ associated with $\mathcal{M}(S)$.
 For each $k>0$, let $\varphi_{k}$ be the uniformization map sending $M(S_{k})$ to $|S_{k}|\mathbb{D}$ such that the vertex on $M(S_{k})$ for the cell containing the center of $S_{k}$ is sent to 0.
@@
 In (4.12), we prove that one can find some $\rho\in[r_{0},r_{1}]$, such that under a conformal map the union of the semi-flowers for the cells intersecting the square concentric to $S$ of side length $\rho$ (shaded in light blue on the right panel) has diameter less than $c(-\log\delta)^{-1/2}|S_{k}|$.
-Further using the line connectivity property and almost planarity in Definition 1.4 and Definition 1.5 gives the claim.
+Further using the line connectivity property and almost planarity condition from Definition 1.4 and Proposition 2.15 gives the claim.
 Proof.
@@
 In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L(\rho)\leq 4c_{2}(-\log\delta)^{-1/2}|S_{k}|$.
-If $\mathring{S}^{\rho}\subseteq\mathring{S}_{k}^{7|S_{k}|/8}$, using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ intersecting $\partial\mathring{S}^{\rho}$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells, and in the planar embedding $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ in Condition 4.
-form a loop around $\mathring{S}^{1.5\delta|S_{k}|}$, and thus encloses all the points $z_{H}$ for $H\in{S}$.
+Now consider the embedding $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ in Proposition 2.15.
+If $\mathring{S}^{\rho}\subseteq\mathring{S}_{k}^{7|S_{k}|/8}$, using the conditions from Definition 1.4 and Proposition 2.15, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ intersecting $\partial\mathring{S}^{\rho}$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells, and in the planar embedding $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ the edges $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ form a loop around $\mathring{S}^{1.5\delta|S_{k}|}$, and thus encloses all the points $z_{H}$ for $H\in{S}$.
 This implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ is separated from infinity by $\cup_{j=1}^{n}\varphi(P_{H_{j}})$, which together with (4.12) implies (4.5).
 If $\mathring{S}^{\rho}\nsubseteq\mathring{S}_{k}^{7{|S_{k}|}/8}$, since $\rho<r_{1}\leq|S_{k}|/4$, $\mathring{S}^{\rho}$ is disjoint from the cell with closest distance to the center of $S_{k}$.
-Using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ either intersecting $\partial\mathring{S}^{\rho}$ or with their vertex on $M(\mathcal{H})$ on the boundary of the surface $M(S_{k})$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells and the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ enclose a loop around all the points $z_{H}$ where $H\in\mathcal{H}(S)$ and the vertex for $H$ is on $M(S_{k})$.
+Using Definition 1.4 and the embedding above, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ either intersecting $\partial\mathring{S}^{\rho}$ or with their vertex on $M(\mathcal{H})$ on the boundary of the surface $M(S_{k})$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells and the edges $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ enclose a loop around all the points $z_{H}$ where $H\in\mathcal{H}(S)$ and the vertex for $H$ is on $M(S_{k})$.
 This together with (4.12) implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ can be separated from 0 by $\partial(|S_{k}|\mathbb{D})$ and some curves of total length $4c_{2}(\log\delta)^{-1/2}|S_{k}|$.
@@
 The following is the analog of Lemma 4.2 in the reverse direction.
-Lemma $4.$
+Lemma 4.3.
 There exists a deterministic constant $c>0$ such that the following is true.
@@
 In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L^{*}(\rho)\leq 4c_{2}(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|$.
-Recall the points $\{z_{H}:H\in\mathcal{H}\}$ and the curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in the planar embedding from the almost planarity condition in Definition 1.5.
+Recall the points $\{z_{H}^{k+3}:H\in\mathcal{V}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ and the curves $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ in the planar embedding from Proposition 2.15.
 Note that by our choice of $r_{1}$, $B(z;\rho)\subseteq|S_{k}|\mathbb{D}$.
-One can find edges $e_{1},...,e_{n}$ forming a loop in $\mathcal{H}$ such that $\gamma_{e_{1}},...,\gamma_{e_{n}}$ make a loop surrounding all the vertices $z_{H}$ with $\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset$, and the union of theses edges has diameter less than $20c_{2}(-\log\delta)^{-1/2}|S_{k}|$.
+One can find edges $e_{1},...,e_{n}$ forming a loop in $\mathcal{H}$ such that $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ make a loop surrounding all the vertices $z_{H}^{k+3}$ with $\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset$, and the union of theses edges has diameter less than $20c_{2}(-\log\delta)^{-1/2}|S_{k}|$.
 Applying  (2.12) once more gives (4.13).
 ∎
-We will also need the version of Lemmas 4.3 and 4.4 for $2^{k},2^{k}]^{2}$ instead of $S_{k}$.
+We will also need the version of Lemmas 4.2 and 4.3 for $2^{k},2^{k}]^{2}$ instead of $S_{k}$.
 With slight abuse of notation, we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$.
@@
 Remark 4.10.
-We note that the proofs of Lemmas 4.3-4.7 can be adapted to the setting of circle packings.
+We note that the proofs of Lemmas 4.2-4.6 can be adapted to the setting of circle packings.
 Therefore, by repeating the arguments in this subsection, it might be possible to give an alternative proof of Proposition 3.9 without using the vertex extremal length in [GJN19].
@@
 Let $\mathcal{H}$ be a random cell configuration satisfying all the conditions in Theorem 1.7 and let $M(\mathcal{H})$ be the Riemann surface constructed from the map associated with $\mathcal{H}$.
-In this section we will define a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ as the limit of functions $\phi_{m}:M(\mathcal{H})\to\mathbb{C}$ for $m\in\mathbb{N}$, which are defined as certain harmonic extensions of an appropriately chosen function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$.
+Recall the function $\phi_{0}$ defined in Section 2.2.
+In this section we will define a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ as the limit of functions $\phi_{m}:M(\mathcal{H})\to\mathbb{C}$ for $m\in\mathbb{N}$, which are defined as certain harmonic extensions of the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$.
 This idea is similar to [GMS22, Section 2.3] and the key difference is that we work with (continuum) harmonic functions on $M(\mathcal{H})$ instead of discrete harmonic functions on the graph $\mathcal{H}$.
-The function $\phi_{0}$ sends each vertex on $M(\mathcal{H})$ to a point in the associated cell of $\mathcal{H}$ and is otherwise piecewise linear.
-More precisely, the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ is defined as follows.
-Given a realization of $\mathcal{H}$, for each cell $H\in\mathcal{H}$, we independently sample a point $\widetilde{c}(H)$ according to the probability measure proportional to the Lebsegue measure on $H$.
-For each triangular face $\Delta$ in $M(\mathcal{H})$ formed by the cells $H_{1},H_{2},H_{3}\in\mathcal{H}$, consider the triangle $\Delta^{\prime}$ in $\mathbb{C}$ formed by connecting $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ using straight line segments, and define $\phi_{0}|_{\Delta}:\Delta\to\Delta^{\prime}$ to be the unique linear function sending the vertices in $\Delta$ to their counterpart in $\Delta^{\prime}$.
-This defines a continuous and piecewise linear function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$.
-See Figure 8 for an illustration.
-Figure 8: An illustration for the definition of the map $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$, which linearly send each triangle on the left panel to the corresponding triangle on the right panel.
-Note that the $\phi_{0}$-images of different equilateral triangles of $M(\mathcal{H})$ might overlap with each other.
 Recall the definition $\widehat{S}_{m}^{z}$ in (2.2) and define the collection of squares
@@
 This is a direct consequence of Lemma 2.6. ∎
-We will also need the following topological properties on $\phi_{0}$.
-Lemma 4.15.
-We have $\phi_{0}(M(\mathcal{H}))=\mathbb{C}$.
-Proof.
-Pick $k$ large enough such that by Definition 1.4, the 4 subgraphs of $\mathcal{H}$ formed by the cells intersecting the 4 sides of $S_{k}$ are connected.
-Since a cell intersecting a vertex of $S_{k}$ is contained in (at least) two of these subgraphs, we get that the set of cells intersecting a side of $S_{k}$ forms a connected graph.
-By further applying Lemma 2.6, we can find a chain of cells $H_{1}\sim H_{2}\sim\cdots\sim H_{n}\sim H_{1}$ in this connected graph such that if we connect $\widetilde{c}(H_{j})$ and $\widetilde{c}(H_{j+1})$ using straight line segments for j=1,,nj=1,...,n with $n+1$ identified with 1, then this creates a loop $\gamma_{n}$ surrounding the square $\widetilde{S}_{k}$ concentric to $S_{k}$ with side length $|S_{k}|/2$, and the winding number of $\gamma_{n}$ around each point in $\widetilde{S}_{k}$ is equal to 1.
-If we consider the vertices for $H_{1},...,H_{n}$ on the surface $M(\mathcal{H})$, then one can find a loop $\widetilde{\gamma}_{n}$ on $M(\mathcal{H})$ connecting these vertices with straight line segments such that $\phi_{0}(\widetilde{\gamma}_{n})=\gamma_{n}$.
-Since $M(\mathcal{H})$ is simply connected, we can continuously shrink $\widetilde{\gamma}_{n}$ to a single point $x$.
-If there exists a point $w\in\widetilde{S}_{k}\backslash\phi_{0}(M(\mathcal{H}))$, then under $\phi_{0}$ image, all of these shrinking loops will have winding number 1 around $w$.
-On the other hand, for loops sufficiently close to $x$, its image under $\phi_{0}$ has winding number 0 around $w$ since $w\notin\phi_{0}(M(\mathcal{H}))$.
-This is a contradiction, which implies that almost surely for large enough $k$, $\widetilde{S}_{k}\subseteq\phi_{0}(\mathcal{H})$, which completes the proof.
-∎
 We call a closed set $D\subseteq M(\mathcal{H})$ a polygon if $\partial D$ is a simple closed curve and can be written as a union of a finite number of line segments on faces of $M(\mathcal{H})$.
@@
 Then we transfer the bound on the the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ to a pointwise bound for $|\phi_{0}-\phi_{\infty}|$.
-The structure of the proof is similar as in [GMS22, Section 2.4], but we need to use a different setup when we conclude the proof of Proposition 4.20 where we we work with the uniformization map $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ as in Lemmas 4.3 and 4.4, and we also need to use regularity results from Section 4.1.
+The structure of the proof is similar as in [GMS22, Section 2.4], but we need to use a different setup when we conclude the proof of Proposition 4.18 where we we work with the uniformization map $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ as in Lemmas 4.2 and 4.3, and we also need to use regularity results from Section 4.1.
 We refer to the paragraph above the latter proof for a proof outline.
@@
 $\limsup_{k\to\infty}\frac{1}{|S_{k}|^{2}}\mathbb{E}\big[\mathrm{Energy}(\phi_{\infty}-\phi_{m_{\varepsilon}};M(S_{k}))|\mathcal{H},\mathcal{D}\big]<e^{-1/\varepsilon^{8}}.$
-Let $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ be the uniformization map defined as in Lemma 4.3 and Lemma $4.$ Using the conformal invariance of Dirichlet energy,
+Let $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ be the uniformization map defined as in Lemma 4.2 and Lemma 4.3. Using the conformal invariance of Dirichlet energy,
 $\limsup_{k\to\infty}\frac{1}{|S_{k}|^{2}}\int_{|S_{k}|\mathbb{D}}\mathbb{E}\Big[|\nabla(\phi_{\infty}\circ\varphi_{k}^{-1})(w)-\nabla(\phi_{m_{\varepsilon}}\circ\varphi_{k}^{-1})(w)|^{2}\big|\mathcal{H},\mathcal{D}\Big]\,dw<e^{-1/\varepsilon^{8}}.$ (4.50)
@@
 This follows from Proposition 4.18 as well as the polynomial growth property in Lemma 4.9.
-Then in Lemmas 4.254.26 and 4.27, we prove that this polynomial must have degree 1, by arguing that if this is not the case, then one can find cells $H_{1},H_{2}$ which are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.3 and Lemma 4.6.
+Then in Lemmas 4.234.24 and 4.25, we prove that this polynomial must have degree 1, by arguing that if this is not the case, then one can find cells $H_{1},H_{2}$ which are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.2 and Lemma 4.5.
 Next we use the ergodic theorem for cell configurations (Lemma 2.4) to prove that for $a,b\in\mathbb{R}$, the average Dirichlet energy of $a\,\mathrm{Re}\,\phi_{\infty}+b\,\mathrm{Im}\,\phi_{\infty}$ is deterministic (Lemma 4.27).
@@
 Fix $\delta\in(0,1)$, and consider a square $S\subseteq S_{k}$ of side length at least $\delta|S_{k}|$.
-For each $z$ contained in some $H\in\mathcal{H}(S)$, if $z$ is inside the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ as in the definition of $\phi_{0}$, then for some $j$, $H_{j}\cap S\neq\emptyset$, and using Lemmas 2.6 and 4.1, it follows that a.s.
+For each $z$ contained in some $H\in\mathcal{H}(S)$, if $z$ is inside the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ as in the definition of $\phi_{0}$, then for some $j$, $H_{j}\cap S\neq\emptyset$, and using Lemmas 2.6 and 2.8, it follows that a.s.
 for large enough $k$, $H_{1},H_{2},H_{3}\subseteq\mathring{S}^{1.5|S|}$ and thus $\phi_{0}^{-1}(z)\subseteq M(\mathring{S}^{2|S|})$.
@@
 The relation (1.5) then follows from (4.89) and (4.93). ∎
-5 Extensions of Theorems 1.7 and 1.8
-In this section, we explain several extensions of Theorems 1.7 and 1.8.
-In Section 5.1, we prove Theorem 5.1, which is a version of Theorems 1.7 and 1.8 with a weaker notion of the line connectivity property in Definition 1.4.
-In Section 5.2, we prove an extension of Theorem 1.7 for a version of circle packing for general planar maps.
-In Section 5.3, we prove an extension of Theorem 1.8 for $p$-angulations with $p\in\{4,5,\dots\}$.
-5.1 An extension with weaker notion of line connectivity
-The following version of Theorem 1.7 and Theorem 1.8, which uses a weaker version of the line connectivity in Definition 1.4, will be used in our future work.
-Theorem 5.1.
-Suppose $\mathcal{H}$ is a cell configuration that satisfies all the conditions in Theorem 1.7 or Theorem 1.8 except for the line connectivity condition in Definition 1.4.
-Suppose there exists a cell configuration $\mathcal{H}^{\rm large}$ satisfying the condition in Definition 1.3 and the moment bound (1.3) along with the following properties:
-  1. 1.
-    The cell configuration $\mathcal{H}^{\rm large}$ is isomorphic to $\mathcal{H}$, in the sense that $\mathcal{H}$ and $\mathcal{H}^{\rm large}$ have isomorphic associated maps; below we let $\phi:\mathcal{H}\to\mathcal{H}^{\rm large}$ denote the isomorphism.
-  2. 2.
-    The convex hull of $H$ is contained in the convex hull of $\phi(H)$ for any $H\in\mathcal{H}$.
-  3. 3.
-    Almost surely, for each given $\varepsilon>0$, there exists $R>0$ such that for $r>R$ and each horizontal or vertical line segment $\ell\subseteq B(0;r)$ with length at least $\varepsilon r$, the following holds.
-For any two cells $H,H^{\prime}\in\mathcal{H}$ intersecting $\ell$, we can find a path $H\sim H_{1}\sim\cdots\sim H_{k}\sim H^{\prime}$, such that $\phi(H_{j})$ intersects $\ell$ for each $j$.
-Then the conclusions in Theorem 1.7 or Theorem 1.8 continue to hold.
-Theorem 5.1 can be viewed as a generalization of Theorem 1.7 and Theorem 1.8 since one can set $\mathcal{H}^{\rm large}=\mathcal{H}$ in the latter setting.
-Proof.
-Most of the proof is identical to that of Theorem 1.7 and Theorem 1.8, and we will only explain the differences.
-Note that the arguments in Section 2.2 do not require the line connectivity property in Definition 1.4 or the almost planarity condition in Defininition 1.5, so the conclusions there continue to hold.
-In particular, by Lemma 2.6 applied to $\mathcal{H}^{\rm large}$ and condition (ii) in the theorem statement, the almost planarity condition in Definition 1.5 holds for $\mathcal{H}^{\rm large}$ as well.
-In the setting of circle packing, the arguments in Section 3 do not directly use the line connectivity property in Definition 1.4.
-The property is used implicitly as we carry the proof of [GMS22, Theorem 1.16]; in the proof in [GMS22], this property is used only in the proof of [GMS22, Lemma 2.19], in the paragraph right after [GMS22, Eq.
-(2.46)].
-To substitute for that, Conditions (ii) and (iii) guarantee that a.s.
-for large enough $k$ and horizontal or vertical line segments $\ell\subseteq S_{k}$ of length $S_{k}$, for any function $g:\mathcal{H}^{\rm large}\to\mathbb{R}$ and $H_{1},H_{2}\in\mathcal{H}(\ell)$,
-$|g(\phi(H_{1}))-g(\phi(H_{2}))|\leq\sum_{H,H^{\prime}\in\mathcal{H}^{\rm large}(\ell):H\sim H^{\prime}}|g(H)-g(H^{\prime})|.$
-The corresponding quantity on the right side can be bounded exactly as [GMS22] since the arguments there only rely on the inputs from Section 2.2, which do not require the line connectivity property.
-[…]
-$L^{\rm large}(\rho)=\sum_{H^{\rm large}\in\mathcal{H}^{\rm large}(\partial\mathring{S}^{\rho})}\mathrm{diam}(\varphi_{k}(P_{H^{\rm large}})).$
-Then  (4.11) holds with $L^{\rm large}(\rho)$ instead of $L(\rho)$ using identical arguments applied to $\mathcal{H}^{\rm large}$, and (4.5) follows from identical arguments using conditions (ii) and (iii).
-For the proof of Lemma 4.15, using conditions (ii) and (iii) along with Lemma 2.6, one can find cells $H_{1}\sim\cdots H_{n}\sim H_{1}$ with identical properties and the conclusion still holds.
-The arguments for Lemma 4.26 are identical, except that at the final step we use that by Lemma 4.1 the set on the right hand side of (4.69) is contained within $\varphi(M(w+[-\delta 2^{k},\delta 2^{k}]^{2}))$, which is a connected subset of the set in (4.66).
-The rest of the proofs in Sections 4.1-4.4 do not require the line connectivity property and can be copied to the setting of Theorem 5.1.
-∎
-5.2 An Extension of Theorem 1.7 to general planar maps
-Let $\mathcal{M}$ be an one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$.
+5 Extensions to general planar maps
+In this section, we prove Theorem 1.8 and Theorem 1.9, extending Theorem 1.6 and Theorem 1.7 to the setting of general planar maps.
+In Section 5.3, we prove Theorem 5.5, which is a version of Theorems 1.6 and 1.7 with a weaker notion of the line connectivity property in Definition 1.4.
+5.1 Proof of Theorem 1.8
+Let $\mathcal{M}$ be a 2-connected simple planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$.
 Recall from Section 1.4 that one can add a vertex $v_{f}$ to every face $f\in\mathcal{F}\mathcal{M}$, and connect the vertices $v_{f}$ to the vertices incident to $f$, to get a triangulation $\mathcal{T}_{\mathcal{M}}$.
@@
 In this section we prove a variant of Theorem 1.6 in this setting.
-Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ as above, and let $\mathcal{T}_{\mathcal{M}}$ be the triangulation generated from $\mathcal{M}$.
-We write $v_{H}\sim f$ if the vertex $v_{H}$ for $H$ in the associated map $\mathcal{M}$ is incident to the face $f\in\mathcal{F}\mathcal{M}$.
-Let
-$\overline{H}_{0}=\bigcup_{f\in\mathcal{F}\mathcal{M}:v_{H_{0}}\sim f}\bigcup_{H\in\mathcal{H}:v_{H}\sim f}H$ (5.1)
-be the union of cells whose vertex is on the same face with $v_{H_{0}}$ and let $d_{0}$ be the maximal degree of the faces incident to $v_{H_{0}}$.
-Theorem 5.2.
-Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ being an one-ended 2-connected simple planar map.
-Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5.
-Further assume
-$\mathbb{E}\Big[\frac{\mathrm{diam}(\overline{H}_{0})^{2}}{\mathrm{area}(H_{0})}\big(\deg(H_{0})+d_{0}\big)^{4}\deg(H_{0})\Big]<\infty,$ (5.2)
-Then the conclusions in Theorem 1.7 holds for the circle packing of $\mathcal{M}$ defined as above.
 […]
@@
 […]
-Lemma $5.$
+Lemma 5.3.
 There are no macroscopic faces of $\mathcal{H}$, in the sense that almost surely for any $\varepsilon>0$, for large enough $k$, for each cell $H_{1},...,H_{j}\in\mathcal{H}$ with $(H_{1}\cup...\cup H_{j})\cap S_{k}\neq\emptyset$ and $v_{H_{1}},...,v_{H_{k}}$ lie on the same face of $\mathcal{M}$, $\mathrm{diam}(H_{1}\cup...\cup H_{j})<\varepsilon|S_{k}|$.
@@
 Since each $\widetilde{H}_{H,k}$ contains the boundary of $H$, we see that the line connectivity property from Definition 1.4 holds.
-The almost planarity condition in Definition 1.5 also holds since one can add vertices and edges directly to the planar map that is close to $\mathcal{H}$ in the sense of Definition 1.5 and there are no macroscopic faces of $\mathcal{H}$ by Lemma $5.$ Furthermore, the moment bound (1.3) for $\widetilde{\mathcal{H}}$ follows from the moment bound (5.2).
+Furthermore, the moment bound (1.2) for $\widetilde{\mathcal{H}}$ follows from the moment bound (1.7).
 Therefore the conditions in Theorem 1.6 hold for $\widetilde{\mathcal{H}}$ and the conclusion follows from Theorem 1.6.
 ∎
-5.3 Extension of Theorem 1.8 to $p$-angulations
-In Theorem 1.8, we worked with the setting where the associated map $\mathcal{M}$ of the cell configuration $\mathcal{H}$ is a triangulation.
-In this section, we briefly explain how to extend Theorem 1.8 to the setting where $\mathcal{M}$ is a $p_{0}$-angulation for $p_{0}\geq 3$.
-Let $\mathcal{M}$ be a one-ended planar map with no boundary.
-Assume that $\mathcal{M}$ has no loops, and every face of $\mathcal{M}$ is adjacent to $p_{0}$ vertices.
-We let $M(\mathcal{M})$ be the Riemann surface constructed in the same way as in the triangulation settings in Section 2.1, where we view each face of $\mathcal{M}$ as a regular $p_{0}$-gon with unit side length.
+5.2 Proof of Theorem 1.9
+Let $\mathcal{M}$ be a planar map with whole-plane topology and let $\mathcal{H}$ be a cell configuration whose associated map is $\mathcal{M}$ as in Theorem 1.9.
+We assume $\mathcal{M}$ only has faces of degree at least 3; we may assume this upon removing loops and collapsing degree 2 faces as described above the theorem statement and observing that if the original map satisfies the assumptions of the theorem then the modified map with only face degrees at least 3 also satisfies these conditions.
+Let $\widetilde{\mathcal{T}}_{\mathcal{M}}$ be the triangulation generated by $\mathcal{M}$ as follows; see Figure 9 for an illustration.
+We add a vertex $v_{f}$ to each face $f\in\mathcal{F}\mathcal{M}$ and a vertex $v_{e}$ to each edge $e\in\mathcal{E}\mathcal{M}$.
+Adding the vertices $v_{e}$ cause each edge $e$ to be split into two edges, i.e., $e$ is replaced by two edges each connecting one end-point of $e$ and $v_{e}$.
+For each face $f$, we connect $v_{f}$ to all the vertices on $\mathcal{M}$ incident to the face $f$ and all the vertices $v_{e}$ where $e$ is an edge of $f$.
+Observe that $\widetilde{\mathcal{T}}_{\mathcal{M}}$ is an infinite plane triangulation without loops and multi-edges.
+Recall the construction of the cell configuration $\widetilde{\mathcal{H}}$ defined in terms of $\mathcal{H}$ as in Section 5.1.
+Let $d_{H}$ be the number of faces and edges of $\mathcal{M}$ incident to the vertex $v_{H}$.
+Let $(\widetilde{H}_{H,k})_{k=0,...,d_{H}}$ be the partition of the cell $H\in\mathcal{H}$ as defined in Section 5.1, except that we partition each cell into $d_{H}+1$ parts instead of $\deg(H)+1$ parts.
+For each edge $e$ (resp.
+face $f$) of $\mathcal{M}$ such that $v_{H}$ is on $e$ (resp.
+$f$), we uniformly randomly set $\widetilde{H}_{H,e}$ (resp.
+$\widetilde{H}_{H,f}$) to be equal to one of $\widetilde{H}_{H,k}$ for $k=1,...,d_{H}$.
+For each vertex $v_{f}$ (resp.
+$v_{e}$) added to generate $\mathcal{T}_{\mathcal{M}}$, let $\widetilde{H}_{f}=\cup_{H:v_{H}\sim f}\widetilde{H}_{H,f}$ (resp.
+$\widetilde{H}_{e}=\cup_{H:v_{H}\sim e}\widetilde{H}_{H,e}$), and let $\widetilde{H}_{H}=\widetilde{H}_{H,0}$.
+This defines a cell configuration $\widetilde{\mathcal{H}}$ with cells $\widetilde{H}_{f}$, $\widetilde{H}_{e}$ and $\widetilde{H}_{H}$ whose associated map is $\widetilde{\mathcal{T}}_{\mathcal{M}}$.
+We further define the conductance of edges $\widetilde{e}$ of $\widetilde{\mathcal{T}}_{\mathcal{M}}$ as follows.
+If $\widetilde{e}$ is connecting a vertex of $\mathcal{M}$ and a vertex $v_{e}$ in $\mathcal{V}\widetilde{\mathcal{T}}_{\mathcal{M}}$, we set $\mathfrak{c}(\widetilde{e})=1$.
+If $\widetilde{e}$ is connecting a vertex of $\mathcal{M}$ and a vertex $v_{f}$ in $\mathcal{V}\widetilde{\mathcal{T}}_{\mathcal{M}}$, we set $\mathfrak{c}(\widetilde{e})=\deg(f)+2$.
+If $\widetilde{e}$ is connecting two vertices $v_{e}$ and $v_{f}$, we set $\mathfrak{c}(\widetilde{e})=\deg(f)+1.5$.
+Note that by our choice of the conductance, the planar map $\mathcal{M}$ is measurable with respect to $(\widetilde{\mathcal{T}}_{\mathcal{M}},\mathfrak{c})$.
+Recall the definition of semi-flowers in Section 4.1 in the setting where $\mathcal{M}$ is a triangulation.
+For a vertex $v$ of $\widetilde{\mathcal{T}}_{\mathcal{M}}$, we define the semi-flower $P_{v}\subseteq M(\mathcal{M})$ as follows.
+First suppose $v$ is also a vertex of $\mathcal{M}$, let $f$ be a face incident to $v$, and let $p\geq 3$ denote the degree of $f$.
+Let $Q_{f}$ be the unit side length regular $p$-gon associated with $f$, and we place $Q_{f}$ on the plane.
+Let $v_{1},...,v_{j_{v,f}}$ be vertices of $Q_{f}$ such that as we glue the polygon $Q_{f}$ according to the adjacency pattern of $\mathcal{M}$ as in the construction of $M(\mathcal{M})$, these vertices are identified with $v$.
+For each such vertex $v_{i}$, let $e_{1},e_{2}$ be the edges of $Q_{f}$ incident to $v_{i}$.
+We mark the points $x_{1},x_{2}$ on $e_{1},e_{2}$ such that they have distance $1/4$ from $v_{i}$.
+Let $e_{1}^{\perp},e_{2}^{\perp}$ be line segments started from $x_{1},x_{2}$ of length $\frac{1}{6\sqrt{3}}$ (the choice of this number is unimportant as long as it is less than $\frac{1}{4\sqrt{3}}$), perpendicular to $e_{1},e_{2}$, and pointing inwards to $Q_{f}$.
+Let $e_{12}$ be the line segment connecting the endpoints of $e_{1}^{\perp},e_{2}^{\perp}$ other than $x_{1},x_{2}$, and let $P_{v,f,i}\subseteq M(\mathcal{M})$ be the part of $Q_{f}$ enclosed by $e_{1},e_{2},e_{1}^{\perp},e_{2}^{\perp}$ and $e_{12}$ and we view $Q_{f}$ as part of $M(\mathcal{M})$.
+We let $P_{v}=\cup_{f\in\mathcal{F}\mathcal{M}:v\sim f}\cup_{i=1}^{j_{v,f}}P_{v,f,i}\subseteq M(\mathcal{M})$.
+Next suppose $v=v_{e}$ where $e$ is an edge on the face $f$ of $\mathcal{M}$.
+First assume $e$ is incident to $f$ and another face $\widetilde{f}$.
+Let $Q_{f}$ be the regular $p$-gon placed on the plane as before.
+Let $x_{3},x_{4}$ be points on $e$ of distance $1/4$ from the midpoint of $e$.
+Let $e_{3}^{\perp},e_{4}^{\perp}$ be line segments perpendicular to $e$ of length $\frac{1}{6\sqrt{3}}$, started from $x_{3},x_{4}$ and pointing inwards to $Q_{f}$.
+Let $e_{34}$ be the line segment connecting the endpoints of $e_{3}^{\perp},e_{4}^{\perp}$ other than $x_{3},x_{4}$.
+Let $P_{e,f}\subseteq M(\mathcal{M})$ be the corresponding part of enclosed by $e,e_{3}^{\perp},e_{4}^{\perp},e_{34}$ when we view $Q_{f}$ as part of $M(\mathcal{M})$.
+Define $P_{e,f^{\prime}}$ analogously, and let $P_{v_{e}}=P_{e,f}\cup P_{e,f^{\prime}}$.
+If $e$ is incident to the face $f$ only, then there is also some other edge $e^{\prime}$ of $Q_{f}$ identified with $e$ in our gluing of $Q_{f}$; in this case we define $P_{e,f}$ and $P_{e^{\prime},f}$ by drawing line segments on $Q_{f}$ in the same way, and let $P_{v_{e}}=P_{e,f}\cup P_{e^{\prime},f}$.
+Finally for a vertex $v_{f}$ where $f$ is a face of $\mathcal{M}$ of degree at least 3, let $P_{v_{f}}=Q_{f}\backslash\cup_{v\in\mathcal{V}\mathcal{M}:v\sim f}P_{v,f}$ be the remaining parts of $Q_{f}$.
+See Figure 9 for an illustration.
+Refer to caption Refer to caption
+Figure 9: Left: An illustration of part of $\widetilde{\mathcal{T}}_{\mathcal{M}}$, where the black edges are edges of the original map $\mathcal{M}$, and the brown edges are the edges added to construct $\widetilde{\mathcal{T}}_{\mathcal{M}}$.
+Right: An illustration of the semi-flowers on the surface $M(\mathcal{M})$.
+The green regions depict the semi-flowers $P_{v}$ where $v$ is a vertex of $\mathcal{M}$, the yellow regions depict the semi-flowers $P_{v_{f}}$ for $f\in\mathcal{F}\mathcal{M}$, and the blue regions are for the semi-flowers $P_{v_{e}}$ with $e\in\mathcal{E}\mathcal{M}$.
+Lemma 5.4.
+Let $M_{0}\subseteq M(\mathcal{M})$ be either equal to $M(\mathcal{M})$ or equal to the closure of a union of some faces of $\mathcal{T}_{\mathcal{M}}$ on $M(\mathcal{M})$ such that $M_{0}$ is homeomorphic to the closed unit disk.
+The conclusions of Lemma 4.1 hold for a conformal map $\varphi:M_{0}\to\mathbb{C}$ or $\varphi:M_{0}\to\mathbb{D}$ and the semi-flowers defined as above.
+Proof.
+Let $v$ be a vertex of $\mathcal{T}_{\mathcal{M}}$ on $M_{0}$.
+We only consider the setting where $v$ is in the interior of $M_{0}$; the boundary case is similar using Schwarz reflection.
+First suppose $v\in\mathcal{V}\mathcal{M}$.
+Let $r\in(0,\frac{{1}}{2})$.
+We define a cone $F_{v}(r)$ as follows.
+For each face $f$ incident to $v$ whose associated regular polygon is $Q_{f}$, let $F_{v,f}(r)$ be the intersection of $Q_{f}$ and the radius $r$ disks centered at the vertex $v_{i}$, where $v_{1},...,v_{j}$ are the vertices identified with $v$ in the gluing of $Q_{f}$.
+Let $F_{v}(r)=\cup_{f\in\mathcal{F}\mathcal{M}:v\sim f}F_{v,f}(r)$ and observe that $F_{v}(r)$ has the same geometry of a Euclidean cone of radius $r$ and some angle $\alpha_{v}\in[\frac{\pi}{3}\deg(v),\pi\deg(v))$, where we use that these disks are disjoint since $r<\frac{1}{2}$.
+Furthermore, from our construction $F_{v}(\frac{1}{16\sqrt{3}})\subseteq P_{v}\subseteq F_{v}(\frac{1}{{2\sqrt{3}}})\subseteq F_{v}(\frac{1}{3})\subseteq M(\mathcal{M})$.
+Therefore under the conformal map $\varphi_{v}:=z\mapsto(3{z})^{\frac{2\pi}{\alpha_{v}}}$, there are some universal constants $c_{0},c_{1}>0$ such that $B(0;c_{0})\subseteq\varphi_{v}(P_{v})\subseteq B(0;1-c_{1}\deg(v)^{-1})$, so the claim (i.e., (4.1)) follows from the Koebe distortion theorem.
+Now suppose $v=v_{f}$ where $f\in\mathcal{F}\mathcal{M}$ has degree $p\geq 3$.
+We place $Q_{f}$ on the plane such that $Q_{f}$ is centered at 0.
+Then
+$B\Big(0;\cos(\frac{\pi}{p})(\frac{1}{2}\cot(\frac{\pi}{p})-\frac{1}{2\sqrt{3}})\Big)\subseteq P_{v_{f}}\subseteq B\Big(0;\sqrt{(\frac{1}{4})^{2}+(\frac{1}{2}\cot(\frac{\pi}{p})-\frac{1}{6\sqrt{3}})^{2}}\Big)\subseteq B(0;\frac{1}{2}\cot(\frac{\pi}{p}))\subseteq Q_{f},$
+and in particular there are universal constants $c_{2},c_{3},c_{4}>0$ such that for some $r>0$, $B(0;c_{2}r)\subseteq P_{v_{f}}\subseteq B(0;c_{3}r(1-c_{4}p^{-1}))\subseteq B(0;c_{3}r)\subseteq Q_{f}$.
+The claim follows again from the Koebe distortion theorem.
+Finally, let $v=v_{e}$ for some edge $e\in\mathcal{E}\mathcal{M}$.
+Then from our construction of $M(\mathcal{M})$, on the surface $M(\mathcal{M})$, $P_{v_{e}}$ is a rectangle of side lengths $\frac{1}{2}$ and $\frac{1}{3\sqrt{3}}$ contained in the interior of a quadrangle of side length $\frac{1}{\sqrt{3}}$ and angles $\pi/3,2\pi/3,\pi/3,2\pi/3$, and the claim is again a consequence of the distortion theorem.
+∎
+Proof of Theorem 1.9.
+Let $\widetilde{\mathcal{H}}$ be the cell configuration constructed as above. Then the moment bound (1.7) implies that in the cell configuration $\widetilde{\mathcal{H}}$,
+$\mathbb{E}\big[\frac{\mathrm{diam}(\widetilde{H}_{0})^{2}}{\mathrm{area}(\widetilde{H}_{0})}\deg(\widetilde{H}_{0})^{4}\big]<\infty;\ \ \mathbb{E}\big[\frac{\mathrm{diam}(\widetilde{H}_{0})^{2}}{\mathrm{area}(\widetilde{H}_{0})}(\sum_{e:H_{0}\sim e}\mathfrak{c}(e))\big]<\infty.$ (5.5)
+Furthermore, using identical arguments in Section 5.1, $\widetilde{\mathcal{H}}$ satisfies all the other conditions in Theorem 1.7.
+In particular, by the same arguments as in [GMS22, Lemma 2.10], for some deterministic constant $C>0$, for all large enough $k$,
+$\sum_{H\in\widetilde{\mathcal{H}}(S_{k})}\mathrm{diam}(H)^{2}(\sum_{e:H_{0}\sim e}\mathfrak{c}(e))<C.$ (5.6)
+Now the identical arguments in Section 4.1 for $\widetilde{\mathcal{H}}$ carry through once we use Lemma 5.4 instead of Lemma 4.1, and we define the function $\phi_{0}$ in Section 4.2 for $\widetilde{\mathcal{H}}$ in the identical way.
+It is elementary to prove that, for $H_{1},H_{2},H_{3}\in\widetilde{\mathcal{H}}$ a triangle $\Delta_{H_{1},H_{2},H_{3}}$ on $M(\mathcal{H})$ where $H_{3}$ is the cell for some vertex $v_{f}$, $\mathrm{Energy}(\phi_{0};\Delta_{H_{1},H_{2},H_{3}})\leq c_{0}\deg(H_{3})(\mathrm{diam}(H_{1})^{2}+\mathrm{diam}(H_{2})^{2}+\mathrm{diam}(H_{3})^{2})$ for some absolute constant $c_{0}>0$, which combined with (5.6) further implies the bound (4.34), i.e., the bound for the Dirichlet energy of $\phi_{0}$ as in Lemma 4.15.
+The rest of the proof is identical to that of Theorem 1.7.
+∎
+5.3 An extension with weaker notion of line connectivity
+The following version of Theorems 1.6-1.9, which uses a weaker version of the line connectivity in Definition 1.4, will be used in our future work.
 Theorem 5.5.
-Let $p_{0}\geq 3$.
-Let $\mathcal{H}$ be a random cell configuration such that its associated map is a one-ended $p_{0}$-angulation as above.
-Let $M(\mathcal{H})$ be the Riemann surface associated with the associated map of $\mathcal{H}$ defined as above.
-Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5.
-Further assume $\mathcal{H}$ satisfies the moment bound (1.3).
-Then the conclusions in Theorem 1.8 still holds.
+Suppose $\mathcal{H}$ is a cell configuration that satisfies all the conditions in one of the settings of Theorems 1.6-1.9 except for the line connectivity condition in Definition 1.4.
+Suppose there exists a cell configuration $\mathcal{H}^{\rm large}$ satisfying the condition in Definition 1.3 and the moment bound (1.2) along with the following properties:
+  1. (i)
+    The cell configuration $\mathcal{H}^{\rm large}$ is isomorphic to $\mathcal{H}$, in the sense that $\mathcal{H}$ and $\mathcal{H}^{\rm large}$ have isomorphic associated maps; below we let $\phi:\mathcal{H}\to\mathcal{H}^{\rm large}$ denote the isomorphism.
+  2. (ii)
+    The convex hull of $H$ is contained in the convex hull of $\phi(H)$ for any $H\in\mathcal{H}$.
+  3. (iii)
+    Almost surely, for each given $\varepsilon>0$, there exists $R>0$ such that for $r>R$ and each horizontal or vertical line segment $\ell\subseteq B(0;r)$ with length at least $\varepsilon r$, the following holds.
+For any two cells $H,H^{\prime}\in\mathcal{H}$ intersecting $\ell$, we can find a path $H\sim H_{1}\sim\cdots\sim H_{k}\sim H^{\prime}$, such that $\phi(H_{j})$ intersects $\ell$ for each $j$.
+Then the conclusions in Theorem 1.6-1.9 continue to hold.
+Theorem 5.5 can be viewed as a generalization of Theorem 1.6-1.9 since one can set $\mathcal{H}^{\rm large}=\mathcal{H}$ in the latter setting.
 Proof.
-The proof is almost identical to the proof of Theorem 1.8.
-The results in Section 2.2 still holds since we did not assume there that the associated map of $\mathcal{H}$ is a triangulation.
-To derive the regularity estimates in Section 4.1, one define the flowers and semi-flowers in the same way as in the triangulation case.
-It is straightforward to check that Lemma 4.2 still holds in this setting, and the rest of the proof in Section 4.1 still carries.
-To define the function $\phi_{0}$ in Section 4.2, we sample the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$ in the same way.
-For each regular $p_{0}$-gon on $M(\mathcal{H})$ formed by the cells $H_{1},...,H_{p_{0}}$, we further uniformly at random choose a way to draw straight lines to partition this face into $p_{0}-2$ triangles, and let $\phi_{0}$ be the piecewise linear function sending each such triangle on $M(\mathcal{H})$ to the corresponding triangle formed by the points $\{\widetilde{c}(H):H\in\mathcal{H}\}$.
-Then one can check that for each face on $M(\mathcal{H})$ formed by $H_{1},...,H_{p_{0}}$, the Dirichlet energy of $\phi_{0}$ on this face is controlled by $C(p_{0})(\mathrm{diam}(H_{1})^{2}+....+\mathrm{diam}(H_{p_{0}})^{2})$ where $C(p_{0})$ is a constant depending only on $p_{0}$.
-This will allow us to bound the Dirichlet energy per area for $\phi_{0}$, and the rest of the proof is identical.
+[…]
+$|g(\phi(H_{1}))-g(\phi(H_{2}))|\leq\sum_{H,H^{\prime}\in\mathcal{H}^{\rm large}(\ell):H\sim H^{\prime}}|g(H)-g(H^{\prime})|.$
+The right side can be bounded exactly as [GMS22] since the arguments there only rely on the inputs from Lemma 2.3-2.7, which do not require the line connectivity property.
+Once this is done, the arguments in Section 2.3 and Section 3 work, except that in the construction of the map $\mathcal{M}_{k,\ell}$ in Lemma 2.14, we connect the cells $H_{j}$ and $H_{j+1}$ there using cells $H$ such that $\phi(H)$ intersects $L_{j,j+1}$.
+In the setting of the Riemann uniformization, Definition 1.4 is used a few times. Let us explain the modifications. To prove Lemma 4.2 in this setting, we define
+$L^{\rm large}(\rho)=\sum_{H^{\rm large}\in\mathcal{H}^{\rm large}(\partial\mathring{S}^{\rho})}\mathrm{diam}(\varphi_{k}(P_{H^{\rm large}})).$
+Then  (4.11) holds with $L^{\rm large}(\rho)$ instead of $L(\rho)$ using identical arguments applied to $\mathcal{H}^{\rm large}$, and (4.5) follows from identical arguments using conditions (ii) and (iii).
+The arguments for Lemma 4.24 are identical, except that at the final step we use that by Lemma 2.8 the set on the right hand side of (4.69) is contained within $\varphi(M(w+[-\delta 2^{k},\delta 2^{k}]^{2}))$, which is a connected subset of the set in (4.66).
+The rest of the proofs in Sections 4.1-4.4 do not require the line connectivity property and can be copied to the setting of Theorem 5.5.
 ∎